### v1

# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

Abstract
Missing data often significantly hamper standard time series analysis, yet in practice they are frequently encountered. In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series. Unlike existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for potential nonlinear dynamics. Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any missing pattern. In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated. As a key advantage, our method mitigates the distributional bias typical of many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series. Leveraging the benign landscape of the optimization formulation, we establish the convergence of an alternating minimization algorithm to critical points. Furthermore, we provide conditions under which the marginal distributions of the underlying time series can be identified. Our numerical experiments, including extensive simulations covering linear and nonlinear time series models and an application to a real-world groundwater dataset laden with missing data, corroborate the practical usefulness of the proposed method.
Keywords— Missing data, Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
1 Introduction
In many scientific fields, data collection technology is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data-collection policies, leading to missing entries in the data (Little and Rubin,, 2019). In time series analysis, missing data pose serious challenges to the application of most statistical tools since it is typically presumed that observations are available at equally spaced time points. Moreover, since the dynamic relationship is of crucial importance in time series analysis, one can not proceed by throwing away the missing data, even if the missing data appear completely at random (i.e., the missing pattern is independent of the series).
To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods can be readily applied for statistical inference in the downstream. Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021). Most of these methods are iterative or EM-type algorithms that alternate between model identification and missing data imputation (see, for example, Chpt. 4 of Peña and Tsay,, 2021). For certain time series models, one can estimate the model parameters simultaneously with the missing data (Harvey and Pierse,, 1984; Gómez et al.,, 1999). Besides the optimal interpolator, ad hoc deterministic imputation methods, such as smoothing with linear interpolation, B-splines and wavelets (see, for instance, Hastie et al.,, 2009), are also used in practice for their simplicity.
However, the optimal interpolator can introduce distributional biases that are detrimental to downstream statistical analysis. To illustrate, consider an AR(1) model, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d. standard normal random variables. It is well-known that the optimal interpolator can be expressed in terms of the inverse autocorrelation function (Peña and Maravall, 1991; see also Chpt. 6 of Peña et al.,, 2011). For instance, suppose that data at time $T$ and $T+1$ are missing; then the optimal interpolators are given by
$\hat{x}_{T}$ $=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},$
$\hat{x}_{T+1}$ $=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$ for $h\geq 0$, it is not difficult to show
$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),$
and
$|\mathbb{E}(\hat{x}_{T}\hat{x}_{T+1})|=\frac{2\phi^{2}+3\phi^{4}+3\phi^{6}+\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}|\mathbb{E}(x_{t}x_{t+1})|<|\mathbb{E}(x_{t}x_{t+1})|.$
Hence, the variance and the autocovariances of the imputed data are different from those of the original AR(1) process. Intuitively, the optimal interpolator for consecutive missing data tends to predict the unconditional mean, leading to a downward bias (in magnitude) of the variances. In general, the joint distribution of the optimally interpolated time series is often quite different from that of the original process, leading to questionable downstream statistical analysis, especially when the number of missing data is non-negligible. Similar biases are also observed for some deterministic imputation methods (Carrizosa et al.,, 2013).
In addition to the distributional bias, the above example also reveals further predicaments of the optimal interpolator. First, in order to compute the optimal interpolator, one often needs to choose a model class (such as the ARMA models) before imputation. However, such decision relies heavily on domain expertise and may be ill-informed when there are many missing values. Second, it may be difficult to compute the optimal interpolator if some nonlinear model is employed. With different missing patterns and different chosen models, the computational procedure may change drastically. Finally, incorporating side information about the missing data into this methodology appears less straightforward. Side information about the missing data restricts the set of admissible imputations. For instance, compositional time series is known to lie in $[0,1]$ and sums to unity at each time point. Besides the optimal interpolator, it is well-known that some deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges (Carrizosa et al.,, 2013).
In this paper, we present a novel procedure for time series imputation, referred to as the temporal Wasserstein imputation (TWI), which circumvents the aforementioned limitations. As a nonparametric method, TWI is highly versatile and can be applied to a wide range of time series data. In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive model (TAR, Tong,, 1983). Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm. This principled algorithmic approach allows for the seamless integration of side information by restricting the admissible imputations, and its extension to multivariate time series is also straightforward. The optimization and statistical underpinnings of TWI are closely related to the optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013), which is briefly reviewed in Section 2. Section 3 maps out the optimization formulation and introduces the computational scheme in detail.
The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the $p$-dimensional marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series (which is stable under stationarity), thereby lowering the distributional bias discussed above. In Section 4, we investigate the optimization and statistical properties of TWI. Our analysis shows that the optimization landscape of TWI is biconvex and the alternating algorithm is guaranteed to converge to a critical point. However, from the statistical viewpoint, missing data imputation suffers from an identification issue: Many distributions (of the stochastic process) may be equally plausible in describing the corrupted series. This raises the question of when and under what conditions TWI can identify the underlying distribution and yield imputations that can be safely used in downstream analysis. In Section 4, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation. The latter condition motivates a simple modification of TWI, termed $k$-TWI, which shows further improvement in imputation quality from TWI in our numerical experiments.
In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings. Across all models, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates. Its advantage is particularly prominent in the presence of nonlinear dynamics, where TWI consistently outperforms the benchmark methods. In Section 6, we apply TWI to a real-world groundwater dataset with many missing data. After imputation, we conduct factor analysis and autoregressive modeling in the downstream. Compared to existing methods, TWI effectively captures the underlying dynamics, producing more reasonable factor estimates as well as more accurate autoregressive predictions. These numerical results underscore the practical utility of TWI.
2 Preliminary: Optimal transport
In this section, we give a brief introduction to the optimal transport (OT) problem, which serves as a prerequisite to the rest of the paper. For a more comprehensive account on the theory of optimal transport, the readers are referred to Villani, (2021, 2008).
Given two probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{p}$, and a cost function $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{+}$, the optimal transport cost between $\mu$ and $\nu$ is defined as
[…]
If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$ for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$. Namely,
[…]
The Wasserstein distance is a metric on the space of probabilities with finite $k$-th moment on $\mathbb{R}^{p}$, and metrizes the weak topology on this space (Villani,, 2021, Chpt. 9). In particular, if $\mu$ and $\{\mu_{t}:t=1,2,\ldots\}$ are probability measures on $\mathbb{R}^{p}$ such that $\mathcal{W}_{k}(\mu_{t},\mu)\rightarrow 0$ as $t\rightarrow\infty$, then $\mu_{t}\Rightarrow\mu$, where $\Rightarrow$ denotes the weak convergence of probability measures (Billingsley,, 1995). Therefore, it can be used as a discrepancy metric between $\mu$ and $\nu$. In fact, the Wasserstein distance offers several advantages over other discrepancy metrics such as the Kullbeck-Leibler divergence or the Kolmogorov-Smirnov distance. In addition to its metric properties, the Wasserstein distance has a nice dual representation and handles support mismatch easily, which makes it suitable in modern machine learning and statistical applications.
If $\mu$ and $\nu$ are discrete measures, for instance $\mu=\sum_{i=1}^{n}p_{i}\delta_{\mathbf{x}_{i}}$ and $\nu=\sum_{j=1}^{m}q_{j}\delta_{\mathbf{y}_{j}}$, where $\{\mathbf{x}_{i}\}_{i=1}^{n},\{\mathbf{y}_{j}\}_{j=1}^{m}$ are in $\mathbb{R}^{p}$, $\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{m}q_{j}=1$, $p_{i}>0$, $q_{j}>0$ for all $i,j$, and $\delta_{\mathbf{x}}$ denotes the dirac measure at $\mathbf{x}$, then the minimization problem becomes the following linear program.
$\mathsf{OTC}(\mu,\nu)=\min_{\{\pi_{ij}\}}$ $\sum_{i,j}L(\mathbf{x}_{i},\mathbf{y}_{j})\pi_{ij}$
s.t. $\sum_{i=1}^{n}\pi_{ij}=q_{j},\quad j=1,\ldots,m,$ (1)
$\sum_{j=1}^{m}\pi_{ij}=p_{i},\quad i=1,\ldots,n,$ (2)
$\pi_{ij}\geq 0.$
The constraints (1) and (2) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints. In the discrete case, the optimal coupling $\{\pi_{ij}\}$, in general, cannot be expressed as the push-forward under an one-to-one map. Instead, the optimal coupling typically involves split of masses for discrete measures. Many computational methods have been proposed to solve this linear program, including the network flow simplex (Orlin,, 1997) and the auction method (Bertsekas,, 1992). When the problem dimensions $n,m$ are large, standard tools for linear programs may suffer from slow computation. Nevertheless, recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
3 Temporal Wasserstein imputation
In this section, we introduce temporal Wasserstein imputation (TWI), the proposed method for time series imputation, through an optimal transport formulation. First, an optimization formulation of the proposed method that directly connects to the optimal transport problem is laid out. Then, we describe an alternating minimization algorithm to impute the missing values. The algorithm also reveals some connections between the proposed method and existing approaches.
3.1 Optimization formulation
Let $\{\tilde{x}_{t}\}$ be the time series of interest, and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$. Denote by $\bigstar$ a generic missing value, and $\mathcal{M}$ the set of time indices corresponding to the missing observations. Then, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise. If $\{\tilde{x}_{t}\}$ is stationary, it follows that its $p$-dimensional marginal distributions, i.e. the probability distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, are stable over time. Let $\mu_{p}$ ($p\in\mathbb{N}$) be the probability measure associated with $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$. Stationarity implies $\mu_{p}$ is independent of the time index $t$.
In light of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distributions remain nearly unchanged before and after a specified time point $n_{1}$. More specifically, we first divide the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$, where $p-1<n_{1}<n-p$, and define the two empirical marginal distributions
$\hat{\mu}_{\mathrm{pre}}(\mathbf{w})=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})},$
$\hat{\mu}_{\mathrm{post}}(\mathbf{w})=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}.$
The discrepancy between the two empirical distributions can then be measured by the optimal transport cost, which is defined as the minimum value of the following linear program.
$\begin{split}\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij}}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t. }&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\\ &\pi_{ij}\geq 0\mbox{ for all $i,j$},\end{split}$ (3)
where $\mathbf{v}_{i}(\mathbf{w}):=(w_{i},w_{i-1},\ldots,w_{i-p+1})^{\top}$, and, for ease of notation, we enumerate the rows and columns of the $(n_{1}-p+2)\times(n-n_{1}-1)$ matrix $\mathbf{\Pi}$ by $\{p-1,p,\ldots,n_{1}\}$ and $\{n_{1}+1,\ldots,n-1\}$, respectively. An important special case of (3) is when $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{k}$ for some $k\geq 1$. In this case, we have $\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})=\mathcal{W}_{k}^{k}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})$.
Our goal is to find the imputation that minimizes the aforementioned discrepancy. Thus, we solve for
$\hat{\mathbf{w}}:=(\hat{w}_{0},\ldots\hat{w}_{n-1})^{\top}=\arg\min_{\mathbf{w}\in\mathcal{C}}\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w})),$ (4)
where the constraint set $\mathcal{C}$ delineates the admissible imputations. Typically, it enforces the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. It may also reflect the side information available. For example, if $x_{t}$ are weekly sales data, the knowledge of monthly sales define a set of linear constraints for the missing values. Another example is when $\{\mathbf{x}_{t}\}$ is a $d$-dimensional compositional time series such as the market share of several companies in an industry. It follows that $\mathbf{x}_{t}^{\top}\mathbf{1}=1$ for all $t$. Then, by defining $\mathcal{C}=\{\mathbf{W}\in\mathbb{R}^{n\times d}:\mathbf{W}\mathbf{1}_{d}=\mathbf{1}_{n}\}$, where $\mathbf{1}_{n}=(1,\ldots,1)^{\top}\in\mathbb{R}^{n}$, the set of admissible imputations are restricted to be compositional as well.
One may also consider the following regularized version of (4).
$\hat{\mathbf{w}}:=(\hat{w}_{0},\ldots\hat{w}_{n-1})^{\top}=\arg\min_{\mathbf{w}\in\mathcal{C}}\left\{\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))+\frac{\lambda}{2}\|\mathbf{w}\|^{2}\right\},$ (5)
where $\lambda>0$ is small. As we will see in the next subsection, the $\ell_{2}$-regularization term makes a subproblem in the alternating minimization algorithm strongly convex, enhancing numerical stability. In addition, in the important special case corresponding to the Wasserstein distance of order 2, i.e., $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{2}$, the regularization allows one to solve the subproblem by finding a fixed point of a linear system. Since $L$ is often chosen to be some power of the Euclidean norm, which induces the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (4) or (5) is called the temporal Wasserstein imputation (TWI) throughout this paper.
In this formulation, it is clear that the temporal Wasserstein imputation, which seeks to equate the marginal distributions on both sides of $n_{1}$, is nonparametric. In addition, with a mildly large $p$, the $p$-dimensional marginal distributions can fully capture the nonlinear dynamics of a Markov process (such as a finite-order TAR model), which would otherwise need to be approximated using many parameters in an ARMA model.
3.2 Alternating minimization algorithm
The optimization landscapes of (4) and (5) are complicated because the optimal transport cost is itself the minimum value of a linear program (3), and there are potentially many variables to optimize. This makes directly optimizing (4) or (5) computationally intractable. To remedy this issue, we propose to solve the nested minimization problem using alternating minimization. Since (4) is a special case of (5) with $\lambda=0$, we focus our discussion on (5).
To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (3) is exactly the discrete optimal transport problem. The minimizer $\hat{\mathbf{\Pi}}$ can be efficiently computed using existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019). Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (3), which corresponds to the inner optimization problem of (5). Next, we optimize (5) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed. This procedure is repeated until convergence. Let
$F(\mathbf{w},\mathbf{\Pi})=\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))+\frac{\lambda}{2}\|\mathbf{w}\|^{2},$ (6)
and the procedure described above can be summarized in Algorithm 1. One advantage of the alternating minimization scheme is that the subproblems (steps (a) and (b) in Algorithm 1) are easy to optimize when $L$ satisfies some regularity conditions. Optimization guarantees are discussed in the next subsection.
Input: Cut-off $n_{1}$, number of lags $p$
Initialization: $\mathbf{w}^{(0)}\in\mathcal{C}$
1 for $k=1,2,\ldots$ do
2      Step (a): Solve for $\hat{\mathbf{\Pi}}^{(k)}$ in problem (3), treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed.
3      Step (b): Solve for $\hat{\mathbf{w}}^{(k)}\in\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}^{(k)})$.
4      Break if convergence criterion is met and set $\hat{\mathbf{w}}=\hat{\mathbf{w}}^{(k)}$
5      end for
     Output: Imputed time series $\hat{\mathbf{w}}$
Algorithm 1 Temporal Wasserstein imputation with alternating minimization
Next, we investigate the properties of Algorithm 1 in an important special case where closed forms can be obtained. Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$. Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where
$\mathbf{H}(\mathbf{\Pi})$ $=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n},\mbox{ and}$
$\mathbf{A}_{h}(\mathbf{\Pi})$ $=\begin{pmatrix}\mathbf{0}_{p-1-h}&&&\\ &\frac{1}{n_{1}-p+2}\mathbf{I}_{n_{1}-p+2}&-\mathbf{\Pi}&\\ &-\mathbf{\Pi}^{\top}&\frac{1}{n-n_{1}-1}\mathbf{I}_{n-n_{1}-1}&\\ &&&\mathbf{0}_{h}\end{pmatrix}\in\mathbb{R}^{n\times n},\quad h=0,1,\ldots,p-1,$
in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero, and $\mathbf{0}_{h}$ is an $h\times h$ matrix of zeros. Hence $F$ is quadratic in $\mathbf{w}$. In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints. Moreover, $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 4 in Section 4). Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004; Parikh et al.,, 2014). If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e.,
$\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$
for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
$\hat{\mathbf{w}}^{(k)}=\mathbf{H}^{-1}\mathbf{K}^{\top}(\mathbf{K}\mathbf{H}^{-1}\mathbf{K}^{\top})^{-1}\mathbf{b},$ (7)
where $\mathbf{H}=\mathbf{H}(\hat{\mathbf{\Pi}}^{(k)})$.
Suppose further that $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$. That is, there is no side information about the missing values. In this case, the solution $(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ to the subproblem in step (b) must satisfy the following system of linear equations which has the same number of unknowns and equations.
$\hat{w}_{s}=&\frac{\sum_{i=s\vee(p-1)}^{s+p-1}\sum_{n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}}{\frac{p\wedge(s+1)}{n_{1}-p+2}+\frac{\lambda}{2}},\quad\mbox{if }0\leq s\leq n_{1}-p+1,s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=s}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}+\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{s+p-1}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{n_{1}-s+1}{n_{1}-p+2}+\frac{s+p-n_{1}+1}{n-n_{1}-1}+\frac{\lambda}{2}},\\ &\hskip 142.26378pt\mbox{if }n_{1}-p+2\leq s\leq n_{1},s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=p-1}^{n_{1}}\sum_{j=s}^{(s+p-1)\wedge(n-1)}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{p\wedge(n-s)}{n-n_{1}-1}+\frac{\lambda}{2}},\quad\mbox{if }n_{1}+1\leq s\leq n-1,s\in\mathcal{M}\\ \hat{w}_{s}=&x_{s},\quad\mbox{if }s\notin\mathcal{M}\end{aligned}\right.$ (8)
where $\hat{\mathbf{\Pi}}=(\hat{\pi}_{ij})$ is treated as fixed. We make a few remarks regarding (8).
Remark 1.
Let us contrast TWI with the optimal interpolator. (8) shows, in general, the temporal Wasserstein imputation is a linear function of many observed data $\{x_{t}:t\notin\mathcal{M}\}$, even when $p$ is small. On the other hand, the optimal interpolator is usually a linear combination of nearby observed data when the underlying process is a finite-order AR model. For instance, suppose the underlying time series is an AR(1) process: $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\{\epsilon_{t}\}$ is a mean-zero, unit-variance white noise. If $x_{n_{1}}=\bigstar$ is the only missing value in the time series, then the optimal interpolator (Peña and Maravall,, 1991) of $x_{n_{1}}$ is
$w_{n_{1},\mathrm{opt}}=\frac{\phi}{1+\phi^{2}}(x_{n_{1}-1}+x_{n_{1}+1}).$
In contrast, (8) shows that the temporal Wasserstein imputation (with $p=2$) is given by
$\hat{w}_{n_{1}}=\frac{1}{n_{1}^{-1}+(n-n_{1}-1)^{-1}}\left(\sum_{t=0}^{n_{1}-1}\hat{\pi}_{t+1,n_{1}+1}x_{t}+\sum_{t=n_{1}+1}^{n-1}\hat{\pi}_{n_{1},t}x_{t}\right),$
which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI shall use all data to learn the underlying dynamics.
Remark 2.
(EM interpretation) Let $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ be a fixed point of the alternating minimization algorithm (Algorithm 1). That is,
$\hat{\mathbf{w}}\in$ $\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),$
$\hat{\mathbf{\Pi}}\in$ $\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
In addition, suppose $\lambda=0$ for a moment. Let $\mathbf{u}=(u_{0},u_{1},\ldots,u_{p-1})$ be a random vector whose distribution is $\hat{\mu}_{\mathrm{pre}}$ and $\mathbf{v}=(v_{0},\ldots,v_{p-1})$ a random vector whose distribution is $\hat{\mu}_{\mathrm{post}}$. Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, which is $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures). Consider an index $s\in\mathcal{M}$ with $p-1\leq s\leq n_{1}-p+1$. By (8), it is not difficult to see that
$\hat{w}_{s}=$ $\frac{1}{p}\sum_{i=s}^{s+p-1}\sum_{j=n_{1}+1}^{n-1}\frac{\hat{\pi}_{ij}}{\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}}\hat{w}_{j-i+s}$
$=$ $\frac{1}{p}\sum_{i=s}^{s+p-1}\sum_{j=n_{1}+1}^{n-1}\mathbb{P}_{\hat{\mathbf{\Pi}}}\left(\mathbf{v}=(\hat{w}_{j},\ldots,\hat{w}_{j-p+1})|\mathbf{u}=(\hat{w}_{i},\ldots,\hat{w}_{i-p+1})\right)\hat{w}_{j-i+s}$
$=$ $\frac{1}{p}\sum_{h=0}^{p-1}\mathbb{E}_{\hat{\mathbf{\Pi}}}(v_{h}|\mathbf{u}=(\hat{w}_{s+h},\ldots,\underbrace{\hat{w}_{s}}_{\mbox{\makebox[0.0pt]{$h$-th element}}},\ldots,\hat{w}_{s+h-p+1})).$ (9)
(9) shows that the temporal Wasserstein imputation $\hat{w}_{s}$ is essentially an average of conditional expectations. Thus, Algorithm 1 is closely related to the EM algorithm: To minimize the Wasserstein distance $\mathbb{E}_{\mathbf{\Pi}}\|\mathbf{u}-\mathbf{v}\|^{2}$, step (b) computes the conditional expectations under the current optimal transport coupling $\hat{\mathbf{\Pi}}$ while step (a) minimizes over the coupling space. Formula like (9) can be derived for other time indices $s$ with appropriate modifications.
Remark 3.
(Barycenter interpretation) The summands in (9) is also known as the barycentric projections (Ambrosio et al.,, 2006; see also Deb et al.,, 2021). Hence the temporal Wasserstein imputation can also be interpreted as an average of barycentric projections under the coupling $\hat{\mathbf{\Pi}}$.
Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series. Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top},t=0,1,\ldots,n-1$. To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (3) by
$\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}.$
Then Algorithm 1 can still be applied. Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 is identical as the scalar case. In the notable special case where $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{2}$, (6) becomes
$F(\mathbf{W},\mathbf{\Pi})=\sum_{l=1}^{d}\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\left[\sum_{h=0}^{p-1}(w_{i-h,l}-w_{j-h,l})^{2}\right]+\frac{\lambda}{2}\|\mathbf{W}\|_{F}^{2},$
where $\|\cdot\|_{F}$ is the Frobenius norm. Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (8) in the case of no side information. Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel, and the imputation only depends on the future and past values of individual series. The cross-sectional information is only exploited when learning the optimal transport coupling in step (a). Thus, the algorithm is fast to implement in the multivariate setting.
We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by existing imputation methods. Their method corrects the resulting statistical pattern by penalizing the deviation of some moments from the theoretical values. For instance, let $\rho_{k}$, $k=1,2,\ldots,k_{0}$ be some pre-specified values for the autocorrelation function at lag $k$, and $m_{k}$, $k=1,2,\ldots,k_{1}$, the pre-specified values for the $k$-th moments. Then imputation is done by solving the following program.
$\min_{\mathbf{w}\in\mathcal{C}}\frac{f(\mathbf{w})}{f(\mathbf{w}^{(0)})}+\sum_{k=1}^{k_{0}}\lambda_{k}\left(\frac{\rho_{k}(\mathbf{w})}{\rho_{k}}-1\right)^{2}+\sum_{k=1}^{k_{1}}\mu_{k}\left(\frac{n^{-1}\sum_{t=0}^{n-1}w_{t}^{k}}{m_{k}}-1\right)^{2},$
where $\rho_{k}(\mathbf{w})$ is the sample autocorrelation function computed from the series $\mathbf{w}$, $f(\mathbf{w})=\sum_{t=1}^{n-1}(w_{t}-w_{t-1})^{2}$, and $\mu_{k},\lambda_{k}\geq 0$ are the penalty strengths. As a result, the moments and the autocorrelation functions of the imputed series will be close to the pre-specified values. However, this approach has a number of drawbacks. First, a good theoretical value for $\rho_{k}$ and $m_{k}$ may not be available in practice, especially when the missing values are many. Second, minimizing $f$ promotes smoothness and a positive correlation between $w_{t}$ and $w_{t-1}$, which does not necessarily align with the underlying time series. Finally, the optimization problem is highly non-convex, leading to intense computational hurdle. In contrast, the temporal Wasserstein imputation is data-driven and completely nonparametric, requiring small amount of tuning. It also enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
In closing this subsection, we discuss the imputation of nonstationary time series. For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and we can apply imputation methods to the first-differenced series $\{y_{t}\}$. In this case, the number of missing data multiplies in the first-differenced series and could compromise the imputation performance. However, for TWI we can mitigate this issue by encoding the information from the raw data via the admissible set $\mathcal{C}$. For example, if $x_{t}$ is observed, then $y_{1}+y_{2}+\ldots+y_{t}=x_{t}-x_{0}$, so we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed). Finally, although we have motivated TWI from the stationarity of the underlying process, we found that it fares quite well with certain nonstationary processes such as cyclic time series. In Section 5, we further explore the application of TWI to nonstationary time series via simulation studies.
4 Theoretical properties
In this section, we investigate the optimization and statistical properties of TWI. Under some general assumptions, we first show Algorithm 1 converges to critical points. Then, the asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$. For general missing patterns, we show that identification of the underlying marginal distributions is possible under the assumptions of asymmetric missing pattern and interpolator stability (see (C1) and (C2) in Section 4.2) using a two-state Markov chain example.
In the following, we maintain the assumption that the index set $\mathcal{M}=\mathcal{M}_{n}$, which corresponds to the missing data, is fixed and independent of the underlying time series. It is worth noting that we allow the number of missing entries to scale with the sample size, i.e., $\sharp(\mathcal{M}_{n})\rightarrow\infty$ as $n\rightarrow\infty$, where $\sharp(E)$ is the cardinality of a set $E$. This asymptotic regime is more suitable for applications where non-negligible portions of the data are missing. In addition, our focus is on the distributional bias induced by the imputations. If $\sharp(\mathcal{M}_{n})=o(n)$, the problem becomes trivial because the empirical marginal distributions of any (bounded) imputations satisfy
$\frac{1}{n}\sum_{t\in\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}+\frac{1}{n}\sum_{t\notin\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}\Rightarrow\mu_{p},$
where $\mathcal{A}_{n}=\{t:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}_{n}\neq\emptyset,t\in\{0,1,\ldots,n-1\}\}$.
4.1 Optimization and statistical guarantees
We begin by studying the optimization properties of Algorithm 1.
Proposition 4.
Suppose $\mathcal{C}$ is a convex set. Then,
  • (a)
    (monotonicity) $F(\mathbf{w}^{(t)},\mathbf{\Pi}^{(t+1)})\leq F(\mathbf{w}^{(t)},\mathbf{\Pi}^{(t)})\leq F(\mathbf{w}^{(t-1)},\mathbf{\Pi}^{(t)})$, $t=1,2,\ldots$;
  • (b)
    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\mathbf{w}^{(t+1)},\mathbf{\Pi}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. That is,
    $\frac{\partial F(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})}{\partial\mathbf{w}}\cdot(\mathbf{w}-\hat{\mathbf{w}})\geq 0,\quad\frac{\partial F(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})}{\partial\mathrm{vec}(\mathbf{\Pi})}\cdot\mathrm{vec}(\mathbf{\Pi}-\hat{\mathbf{\Pi}})\geq 0$
    for any $\mathbf{w}\in\mathcal{C}$ and $\mathbf{\Pi}$ satisfying the coupling constraints.
If, in addition, the cost function $L$ takes the form $L(\mathbf{u},\mathbf{v})=\sum_{k=1}^{p}\ell(u_{k}-v_{k})$, where $\mathbf{u}=(u_{1},\ldots,u_{p})^{\top}$, $\mathbf{v}=(v_{1},\ldots,v_{p})^{\top}$, and $\ell$ is twice differentiable and convex, then
  • (c)
    $F(\mathbf{w},\mathbf{\Pi})$ is biconvex in $\mathbf{w}$ and $\mathbf{\Pi}$.
Proof.
Clearly, (a) holds by construction of the algorithm. Since Algorithm 1 is an instance of the Gauss-Siedel algorithm, (b) follows from Grippo and Sciandrone, (2000). Thus it remains to prove (c). Optimizing $\mathbf{\Pi}$ while holding $\mathbf{w}$ fixed is the optimal transport problem, which is clearly convex. Now fix $\mathbf{\Pi}$. Note that
$F(\mathbf{w},\mathbf{\Pi})=$ $\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\sum_{h=0}^{p-1}\ell(w_{i-h}-w_{j-h})+\frac{\lambda}{2}\|\mathbf{w}\|^{2}$
$=$ $\sum_{h=0}^{p-1}\underbrace{\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell(w_{i-h}-w_{j-h})}_{F_{h}(\mathbf{w},\mathbf{\Pi})}+\frac{\lambda}{2}\|\mathbf{w}\|^{2}.$
It suffices to show each $F_{h}$ is convex in $\mathbf{w}$. In the following, we only prove the case of $F_{0}$ since the arguments for the other cases are similar. Let $\mathbf{u}=(w_{p-1},w_{1},\ldots,w_{n_{1}})^{\top}$ and $\mathbf{v}=(w_{n_{1}+1},\ldots w_{n-1})^{\top}$. Then
$\frac{\partial F_{0}}{\partial\mathbf{u}}$ $=\left(\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell^{\prime}(w_{i}-w_{j}):i=p-1,\ldots,n_{1}\right)^{\top},$
$\frac{\partial F_{0}}{\partial\mathbf{v}}$ $=\left(-\sum_{i=p-1}^{n_{1}}\pi_{ij}\ell^{\prime}(w_{i}-w_{j}):j=n_{1}+1,\ldots,n-1\right)^{\top}.$
Moreover,
$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}^{2}}=$ $\mathrm{diag}\left(\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):i=p-1,\ldots,n_{1}\right),$ (10)
$\frac{\partial^{2}F_{0}}{\partial\mathbf{v}^{2}}=$ $\mathrm{diag}\left(\sum_{i=p-1}^{n_{1}}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):j=n_{1}+1,\ldots,n-1\right),$ (11)
$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}\partial\mathbf{v}}=$ $-\left(\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j})\right)_{p-1\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}.$ (12)
Since $\ell$ is convex, ′′$\ell^{\prime\prime}\geq 0$. From (10)–(12), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries. It follows from the Gershgorin circle theorem that $\mathbf{H}$ is positive semi-definite. Thus $F_{0}$ is convex in $\mathbf{w}$. ∎
Remark 5.
It is straightforward to generalize Proposition 4 for multivariate time series. Details are omitted.
Due to the biconvexity of $F$, one can employ many off-the-shelf solvers for the subproblems in Algorithm 1. Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$. Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
In general, not all critical points of $F$ yield desirable imputations, and even its global minimizer may not correspond to the correct marginal distributions. For (an extreme) example, imputing a fixed constant when all data are missing results in a zero transport cost, which is clearly arbitrary. However, when the missing data happen to lie on the same side of $n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 6 below. Before stating the proposition, we make the following assumptions, which are satisfied by many linear and nonlinear time series models.
The underlying time series $\{\tilde{x}_{t}\}$ is stationary and ergodic.
$\mathbb{E}|\tilde{x}_{t}|^{k}<\infty$ for some $k\geq 1$.
Proposition 6.
Assume (A1) and (A2) hold. Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$. Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$. If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$, subject to relevant constraints such as the coupling conditions and with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, satisfies
$\mathcal{W}_{k}\left(\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)+\mathcal{W}_{k}\left(\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)\rightarrow 0\quad\mbox{almost surely},$
where $\mu_{p}$ is the $p$-dimensional marginal distribution of $\{\tilde{x}_{t}\}$.
Proposition 6 implies the empirical marginal distributions of the imputed time series converge to the underlying marginal distribution in the Wasserstein distance. This implies
$\hat{\mu}_{\mathrm{pre}}(\hat{\mathbf{w}})=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}\Rightarrow\mu_{p},\mbox{ and}\quad\hat{\mu}_{\mathrm{post}}(\hat{\mathbf{w}})=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}\Rightarrow\mu_{p}.$
In fact, convergence in Wasserstein distance further ensures for any continuous function $\varphi$ with $|\varphi(x)|\leq C(1+|x|^{k})$ for all $x\in\mathbb{R}$ and some $C<\infty$, we have
[…]
In practice, we may not have all missing data lying on the same side of $n_{1}$, even though we have some freedom to choose $n_{1}$. Unfortunately, asymptotic consistency results for general missing pattern that are in similar spirits as Proposition 6 can hardly be derived. The reason, as we shall see in the next subsection, is that matching the underlying marginal distribution by equating the empirical marginal distributions essentially suffers from an identification problem.
4.2 General missing pattern: A two-state Markov chain example
In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example. Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations. We also show that, under two favorable conditions, the correct marginal distributions can still be pinned down by TWI. To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
Suppose the underlying time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain. Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by
$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.2,$
$\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.6.$
The transition probabilities are specified to facilitate discussions; the main results below are not specific to these choices. It is easy to check that $\{\tilde{x}_{t}\}$ has the following stationary distribution.
$\mathbb{P}(\tilde{x}_{t}=1)=0.75,\quad\mathbb{P}(\tilde{x}_{t}=0)=0.25.$
Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=0.6,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=0.15,$ (13)
$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=0.15,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=0.1.$
Without loss of generality, we set $n_{1}=0$. Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations. Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$. Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$, $q_{2}$ are some integers larger than 2. Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ (14)
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$
and if $t\geq n_{1}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ (15)
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$
for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$. Note that $\mathbb{P}$ in (14) and (15) should be interpreted as proportions since the imputation may still depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined. Imputation methods corresponding to varying $a_{1},a_{2},b_{1},b_{2}$ will lead to different 2-dimensional marginal distributions. Indeed, from (13)–(15), we can compute the 2-dimensional marginal distributions of the imputed series. For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-q_{1}^{-1})+\frac{3}{4q_{1}}a_{1},$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))$ $=0.15(1-q_{1}^{-1})+\frac{1}{4q_{1}}b_{1},$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))$ $=0.15(1-q_{1}^{-1})+\frac{3}{4q_{1}}(1-a_{1}),$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))$ $=0.1(1-q_{1}^{-1})+\frac{1}{4q_{1}}(1-b_{1}).$
For $\{\hat{w}_{t}\}_{0\leq t<\infty}$, we can obtain similar expressions with $a_{1},b_{1},q_{1}$ replaced by $a_{2},b_{2},q_{2}$. The “correct” imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting 2-dimensional marginal distributions equal the marginal distribution of the underlying series, given in (13). To achieve such identification, the correct imputation must choose
$a_{1}^{*}=a_{2}^{*}=0.8,\quad b_{1}^{*}=b_{2}^{*}=0.6.$ (16)
Note that this calibration does not depend on $q_{1},q_{2}$; it only depends on the distributional properties of the underlying process.
Assuming the optimization is carried out perfectly [^1] This is where we ignore finite sample complications and assume a zero Wasserstein loss is achieved. One may use, for example, the hamming distance $L(x,y)=\mathbf{1}\{x\neq y\}$ as the ground metric for the discrete data., the temporal Wasserstein imputation (with $p\geq 2$) calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$. Hence, $\{a_{1},a_{2},b_{1},b_{2}\}$ satisfies
[…]
Clearly, if $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ satisfies the above system of equations. Therefore, identification is not possible in this case, which leads to our first condition.
Asymmetric missing pattern: $q_{1}\neq q_{2}$.
However, condition (C1) is still inadequate for identification because the equations in (17) are not linearly independent. The lack of uniqueness of the solutions to (17) is at the crux of the identification issue of the imputation problem. The blue lines in Figure 1 plots the set of solutions to (17). Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same.
Figure 1: Set of calibrations that correspond to temporal Wasserstein imputation solutions (blue line) and the correct calibrations $(a_{1}^{*},a_{2}^{*})$ and $(b_{1}^{*},b_{2}^{*})$ (red dots). The dashed orange line is the 45 degree line. $q_{1}=3$ and $q_{2}=4$ are used to plot this figure.
Nonetheless, (16) hints an additional requirement for the correct imputation, giving rise to the following condition.
Interpolator stability: $a_{1}=a_{2}$, $b_{1}=b_{2}$.
Intuitively, (C2) requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$. Together, (C1) and (C2) imply a unique solution to the system (17) and are sufficient for identification.
Proposition 7.
Under (C1) and (C2), the only solution to (17) is
$a_{1}=a_{2}=a_{1}^{*},\quad b_{1}=b_{2}=b_{1}^{*}.$
That is, the only marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the set of temporal Wasserstein imputations at the correct calibrations.
The above example illustrates when TWI can potentially identify the underlying marginal distribution. The crucial interpolator stability condition (C2) thus provides some guidance on the selection of TWI solutions. Indeed, Algorithm 1 may converge to varying solutions depending on the initialization. (C2) implies that we shall choose the solution so that the relationship between the imputed values and the neighboring observations stays stable before and after $n_{1}$. In higher dimensions, summary statistics can be used. For example, fit a linear regression of imputed values on neighboring values using data before and after $n_{1}$ and compare the similarity of the coefficients. To promote interpolator stability, we suggest running TWI with different cut-off $n_{1}$, using the previously obtained imputations as initializations. This leads to a simple procedure termed $k$-TWI henceforth, which is summarized in Algorithm 2 below.
Input: Cut-off points $n_{1},n_{2},\ldots,n_{k}$, number of lags $p$
Initialization: $\hat{\mathbf{w}}^{(0)}\in\mathcal{C}$
1 for $h=1,2,\ldots,k$ do
2      Run Algorithm 1 with initialization $\hat{\mathbf{w}}^{(0)}$ and cut-off $n_{h}$ to obtain $\hat{\mathbf{w}}$
3      Set $\hat{\mathbf{w}}^{(0)}=\hat{\mathbf{w}}$
4      end for
     Output: Imputed time series $\hat{\mathbf{w}}$
Algorithm 2 $k$-TWI
5 Simulation Studies
In this section, we apply the proposed temporal Wasserstein imputation to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance. For comparison, we employ alternative imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations. In addition, the following two missing patterns are considered.
  • Pattern I: 300 observations are omitted at random, which corresponds to sporadic missing data.
  • Pattern II: For every 20 observations, 6 consecutive observations are omitted, which mimics patches of missing data.
The total length of the time series is set to $n=1000$. Thus, for both missing patterns, 30% of the data are missing. For each model and missing pattern considered in this section, 1000 Monte Carlo simulations are carried out to compute the measure of performance.
We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward. Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (7). In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($0.25n$, $0.5n$, and $0.75n$) to the simulated data. Both linear interpolation and Kalman smoothing are used as initialization of TWI in our experiments. Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R. The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series. The details are documented in Appendix B.
5.1 Linear and nonlinear univariate time series
In this subsection, we focus on univariate time series exhibiting linear or nonlinear dynamic patterns. Specifically, we consider the following time series models. In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d. sequences of standard Gaussian variables.
Model 1 (AR).
$x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\phi$ is set to 0.8.
Model 2 (ARMA).
$(1-\phi_{1}B)x_{t}=(1+\phi_{2}B)\epsilon_{t}$, where $B$ is the backshift operator and $(\phi_{1},\phi_{2})=(0.8,-0.6)$.
Model 3 (TAR; Tong,, 1983).
$x_{t}=\left\{\begin{array}[]{cc}\phi_{1}x_{t-1}+\epsilon_{t},&x_{t-1}\leq\tau\\ \phi_{2}x_{t-1}+0.5\epsilon_{t},&x_{t-1}>\tau\end{array}\right.,$
where $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$
Model 4 (I(1) process).
Data $\{x_{t}\}$ are generated according to
$z_{t}=$ $\phi z_{t-1}+0.5\epsilon_{t,1},$
$x_{t}=$ $x_{t-1}+z_{t}+\epsilon_{t,2},$
with $\phi=-0.7$.
Model 5 (Deterministic cyclic series with noise (CYC)).
$x_{t}=10\cos(t(0.23\pi))+6\cos(t(0.17\pi))+0.5\epsilon_{t}.$
Models 1 and 2 are linear time series models commonly seen in practice. Kalman smoothing is particularly adept at imputing such series. Note that Model 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution, the key objects TWI and $k$-TWI work with, is inadequate to fully characterize the distribution of the underlying process. Model 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018). Model 4 generates an I(1) series which typically produces a highly persistent stochastic trend. To remove the unit root, difference transform is usually applied to the data $\{x_{t}\}$; that is, one often deals with $\nabla x_{t}=x_{t}-x_{t-1}$ instead of the original series. However, this may result in excessive missing values in $\{\nabla x_{t}\}$, which can cause difficulty in imputation. Finally, Model 5 features a deterministic cyclic trend with non-standard periods. Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
Three measures of performance are considered. First, we report the Wasserstein distance (of order 2) between the empirical 3-dimensional marginal distributions of the imputed series and the full data. Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{x_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$. We report
$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{({x}_{t}^{(i)},{x}_{t-1}^{(i)},{x}_{t-2}^{(i)})}\right),$ (18)
as a measure of how well the imputed series approximates the original series. Second, we compare the root mean square estimation errors (RMSE) of the estimated autocovariance functions. Define $\gamma(\mathbf{z},h)=(n-h)^{-1}\sum_{t=h}^{n-1}(z_{t}-\bar{z})(z_{t-h}-\bar{z})$, where $\mathbf{z}=(z_{0},z_{1},\ldots,z_{n-1})$ and $\bar{z}=n^{-1}\sum_{t=0}^{n-1}z_{t}$. Then the RMSE is defined as
$\mathrm{RMSE}_{\gamma}(h)=\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}(\gamma(\hat{\mathbf{w}}^{(i)},h)-\gamma_{\mathbf{x}}(h))^{2}},$ (19)
where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(x_{0}^{(i)},\ldots,x_{n-1}^{(i)})$. Third, for Models 14, the model parameters are estimated using the imputed series, and their root mean square errors are reported.
The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1. Across all five models employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series. While other methods may yield good imputations for some specific models, the result is not consistent across all models. For instance, linear and spline interpolations show low Wasserstein loss for Model 1, but their performance is less competitive when applied to other models. Similarly, the Kalman smoothing performs well under Model 1 with missing pattern I, but its advantage diminishes under missing pattern II. In contrast, TWI methods consistently deliver solid performance across models and missing patterns, provided they are carefully initialized.
Additionally, Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often siginificantly. For example, for Model 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74). Similarly, for Model 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79). Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed (${x}_{t}=\bigstar$ or ${x}_{t-1}=\bigstar$) under Model 3, alongside $(x_{t},x_{t-1})$ of the original data. While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth. Figure 3 similarly plots $\hat{w}_{t}$ against $\hat{w}_{t-h}$ (and $x_{t}$ against $x_{t-h}$) for $h=1,2,6,10$ under Model 5. As shown in the figure, TWI markedly improves from its linear initialization by “denoising” the marginal distributions, leading to a distribution that closely match the ground truth.
For the nonstationary Model 4 (I(1)), we apply the methods to obtain an imputation $\{\hat{w}_{t}\}$ and focus on the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$. Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$. As noted before, there are many more missing values in $\{\nabla x_{t}\}$ than originally in $\{x_{t}\}$. In this case, TWI still consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations. TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization. Similar notations are used for $k$-TWI. For Model 4 (I(1)), the loss is computed using the first-differenced series. The smallest value for each model and missing pattern is marked in boldface.
Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
Missing pattern I
AR 0.41 0.41 0.42 0.50 0.40 0.44 0.41 0.44
ARMA 0.48 0.58 0.47 0.48 0.40 0.38 0.40 0.38
TAR 1.12 1.51 1.38 1.13 0.96 0.81 0.74 0.63
I(1) 0.71 0.75 0.62 4.38 0.58 0.53 0.54 0.50
CYC 1.96 0.82 0.77 1.97 0.79 0.77 0.60 0.70
NLVAR 2.98 3.42 3.01 2.94 2.25 2.14 2.19 2.12
ALR (×$\times 10$) 0.70 1.00 4.17 0.55 0.43 0.37 0.38 0.35
Missing pattern II
AR 0.44 0.59 0.50 0.76 0.39 0.44 0.43 0.45
ARMA 0.47 1.29 0.51 0.53 0.36 0.34 0.40 0.35
TAR 1.04 3.11 1.25 0.99 0.84 0.73 0.76 0.61
I(1) 0.67 0.65 0.60 8.35 0.51 0.45 0.49 0.44
CYC 2.58 1.81 0.77 3.79 2.62 1.60 0.62 0.72
NLVAR 3.00 6.07 3.33 3.67 2.13 2.02 2.31 2.04
ALR (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
(a) Marginal distribution under missing pattern I
(b) Marginal distribution under missing pattern II
Figure 2: Scatter plot of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
(a) $h=1$
(b) $h=2$
(c) $h=6$
(d) $h=10$
Figure 3: Scatter plots of $(x_{t-h},x_{t})$ of the original data and $(\hat{w}_{t-h},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 5 (CYC) under missing pattern I. $h\in\{1,2,6,10\}$.
Next, we examine the sample autocovariance functions (ACFs) estimated from the imputed data. Table 2 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$. For Model 1 (AR), most methods perform reasonably well. Linear interpolation produces accurate ACF estimates since the positive AR coefficient intrinsically favors linearity, while the scalar filter exhibits largest estimation errors. For Model 2 (ARMA), TWI improves from linear and Kalman initializations and yields more accurate ACF estimates: TWI reduces the biases in Kalman smoothing when $h=0$ and in linear interpolation when $h=1$, and $h=2$. For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series exhibit the lowest RMSEs, with substantial improvements over the initializations. Notably, the scalar filter also yields accurate ACFs for $h=1,2$. Figure 4 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
The advantage of TWI becomes even more pronounced in nonstationary models 4 (I(1)) and 5 (CYC), for which it achieves low ACF estimation errors. The scalar filter performs poorly for these models, which is likely the consequence of using many indicators in the intervention analysis. Among the benchmarks, Kalman filter produces the best result for these two models, but it can be further improved by the proposed TWI procedure.
(a) missing pattern I
(b) missing pattern II
Figure 4: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR). The blue line is the smoothed density estimated from the histogram of the ACFs estimated from the ground truth.
Table 2: $\mathrm{RMSE}_{\gamma}(h)$, defined in (19), for $h=0,1,2$, with $h$ denoting lag. The results are averaged over 1000 simulations, and the smallest values in each row are marked in boldface. See Table 1 for notations.
Model $h$ Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
Missing pattern I
AR 0 0.32 0.29 0.37 0.50 0.30 0.31 0.32 0.32
1 0.26 0.34 0.29 0.47 0.27 0.30 0.28 0.31
2 0.26 0.27 0.29 0.38 0.25 0.28 0.27 0.29
ARMA 0 0.14 0.19 0.29 0.31 0.13 0.13 0.22 0.18
1 0.22 0.44 0.07 0.08 0.14 0.10 0.07 0.07
2 0.08 0.08 0.07 0.07 0.06 0.06 0.07 0.07
TAR 0 0.59 0.77 0.73 1.07 0.65 0.56 0.56 0.45
1 0.81 1.68 0.19 0.14 0.46 0.32 0.17 0.13
2 0.40 0.31 0.16 0.11 0.24 0.16 0.10 0.09
I(1) 0 0.52 0.40 0.48 19.41 0.42 0.35 0.40 0.34
1 0.35 0.51 0.25 9.32 0.28 0.25 0.22 0.21
2 0.17 0.35 0.11 1.28 0.15 0.14 0.11 0.11
CYC 0 10.94 2.40 3.90 4.67 1.81 1.25 0.68 0.62
1 9.05 1.72 3.33 5.88 1.47 1.04 0.54 0.51
2 1.90 0.46 0.81 2.13 0.41 0.33 0.17 0.18
Missing pattern II
AR 0 0.38 0.70 0.55 0.78 0.33 0.36 0.43 0.39
1 0.28 0.79 0.42 0.70 0.29 0.34 0.37 0.36
2 0.26 0.66 0.36 0.65 0.27 0.31 0.34 0.34
ARMA 0 0.14 1.35 0.31 0.32 0.13 0.11 0.24 0.17
1 0.17 1.49 0.08 0.09 0.10 0.08 0.07 0.07
2 0.15 1.05 0.07 0.08 0.09 0.07 0.07 0.07
TAR 0 0.49 5.39 0.71 1.08 0.72 0.68 0.73 0.65
1 0.75 6.04 0.34 0.11 0.31 0.20 0.09 0.09
2 0.77 4.36 0.34 0.10 0.42 0.28 0.19 0.14
I(1) 0 0.47 0.27 0.47 51.84 0.38 0.30 0.36 0.29
1 0.19 0.43 0.18 0.56 0.17 0.14 0.14 0.13
2 0.08 0.08 0.07 0.52 0.08 0.08 0.07 0.07
CYC 0 7.46 7.56 1.96 15.21 12.27 6.49 1.26 1.03
1 2.63 3.99 1.45 16.23 9.57 5.07 0.98 0.83
2 11.18 5.09 0.79 7.90 1.28 1.04 0.28 0.31
Finally, we turn to the model parameters estimated from the imputed series. Table 3 presents the RMSEs for each model parameter. For Model 1 (AR) and Model 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements. For the nonlinear Model 3 (TAR), the results are drastically different. TWI produces the most accurate parameter estimates and can achieve approximately 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing. In some cases, such as with linear initialization under missing pattern II, the improvement far exceeds 50%. These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics. For Model 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization. Figure 5 shows the histogram of the estimated parameter $\phi_{1}$ in Model 4. It is visible from the figure that TWI helps neutralize the biases from the initialization and produces reliable estimates.
Table 3: Estimation root mean square errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
Missing pattern I
AR $\phi$ 0.06 0.06 0.05 0.03 0.03 0.04 0.03 0.04
ARMA $\phi_{1}$ 0.19 0.47 0.12 0.08 0.17 0.16 0.11 0.11
$\phi_{2}$ 0.46 0.95 0.18 0.10 0.33 0.26 0.16 0.15
TAR $\phi_{1}$ 1.02 1.67 0.82 0.63 0.71 0.58 0.53 0.44
$\phi_{2}$ 0.08 0.07 0.20 0.04 0.05 0.03 0.03 0.02
$\tau$ 0.01 0.75 0.35 0.20 0.05 0.05 0.03 0.03
I(1) $\phi$ 0.96 0.47 0.45 0.68 0.48 0.32 0.30 0.24
NLVAR $\phi_{11}$ 0.19 0.27 0.15 0.03 0.06 0.03 0.05 0.03
$\phi_{12}$ 3.10 3.16 2.87 1.67 1.34 1.02 1.22 0.97
$\phi_{21}$ 0.03 0.07 0.03 0.03 0.03 0.03 0.03 0.03
$\phi_{21}$ 0.18 0.23 0.10 0.04 0.05 0.04 0.04 0.04
Missing pattern II
AR $\phi$ 0.05 0.07 0.04 0.04 0.03 0.02 0.03 0.02
ARMA $\phi_{1}$ 0.04 0.08 0.09 0.08 0.05 0.06 0.09 0.09
$\phi_{2}$ 0.16 0.61 0.11 0.10 0.12 0.11 0.11 0.12
TAR $\phi_{1}$ 0.72 2.44 0.38 0.25 0.25 0.20 0.29 0.21
$\phi_{2}$ 0.09 0.12 0.05 0.03 0.06 0.05 0.02 0.01
$\tau$ 0.01 1.72 0.01 0.01 0.01 0.01 0.01 0.01
I(1) $\phi$ 0.08 1.08 0.11 0.70 0.09 0.10 0.10 0.10
NLVAR $\phi_{11}$ 0.19 0.44 0.08 0.02 0.02 0.01 0.02 0.01
$\phi_{12}$ 2.67 3.51 1.68 1.26 0.61 0.46 0.61 0.43
$\phi_{21}$ 0.02 0.09 0.03 0.03 0.06 0.05 0.04 0.03
$\phi_{21}$ 0.15 0.34 0.05 0.04 0.04 0.04 0.04 0.04
(a) missing pattern I
(b) missing pattern II
Figure 5: Histogram of estimated model parameter $\phi_{1}$ when the data are generated from Model 4 (I(1)). The blue line is the smoothed density estimated from the histogram of the parameter estimated from the ground truth.
In conclusion, under the univariate time series models considered in this subsection, which include linear, nonlinear, and nonstationary time series, TWI consistently fares well in all three comparing metrics. It can capture nonlinear dynamics as illustrated by the low Wasserstein error in the imputed marginal distributions, which provides substantial improvements in downstream ACF and model estimates.
5.2 Multivariate time series
In this subsection, we consider the following two multivariate nonlinear time series models.
Model 6 (Nonlinear VAR; NLVAR).
$x_{t,1}=$ $\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}$
$x_{t,2}=$ $\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
where $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ is the centered sigmoid function, and $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
Model 7 (Additive logistic model; AL, Brunsdon and Smith,, 1998).
$\begin{pmatrix}y_{t,1}\\ y_{t,2}\end{pmatrix}=\begin{pmatrix}0.1\\ 0.1\end{pmatrix}+\begin{pmatrix}0.7&-0.5\\ 0&-0.7\end{pmatrix}\begin{pmatrix}y_{t-1,1}\\ y_{t-1,2}\end{pmatrix}+0.2\begin{pmatrix}\epsilon_{t,1}\\ \epsilon_{t,2}\end{pmatrix},$
and the observed data are
$x_{t,j}=\left\{\begin{aligned} &\frac{\exp(y_{t,j})}{1+\exp(y_{t,1})+\exp(y_{t,2})},\quad j=1,2,\\ &\frac{1}{1+\exp(y_{t,1})+\exp(y_{t,2})},\quad j=3.\end{aligned}\right.$
Model 6 is a VAR model with a nonlinear term. In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function. Model 7 generates the so-called compositional time series, which refers to a multivariate time series $\{\mathbf{x}_{t}\}$, where $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$, $d>1$, such that for each $t$, $\sum_{j=1}^{d}x_{t,j}=1$, and $x_{t,j}\geq 0$ for all $j$. Such data are encountered in a wide array of fields such as the environmental sciences, economics, geology and political science. Here, we adopt the additive logistic (AL) model of Brunsdon and Smith, (1998) in Model 7. The AL model postulates that after some nonlinear transformation, the data follow a standard VARMA model. In fact, one can view compositional data as a time series on some low-dimensional manifold. Alternative models for composition time series include the recent works of Zheng and Chen, (2017), Harvey et al., (2024), and Zhu and Müller, (2024).
Focus on Model 6 (NLVAR) first. The Wasserstein distances between the 3-dimensional marginal distributions of the imputed and the original series, defined in (18), are again reported Table 1. TWI methods clearly produce the lowest Wasserstein distances. As depicted in Figure 6, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships. Next, we assess the effectiveness of the imputed series in estimating the model parameters, assuming the nonlinearity is known. That is, we estimate $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$ by plugging in the imputed series. The root mean square estimation errors are documented in Table 3. Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$. It attains more than 50% reduction in RMSE from initialization under both missing patterns I and II, and can achieve 50% lower RMSE than the scalar filter, the best-performing benchmark.
(a) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,1}$
(b) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,1}$
(c) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,2}$
(d) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,2}$
Figure 6: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 (NLVAR) under missing pattern II.
Next, we turn to Model 7 (AL). Since the components must sum to unity, they exhibit negative correlation when one component approaches the endpoints 0 and 1. However, such correlation tends to weaken when the series are near the midpoint. Table 1 records the Wasserstein distance between the 3-dimensional marginal distribution of the imputed and original series. With the smallest Wasserstein loss, TWI methods again demonstrate their ability to adapt to nonlinear dependence structures.
It is worth noting that while the linear interpolation results in a compositional series, other methods, such as the Kalman smoothing, do not generally maintain this compositional property. By restricting the admissible set, TWI ensures that the imputed series remain compositional. As shown in Figure 7, linear interpolation falsely promotes negative correlation between $\hat{w}_{t,3}$ and $\hat{w}_{t-1,2}$ and Kalman smoothing often deviates from the plausible range of a compositional series. TWI amends these peculiarities, leading to imputations that align more closely with the underlying data.
Figure 7: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data (plotted in blue) and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series (plotted in red) for $i,j=1,2,3$, $i\neq j$, when the data are generated from Model 7 (AL) under missing pattern I.
6 Groundwater data application
In this section, we apply the proposed temporal Wasserstein imputation to a groundwater dataset from Taiwan. Groundwater serves as a critical water source for Taiwan, and its management is of considerable importance to the policymakers as well as the industrial sector, particularly due to the island’s steep topography and varying rainfall distribution (Hsu et al.,, 2020). However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells. To address this, imputation of missing values is essential to facilitate reliable downstream analyses.
6.1 Preprocessing and exploratory analysis
Our goal is to study the seasonal groundwater variations, for which we first transform the data into monthly records. The data contain hourly observations from 352 monitoring sites between October, 1992 and August, 2020. For each site, we define monthly groundwater level by the average of the observations in a given month. If more than one-third of the observations for a month are missing, we classify that month as missing. For our analysis, we focus on sites with a less-than-median amount of missing data, reducing the sample to 176 sites. Figure 8 shows the monthly groundwater levels for these sites, after subtracting the mean and normalized by the standard deviation. We make two observations. First, many series exhibit strong seasonality, which is due to the rainfall pattern in Taiwan. There is more rain in the summer due to monsoon and typhoons, while winters are generally drier. Second, there are many outlying values and anomalies in the observed series. For instance, there are abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean. For the purpose of our analysis, we treat such extreme values as missing data.
Figure 8: Time plot of the groundwater levels of the 176 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites. Figure 9 shows the fluctuations in the percentage of sites with missing data over time. Although missing data are mostly prevalent in the early years with more than 50% of the sites generating missing data, there are some time during which the percentage surges, such as between 2003 and 2005. In addition, the percentage of missing data is gradually increasing in more recent years. Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
Figure 9: Monthly percentage of sites recording a missing value.
6.2 Results
Similar to Section 5, we employ linear interpolation, spline interpolation, Kalman smoothing, and the scalar filter as benchmark imputation methods. For TWI and $k$-TWI, linear interpolation and Kalman smoothing are again used as initializations.
We consider two downstream tasks using the imputed series. First, we perform dimension reduction via the time series factor model of Lam et al., (2011). The factor model postulates that the observations $\mathbf{x}_{t}\in\mathbb{R}^{176}$ are driven by a small number of latent factor processes. Specifically,
$\mathbf{x}_{t}=\mathbf{A}\mathbf{f}_{t}+\bm{\epsilon}_{t},$
where $\mathbf{f}_{t}\in\mathbb{R}^{r}$, $r\ll 176$, is the factor process and $\{\bm{\epsilon}_{t}\}$ is white noise. Factor modeling is helpful for understanding the high-dimensional time series and its complex temporal patterns. However, estimating the factor model requires eigenanalysis of the autocovariance matrices $\mathbf{\Sigma}(h)=\mathbb{E}(\mathbf{x}_{t}\mathbf{x}_{t-h})$, which becomes infeasible when the series are severely corrupted by missing data.
Figure 10 shows the leading two factors estimated using series imputed by the benchmark methods. The linear and spline interpolation methods, both favoring trend preservation, result in the first estimated factor displaying strong trends, while the second factor reflects more the seasonal patterns. Notably, this is reversed when factors are estimated based on series imputed by Kalman smoothing and scalar filter. In these cases, the leading factor process exhibits strong seasonality, and the second factor displays trend patterns with a slight U-shape, which is quite different from those estimated with linear and spline methods.
Figure 10: Factor processes estimated using series imputed by the benchmark methods. The dashed red line marks January 2002.
Figure 11 displays the leading two factors estimated using series imputed by TWI and $k$-TWI. Note that while linear interpolation has led to a strong trend factor, as shown in Figure 10, the leading estimated factor of TWI initialized by linear interpolations exhibits significant seasonal pattern, which aligns more closely with those estimated by Kalman smoothing and scalar filter. Moreover, the second estimated factor also shows trends resembling those of the Kalman smoothing and scalar filter, but with a more apparent U-shape. To compare TWI with Kalman smoothing, we focus on the data before January 2002 (marked by the red dashed lines in Figure 10 and 11). Recall that missing data are prevalent during this period (Figure 9). Factors estimated by Kalman smoothing are affected by the abundance of missing data and tend to show smaller variations in the early periods, which may be the result of small variances in the optimal interpolator. In contrast, factors estimated using data imputed by TWI (initialized by Kalman smoothing) show consistent variational patterns across time and appear more trustworthy in understanding the dynamics. Overall, TWI took a more holistic view of the series which accounts for the prominent seasonality and trend in the data.
Figure 11: Factor processes estimated using series imputed by TWI. The dashed red line marks January 2002.
Our second downstream task concerns autoregreesive modeling. Specifically, we fit an AR(6) model using data imputed by the methods considered. The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including missing values. Then, we assess the impact of imputation quality by the out-of-sample prediction performance in predicting the remaining data. Note that the missing data in the test data are removed for fair comparison. Figure 12 presents, in violin plots, the distribution of the root mean squared prediction errors (RMSE) for each method across the 176 series under study. The TWI-type methods effectively alleviate the large RMSEs caused by the benchmark imputation methods and produce lower RMSEs in general.
Figure 12: The root mean squared prediction errors (RMSEs) distribution for each method across the 176 series, ranked by median from left to right.
In conclusion, the temporal Wasserstein imputation preserves important temporal features such as the seasonal patterns and enables complex downstream statistical analyses such as factor modeling. Our analysis show that the results obtained from TWI appear reliable. Thus, we expect the series imputed by TWI are also suitable for further analyses, especially when nonlinear models are employed.
7 Conclusion
In this work, we proposed an imputation method, the temporal Wasserstein imputation, for imputing missing entries in time series data. The proposed method complements the existing approaches by effectively capturing nonlinear dynamics and reducing the distributional bias typically observed for the conventional methods in handling missing time series data. Hence, the imputation of the proposed method is more reliable in downstream analysis. Our simulation studies and the analysis of groundwater data support the practical utility of the proposed method.
Appendix A Proof of Proposition 6
Lemma 8.
Suppose $\{\tilde{x}_{t}\}$ satisfies (A1) and (A2). Then
$\mathcal{W}_{k}\left(\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})},\mu_{p}\right)\xrightarrow{n\rightarrow\infty}0\quad\mathrm{a.s.},$
where $\tilde{\mathbf{x}}=(\tilde{x}_{0},\ldots,\tilde{x}_{n-1})$.
Proof.
By Theorem 7.12 of Villani, (2021), it suffices to show that with probability one,
$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}\Rightarrow\mu_{p},$ (20)
and
$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|^{k}.$ (21)
By Theorem 36.4 of Billingsley, (1995), the processes $\{\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\}_{t}$ and $\{\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\}_{t}$ are stationary and ergodic for any $\mathbf{h}\in\mathbb{R}^{p}$, where $\mathrm{i}=\sqrt{-1}$. Thus (21) follows immediately from the Ergodic theorem. To show (20), it suffices to show, with probability one,
$\hat{\varphi}_{n}(\mathbf{h}):=\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\rightarrow\int\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v})d\mu_{p}(\mathbf{v})=:\varphi(\mathbf{h})\quad\mbox{for all }\mathbf{h}\in\mathbb{R}^{p}.$
By the Ergodic theorem, we have
$\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q}),\quad\mbox{for all }\mathbf{q}\in\mathbb{Q}^{p},$ (22)
with probability one. Let
=$\mathcal{E}=\{\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q})\mbox{ for all }\mathbf{q}\in\mathbb{Q}^{p}\}\cap\left\{\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|\right\},$
noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2). By the Ergodic theorem and (22), $\mathbb{P}(\mathcal{E})=1$. Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary. Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
$|\varphi(\mathbf{h})-\varphi(\mathbf{q})|<\epsilon,$ (23)
for all $\mathbf{q}\in\mathbb{Q}^{p}$ with $\|\mathbf{h}-\mathbf{q}\|<\delta$. In addition, using $|\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{z})-1|\leq\|\mathbf{h}\|\|\mathbf{z}\|$, we have
$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (24)
for all $\mathbf{h}$, $\mathbf{q}$ such that $\|\mathbf{h}-\mathbf{q}\|<\delta$. Combining (23)–(24), we have
$|\hat{\varphi}_{n}(\mathbf{h})-\varphi(\mathbf{h})|\leq$ $|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}(\mathbf{q})|+|\hat{\varphi}_{n}(\mathbf{q})-\varphi(\mathbf{q})|+|\varphi(\mathbf{q})-\varphi(\mathbf{h})|$
$\leq$ $\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|+o(1)+\epsilon$
on $\omega\in\mathcal{E}$. Note, in particular, that $\epsilon$ and $\delta$ do not depend on $\mathbf{h}$. Letting $\epsilon\rightarrow 0$ and $\delta\rightarrow 0$ proves (20) holds on $\mathcal{E}$, and hence with probability one. ∎
Proof of Proposition 6.
We only consider the case $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ here as the other case follows from a similar argument. Define
$\hat{\mu}_{\mathrm{pre}}$ $:=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}=:\mu_{\mathrm{pre}},$
$\hat{\mu}_{\mathrm{post}}$ $:=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}$
$\mu_{\mathrm{post}}$ $:=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}$
Since the Wasserstein distance of order $k$ is a metric on the space of probability measures with finite $k$-th moments, we have
$\mathcal{W}_{k}(\hat{\mu}_{\mathrm{post}},\mu_{p})\leq$ $\mathcal{W}_{k}(\hat{\mu}_{\mathrm{post}},\hat{\mu}_{\mathrm{pre}})+\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p})$
$\leq$ $\mathcal{W}_{k}({\mu}_{\mathrm{post}},{\mu}_{\mathrm{pre}})+\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p})$
$\leq$ $\mathcal{W}_{k}(\mu_{\mathrm{post}},\mu_{p})+2\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p}).$
Now the desired result follows from Lemma 8. ∎
Appendix B Scalar filter algorithm
The following summarizes the scalar filter algorithm of Peña and Tsay, (2021).
  • Initialize the imputation $\hat{\mathbf{w}}^{(0)}=(\hat{w}_{0}^{(0)},\hat{w}_{1}^{(0)},\ldots,\hat{w}_{n-1}^{(0)})^{\top}$ with $\hat{w}_{t}^{(0)}=0$ if $t\in\mathcal{M}$.
  • For each $s\in\mathcal{M}$, let $I^{(s)}_{t}$ be the indicator function. That is,
    $I_{t}^{(s)}=\left\{\begin{array}[]{cc}1,&t=s\\ 0,&t\neq s\end{array}\right.$
    and estimate the intervention model using $\hat{\mathbf{w}}^{(0)}$. For example, one can fit an ARX model with $\{I_{t}^{(s)}:s\in\mathcal{M}\}$ as exogenous predictors.
  • Let $\hat{\beta}^{(s)}$ be the intervention effect associated with $I_{t}^{(s)}$. Estimate the missing data by
    $\hat{w}_{s}=\hat{w}_{s}^{(0)}-\hat{\beta}^{(s)},\quad s\in\mathcal{M}.$
As one can see from the algorithm above, the scalar filter algorithm proposed by Peña and Tsay, (2021) is based on intervention analysis. It can be iterated in order to improve the estimates. However, no convergence guarantee is provided by the authors.
When fitting an ARX model with the indicators as exogenous predictors, if there are many missing values, standard approaches may be infeasible or unreliable due to near collinearity. In this case, we employ a small $\ell_{2}$ penalty to avoid numerical instability.

### v2

@@
 Abstract
-Missing data often significantly hamper standard time series analysis, yet in practice they are frequently encountered.
+Missing data can significantly hamper standard time series analysis, yet in practice they are frequently encountered.
 In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series.
 Unlike existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for potential nonlinear dynamics.
-Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any missing pattern.
+Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any non-systematic missing pattern.
 In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated.
@@
 Leveraging the benign landscape of the optimization formulation, we establish the convergence of an alternating minimization algorithm to critical points.
-Furthermore, we provide conditions under which the marginal distributions of the underlying time series can be identified.
-Our numerical experiments, including extensive simulations covering linear and nonlinear time series models and an application to a real-world groundwater dataset laden with missing data, corroborate the practical usefulness of the proposed method.
-Keywords— Missing data, Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
+We also provide conditions under which the marginal distributions of the underlying time series can be identified.
+Numerical experiments, including extensive simulations covering linear and nonlinear time series models and a real-world groundwater dataset laden with missing values, corroborate the practical usefulness of the proposed method.
+Keywords: Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
 1 Introduction
-In many scientific fields, data collection technology is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data-collection policies, leading to missing entries in the data (Little and Rubin,, 2019).
-In time series analysis, missing data pose serious challenges to the application of most statistical tools since it is typically presumed that observations are available at equally spaced time points.
+In many scientific fields, data collection is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data collection policies, leading to missing entries in the data (Little and Rubin,, 2019).
+In time series analysis, missing data pose serious challenges to the application of many statistical tools since it is typically presumed that observations are available at equally spaced time points.
 Moreover, since the dynamic relationship is of crucial importance in time series analysis, one cannot proceed by throwing away the missing data, even if the missing data appear completely at random (i.e., the missing pattern is independent of the series).
-To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods can be readily applied for statistical inference in the downstream.
-Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021).
-Most of these methods are iterative or EM-type algorithms that alternate between model identification and missing data imputation (see, for example, Chpt.
+To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods becomes readily applicable for statistical inference in the downstream.
+Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021; McElroy and Politis,, 2020, 2022; McElroy,, 2022).
+Since the optimal interpolator would depend on the probabilistic structure of the underlying process, most methods involve EM-type iterations that alternate between model identification and missing data imputation (see, for example, Chpt.
 4 of Peña and Tsay,, 2021).
-For certain time series models, one can estimate the model parameters simultaneously with the missing data (Harvey and Pierse,, 1984; Gómez et al.,, 1999).
 Besides the optimal interpolator, ad hoc deterministic imputation methods, such as smoothing with linear interpolation, B-splines and wavelets (see, for instance, Hastie et al.,, 2009), are also used in practice for their simplicity.
-However, the optimal interpolator can introduce distributional biases that are detrimental to downstream statistical analysis.
-To illustrate, consider an AR(1) model, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d.
+However, the optimal interpolator can cause distributional biases that are detrimental to downstream statistical analysis.
+To illustrate, consider an AR(1) process, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d.
 standard normal random variables.
-It is well-known that the optimal interpolator can be expressed in terms of the inverse autocorrelation function (Peña and Maravall, 1991; see also Chpt.
+The optimal interpolator can be expressed in terms of the inverse autocorrelation functions (Peña and Maravall, 1991; McElroy and Politis, 2022; see also Chpt.
 6 of Peña et al.,, 2011).
-For instance, suppose that data at time $T$ and $T+1$ are missing; then the optimal interpolators are given by
-$\hat{x}_{T}$ $=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},$
-$\hat{x}_{T+1}$ $=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
-Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$ for $h\geq 0$, it is not difficult to show
-$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),$
-and
+For instance, suppose that data at time indices $T$ and $T+1$ are missing.
+Then the optimal interpolators are given by
+$\hat{x}_{T}=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},\quad\hat{x}_{T+1}=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
+Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$, for $h\geq 0$, it is not difficult to show that
+$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),\mbox{ and}$
 $|\mathbb{E}(\hat{x}_{T}\hat{x}_{T+1})|=\frac{2\phi^{2}+3\phi^{4}+3\phi^{6}+\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}|\mathbb{E}(x_{t}x_{t+1})|<|\mathbb{E}(x_{t}x_{t+1})|.$
 Hence, the variance and the autocovariances of the imputed data are different from those of the original AR(1) process.
-Intuitively, the optimal interpolator for consecutive missing data tends to predict the unconditional mean, leading to a downward bias (in magnitude) of the variances.
 In general, the joint distribution of the optimally interpolated time series is often quite different from that of the original process, leading to questionable downstream statistical analysis, especially when the number of missing data is non-negligible.
@@
 In addition to the distributional bias, the above example also reveals further predicaments of the optimal interpolator.
-First, in order to compute the optimal interpolator, one often needs to choose a model class (such as the ARMA models) before imputation.
-However, such decision relies heavily on domain expertise and may be ill-informed when there are many missing values.
+First, in order to compute the optimal interpolator, one often needs to choose a model class before imputation.
+However, such a decision relies heavily on domain expertise and may be ill-informed when there are many missing values.
+While it is possible, for certain time series models, to estimate the model parameters without imputation, it still engages in model specification before imputation.
 Second, it may be difficult to compute the optimal interpolator if some nonlinear model is employed.
@@
 Side information about the missing data restricts the set of admissible imputations.
-For instance, compositional time series is known to lie in $[0,1]$ and sums to unity at each time point.
-Besides the optimal interpolator, it is well-known that some deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges (Carrizosa et al.,, 2013).
-In this paper, we present a novel procedure for time series imputation, referred to as the temporal Wasserstein imputation (TWI), which circumvents the aforementioned limitations.
+For instance, a compositional time series is known to lie in $[0,1]$ and sums to unity at each time point.
+Besides the optimal interpolator, many deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges.
+In this paper, we present the temporal Wasserstein imputation (TWI), a novel approach for time series imputation which circumvents the aforementioned limitations.
 As a nonparametric method, TWI is highly versatile and can be applied to a wide range of time series data.
-In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive model (TAR, Tong,, 1983).
+In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive processes (TAR, Tong,, 1983).
 Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm.
@@
 Section 3 maps out the optimization formulation and introduces the computational scheme in detail.
-The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the $p$-dimensional marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series (which is stable under stationarity), thereby lowering the distributional bias discussed above.
+The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series, thereby lowering the distributional bias discussed above.
 In Section 4, we investigate the optimization and statistical properties of TWI.
@@
 However, from the statistical viewpoint, missing data imputation suffers from an identification issue: Many distributions (of the stochastic process) may be equally plausible in describing the corrupted series.
-This raises the question of when and under what conditions TWI can identify the underlying distribution and yield imputations that can be safely used in downstream analysis.
-In Section 4, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation.
+This raises the question of whether and under what conditions TWI can identify the underlying marginal distributions.
+In Section 4.2, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation.
 The latter condition motivates a simple modification of TWI, termed $k$-TWI, which shows further improvement in imputation quality from TWI in our numerical experiments.
 In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings.
-Across all models, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates.
-Its advantage is particularly prominent in the presence of nonlinear dynamics, where TWI consistently outperforms the benchmark methods.
+Across all processes considered, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates.
+Its advantage is particularly prominent in the presence of nonlinear dynamics.
 In Section 6, we apply TWI to a real-world groundwater dataset with many missing data.
-After imputation, we conduct factor analysis and autoregressive modeling in the downstream.
-Compared to existing methods, TWI effectively captures the underlying dynamics, producing more reasonable factor estimates as well as more accurate autoregressive predictions.
-These numerical results underscore the practical utility of TWI.
+After imputation, we conduct independent component analysis and autoregressive modeling in the downstream.
+Compared to existing methods, TWI effectively captures the underlying dynamics, producing interpretable components as well as accurate autoregressive predictions.
+These numerical results support the practical contributions of TWI.
 2 Preliminary: Optimal transport
-In this section, we give a brief introduction to the optimal transport (OT) problem, which serves as a prerequisite to the rest of the paper.
+In this section, we briefly review the optimal transport (OT) problem, which serves as a prerequisite to the rest of the paper.
 For a more comprehensive account on the theory of optimal transport, the readers are referred to Villani, (2021, 2008).
-Given two probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{p}$, and a cost function $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{+}$, the optimal transport cost between $\mu$ and $\nu$ is defined as
+Given two probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{p}$ and a cost function $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{+}$, the optimal transport cost between $\mu$ and $\nu$ is defined as
 […]
 […]
-If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$ for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$.
+If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$, for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$.
 Namely,
@@
 9).
-In particular, if $\mu$ and $\{\mu_{t}:t=1,2,\ldots\}$ are probability measures on $\mathbb{R}^{p}$ such that $\mathcal{W}_{k}(\mu_{t},\mu)\rightarrow 0$ as $t\rightarrow\infty$, then $\mu_{t}\Rightarrow\mu$, where $\Rightarrow$ denotes the weak convergence of probability measures (Billingsley,, 1995).
+In particular, if $\mu$ and $\{\mu_{t}:t=1,2,\ldots\}$ are probability measures on $\mathbb{R}^{p}$ such that $\mathcal{W}_{k}(\mu_{t},\mu)\rightarrow 0$ as $t\rightarrow\infty$, then $\mu_{t}\Rightarrow\mu$ and $\int\|\mathbf{x}\|^{k}\mu_{t}(d\mathbf{x})\rightarrow\int\|\mathbf{x}\|^{k}\mu(d\mathbf{x})$, where $\Rightarrow$ denotes the weak convergence of probability measures (Billingsley,, 1995).
 Therefore, it can be used as a discrepancy metric between $\mu$ and $\nu$.
-In fact, the Wasserstein distance offers several advantages over other discrepancy metrics such as the Kullbeck-Leibler divergence or the Kolmogorov-Smirnov distance.
-In addition to its metric properties, the Wasserstein distance has a nice dual representation and handles support mismatch easily, which makes it suitable in modern machine learning and statistical applications.
+In fact, the Wasserstein distance offers several advantages over other discrepancy metrics such as the Kullbeck-Leibler divergence and the Kolmogorov-Smirnov distance.
+For instance, the Wasserstein distance has a nice dual representation and handles support mismatch easily, making it suitable for machine learning and statistical applications.
 If $\mu$ and $\nu$ are discrete measures, for instance $\mu=\sum_{i=1}^{n}p_{i}\delta_{\mathbf{x}_{i}}$ and $\nu=\sum_{j=1}^{m}q_{j}\delta_{\mathbf{y}_{j}}$, where $\{\mathbf{x}_{i}\}_{i=1}^{n},\{\mathbf{y}_{j}\}_{j=1}^{m}$ are in $\mathbb{R}^{p}$, $\sum_{i=1}^{n}p_{i}=\sum_{j=1}^{m}q_{j}=1$, $p_{i}>0$, $q_{j}>0$ for all $i,j$, and $\delta_{\mathbf{x}}$ denotes the dirac measure at $\mathbf{x}$, then the minimization problem becomes the following linear program.
-$\mathsf{OTC}(\mu,\nu)=\min_{\{\pi_{ij}\}}$ $\sum_{i,j}L(\mathbf{x}_{i},\mathbf{y}_{j})\pi_{ij}$
-s.t. $\sum_{i=1}^{n}\pi_{ij}=q_{j},\quad j=1,\ldots,m,$ (1)
-$\sum_{j=1}^{m}\pi_{ij}=p_{i},\quad i=1,\ldots,n,$ (2)
-$\pi_{ij}\geq 0.$
-The constraints (1) and (2) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints.
-In the discrete case, the optimal coupling $\{\pi_{ij}\}$, in general, cannot be expressed as the push-forward under an one-to-one map.
-Instead, the optimal coupling typically involves split of masses for discrete measures.
-Many computational methods have been proposed to solve this linear program, including the network flow simplex (Orlin,, 1997) and the auction method (Bertsekas,, 1992).
-When the problem dimensions $n,m$ are large, standard tools for linear programs may suffer from slow computation.
-Nevertheless, recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
+$\mathsf{OTC}(\mu,\nu)=\min_{\{\pi_{ij}\},\pi_{ij}\geq 0}$ $\sum_{i,j}L(\mathbf{x}_{i},\mathbf{y}_{j})\pi_{ij}$
+s.t. $\sum_{i=1}^{n}\pi_{ij}=q_{j},\quad j=1,\ldots,m,$ $\sum_{j=1}^{m}\pi_{ij}=p_{i},\quad i=1,\ldots,n.$ (1)
+The constraints in (1) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints.
+In the discrete case, the optimal coupling $\{\pi_{ij}\}$ typically involves split of masses for discrete measures and is no longer the push-forward under an one-to-one map.
+Recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
 3 Temporal Wasserstein imputation
-In this section, we introduce temporal Wasserstein imputation (TWI), the proposed method for time series imputation, through an optimal transport formulation.
-First, an optimization formulation of the proposed method that directly connects to the optimal transport problem is laid out.
+In this section, we introduce the proposed temporal Wasserstein imputation (TWI) method for time series imputation.
+First, we provide an optimization formulation of the proposed method that directly connects to the optimal transport problem.
 Then, we describe an alternating minimization algorithm to impute the missing values.
-The algorithm also reveals some connections between the proposed method and existing approaches.
+The algorithm also reveals some connections between the proposed method and some existing approaches.
 3.1 Optimization formulation
-Let $\{\tilde{x}_{t}\}$ be the time series of interest, and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$.
+Let $\{\tilde{x}_{t}\}$ be the time series of interest and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$.
 Denote by $\bigstar$ a generic missing value, and $\mathcal{M}$ the set of time indices corresponding to the missing observations.
-Then, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise.
-If $\{\tilde{x}_{t}\}$ is stationary, it follows that its $p$-dimensional marginal distributions, i.e.
-the probability distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, are stable over time.
-Let $\mu_{p}$ ($p\in\mathbb{N}$) be the probability measure associated with $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$.
-Stationarity implies $\mu_{p}$ is independent of the time index $t$.
-In light of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distributions remain nearly unchanged before and after a specified time point $n_{1}$.
-More specifically, we first divide the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$, where $p-1<n_{1}<n-p$, and define the two empirical marginal distributions
-$\hat{\mu}_{\mathrm{pre}}(\mathbf{w})=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})},$
-$\hat{\mu}_{\mathrm{post}}(\mathbf{w})=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}.$
-The discrepancy between the two empirical distributions can then be measured by the optimal transport cost, which is defined as the minimum value of the following linear program.
-$\begin{split}\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij}}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t.
-}&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\\ &\pi_{ij}\geq 0\mbox{ for all $i,j$},\end{split}$ (3)
+Thus, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise.
+We discuss below the case that $x_{t}$ is stationary.
+The case of unit-root nonstationary series is discussed at the end of this section.
+If $\{\tilde{x}_{t}\}$ is stationary, then its $p$-dimensional marginal distribution $\mu_{p}$, i.e.
+the distributions of $(\tilde{x}_{t},\tilde{x}_{t-1},\ldots,\tilde{x}_{t-p+1})$, does not depend on $t$, where $p\in\mathbb{N}$.
+In view of this property, when imputing the missing values in $(x_{0},\ldots,x_{n-1})$, we seek an imputation $\mathbf{w}=(w_{0},w_{1},\ldots,w_{n-1})^{\top}$ whose marginal distribution remains nearly unchanged.
+More specifically, we first split the time series into two parts $\{w_{0},w_{1},\ldots,w_{n_{1}}\}$ and $\{w_{n_{1}+1},w_{n_{1}+2},\ldots,w_{n-1}\}$ and define the two empirical marginal distributions
+$\hat{\mu}_{\mathrm{pre}}(\mathbf{w})=\sum_{t=p-1}^{n_{1}}\frac{\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}}{n_{1}-p+2},\quad\hat{\mu}_{\mathrm{post}}(\mathbf{w})=\sum_{t=n_{1}+1}^{n-1}\frac{\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}}{n-n_{1}-1},$
+where $p-1<n_{1}<n-p$, and the values of $p$ and $n_{1}$ used will be discussed later.
+The discrepancy between the two empirical distributions can then be measured by the optimal transport cost, which is the minimum value of the following linear program.
+$\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij},\pi_{ij}\geq 0}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t.
+}&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\quad\mbox{for all }j\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\quad\mbox{for all }i,\end{aligned}\right\}$ (2)
 where $\mathbf{v}_{i}(\mathbf{w}):=(w_{i},w_{i-1},\ldots,w_{i-p+1})^{\top}$, and, for ease of notation, we enumerate the rows and columns of the $(n_{1}-p+2)\times(n-n_{1}-1)$ matrix $\mathbf{\Pi}$ by $\{p-1,p,\ldots,n_{1}\}$ and $\{n_{1}+1,\ldots,n-1\}$, respectively.
-An important special case of (3) is when $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{k}$ for some $k\geq 1$.
+An important special case of (2) is when $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{k}$, for some $k\geq 1$.
 In this case, we have $\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})=\mathcal{W}_{k}^{k}(\hat{\mu}_{\mathrm{pre}},\hat{\mu}_{\mathrm{post}})$.
@@
 It may also reflect the side information available.
-For example, if $x_{t}$ are weekly sales data, the knowledge of monthly sales define a set of linear constraints for the missing values.
+For example, if $x_{t}$ are monthly sales data, the knowledge of quarterly sales define a set of linear constraints for the missing values.
 Another example is when $\{\mathbf{x}_{t}\}$ is a $d$-dimensional compositional time series such as the market share of several companies in an industry.
@@
 $\hat{\mathbf{w}}:=(\hat{w}_{0},\ldots\hat{w}_{n-1})^{\top}=\arg\min_{\mathbf{w}\in\mathcal{C}}\left\{\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))+\frac{\lambda}{2}\|\mathbf{w}\|^{2}\right\},$ (4)
-where $\lambda>0$ is small.
+where the regularization parameter $\lambda>0$ is small.
 As we will see in the next subsection, the $\ell_{2}$-regularization term makes a subproblem in the alternating minimization algorithm strongly convex, enhancing numerical stability.
@@
 In this formulation, it is clear that the temporal Wasserstein imputation, which seeks to equate the marginal distributions on both sides of $n_{1}$, is nonparametric.
-In addition, with a mildly large $p$, the $p$-dimensional marginal distributions can fully capture the nonlinear dynamics of a Markov process (such as a finite-order TAR model), which would otherwise need to be approximated using many parameters in an ARMA model.
+With a mildly large $p$, the $p$-dimensional marginal distributions can sometimes fully characterize the nonlinear dynamics of a Markov process (such as a finite-order threshold AR process), which would otherwise need many parameters using ARMA approximation.
 3.2 Alternating minimization algorithm
 The optimization landscapes of (3) and (4) are complicated because the optimal transport cost is itself the minimum value of a linear program (2), and there are potentially many variables to optimize.
-This makes directly optimizing (4) or (5) computationally intractable.
-To remedy this issue, we propose to solve the nested minimization problem using alternating minimization.
+To tackle this issue, we propose to solve the nested minimization problem using alternating minimization.
 Since (3) is a special case of (4) with $\lambda=0$, we focus our discussion on (4).
 To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (2) is exactly the discrete optimal transport problem.
-The minimizer $\hat{\mathbf{\Pi}}$ can be efficiently computed using existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019).
-Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (3), which corresponds to the inner optimization problem of (5).
+Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (2) efficiently by some existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019).
+This corresponds to the inner optimization problem of (4).
 Next, we optimize (4) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed.
@@
 Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$.
-Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where
-$\mathbf{H}(\mathbf{\Pi})$ $=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n},\mbox{ and}$
+Then it is not difficult to show that $F(\mathbf{w},\mathbf{\Pi})=\mathbf{w}^{\top}\mathbf{H}(\mathbf{\Pi})\mathbf{w}$, where $\mathbf{H}(\mathbf{\Pi})=\sum_{h=0}^{p-1}\mathbf{A}_{h}(\mathbf{\Pi})+\frac{\lambda}{2}\mathbf{I}_{n}$ and
 $\mathbf{A}_{h}(\mathbf{\Pi})$ $=\begin{pmatrix}\mathbf{0}_{p-1-h}&&&\\ &\frac{1}{n_{1}-p+2}\mathbf{I}_{n_{1}-p+2}&-\mathbf{\Pi}&\\ &-\mathbf{\Pi}^{\top}&\frac{1}{n-n_{1}-1}\mathbf{I}_{n-n_{1}-1}&\\ &&&\mathbf{0}_{h}\end{pmatrix}\in\mathbb{R}^{n\times n},\quad h=0,1,\ldots,p-1,$
@@
 Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004; Parikh et al.,, 2014).
-If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e.,
-$\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$
-for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
+If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e., $\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$ for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
 $\hat{\mathbf{w}}^{(k)}=\mathbf{H}^{-1}\mathbf{K}^{\top}(\mathbf{K}\mathbf{H}^{-1}\mathbf{K}^{\top})^{-1}\mathbf{b},$ (6)
@@
 $w_{n_{1},\mathrm{opt}}=\frac{\phi}{1+\phi^{2}}(x_{n_{1}-1}+x_{n_{1}+1}).$
-In contrast, (8) shows that the temporal Wasserstein imputation (with $p=2$) is given by
+In contrast, (7) shows that the temporal Wasserstein imputation (with $p=2$ and $\lambda=0$) is given by
 $\hat{w}_{n_{1}}=\frac{1}{n_{1}^{-1}+(n-n_{1}-1)^{-1}}\left(\sum_{t=0}^{n_{1}-1}\hat{\pi}_{t+1,n_{1}+1}x_{t}+\sum_{t=n_{1}+1}^{n-1}\hat{\pi}_{n_{1},t}x_{t}\right),$
-which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI shall use all data to learn the underlying dynamics.
+which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI uses all data to learn the underlying dynamics.
 Remark 2.
 (EM interpretation) Let $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ be a fixed point of the alternating minimization algorithm (Algorithm 1). That is,
-$\hat{\mathbf{w}}\in$ $\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),$
-$\hat{\mathbf{\Pi}}\in$ $\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
+$\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),\quad\hat{\mathbf{\Pi}}\in\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
 In addition, suppose $\lambda=0$ for a moment.
 Let $\mathbf{u}=(u_{0},u_{1},\ldots,u_{p-1})$ be a random vector whose distribution is $\hat{\mu}_{\mathrm{pre}}$ and $\mathbf{v}=(v_{0},\ldots,v_{p-1})$ a random vector whose distribution is $\hat{\mu}_{\mathrm{post}}$.
-Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, which is $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures).
+Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, identified as $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures).
 Consider an index $s\in\mathcal{M}$ with $p-1\leq s\leq n_{1}-p+1$.
@@
 $\hat{w}_{s}=$ $\frac{1}{p}\sum_{i=s}^{s+p-1}\sum_{j=n_{1}+1}^{n-1}\frac{\hat{\pi}_{ij}}{\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}}\hat{w}_{j-i+s}$
-$=$ $\frac{1}{p}\sum_{i=s}^{s+p-1}\sum_{j=n_{1}+1}^{n-1}\mathbb{P}_{\hat{\mathbf{\Pi}}}\left(\mathbf{v}=(\hat{w}_{j},\ldots,\hat{w}_{j-p+1})|\mathbf{u}=(\hat{w}_{i},\ldots,\hat{w}_{i-p+1})\right)\hat{w}_{j-i+s}$
-$=$ $\frac{1}{p}\sum_{h=0}^{p-1}\mathbb{E}_{\hat{\mathbf{\Pi}}}(v_{h}|\mathbf{u}=(\hat{w}_{s+h},\ldots,\underbrace{\hat{w}_{s}}_{\mbox{\makebox[0.0pt]{$h$-th element}}},\ldots,\hat{w}_{s+h-p+1})).$ (9)
+$=$ $\frac{1}{p}\sum_{h=0}^{p-1}\mathbb{E}_{\hat{\mathbf{\Pi}}}(v_{h}|\mathbf{u}=(\hat{w}_{s+h},\ldots,\hat{w}_{s},\ldots,\hat{w}_{s+h-p+1})).$ (8)
 (8) shows that the temporal Wasserstein imputation $\hat{w}_{s}$ is essentially an average of conditional expectations.
@@
 Formula like (8) can be derived for other time indices $s$ with appropriate modifications.
-Remark 3.
-(Barycenter interpretation) The summands in (9) is also known as the barycentric projections (Ambrosio et al.,, 2006; see also Deb et al.,, 2021).
-Hence the temporal Wasserstein imputation can also be interpreted as an average of barycentric projections under the coupling $\hat{\mathbf{\Pi}}$.
 Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series.
-Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top},t=0,1,\ldots,n-1$.
-To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (3) by
-$\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}.$
+Suppose the multivariate time series of interest is $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$.
+To compute the imputation $\mathbf{W}=(\mathbf{w}_{0},\mathbf{w}_{1},\ldots,\mathbf{w}_{n-1})$ where $\mathbf{w}_{t}=(w_{t,1},\ldots,w_{t,d})^{\top}$, we may replace $\mathbf{v}_{i}(\mathbf{w})$ in (2) by $\mathbf{v}_{i}(\mathbf{W}):=(w_{i,1},\ldots,w_{i-p+1,1},w_{i,2},\ldots,w_{i-p+1,2},\ldots,w_{i-p+1,d})^{\top}$.
 Then Algorithm 1 can still be applied.
@@
 Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (7) in the case of no side information.
-Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel, and the imputation only depends on the future and past values of individual series.
-The cross-sectional information is only exploited when learning the optimal transport coupling in step (a).
+Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel to lower the computational overhead and only the values of individual series are involved in computing the imputation in each component.
+The cross-sectional information is exploited when learning the optimal transport coupling in step (a).
 Thus, the algorithm is fast to implement in the multivariate setting.
-We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by existing imputation methods.
-Their method corrects the resulting statistical pattern by penalizing the deviation of some moments from the theoretical values.
-For instance, let $\rho_{k}$, $k=1,2,\ldots,k_{0}$ be some pre-specified values for the autocorrelation function at lag $k$, and $m_{k}$, $k=1,2,\ldots,k_{1}$, the pre-specified values for the $k$-th moments.
-Then imputation is done by solving the following program.
-$\min_{\mathbf{w}\in\mathcal{C}}\frac{f(\mathbf{w})}{f(\mathbf{w}^{(0)})}+\sum_{k=1}^{k_{0}}\lambda_{k}\left(\frac{\rho_{k}(\mathbf{w})}{\rho_{k}}-1\right)^{2}+\sum_{k=1}^{k_{1}}\mu_{k}\left(\frac{n^{-1}\sum_{t=0}^{n-1}w_{t}^{k}}{m_{k}}-1\right)^{2},$
-where $\rho_{k}(\mathbf{w})$ is the sample autocorrelation function computed from the series $\mathbf{w}$, $f(\mathbf{w})=\sum_{t=1}^{n-1}(w_{t}-w_{t-1})^{2}$, and $\mu_{k},\lambda_{k}\geq 0$ are the penalty strengths.
-As a result, the moments and the autocorrelation functions of the imputed series will be close to the pre-specified values.
+We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by imputation.
+Their method is based on penalizing the deviation of some moments from the theoretical values.
+As a result, the moments, such as the autocorrelation functions, of the imputed series are close to the pre-specified values.
 However, this approach has a number of drawbacks.
-First, a good theoretical value for $\rho_{k}$ and $m_{k}$ may not be available in practice, especially when the missing values are many.
-Second, minimizing $f$ promotes smoothness and a positive correlation between $w_{t}$ and $w_{t-1}$, which does not necessarily align with the underlying time series.
-Finally, the optimization problem is highly non-convex, leading to intense computational hurdle.
-In contrast, the temporal Wasserstein imputation is data-driven and completely nonparametric, requiring small amount of tuning.
-It also enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
-In closing this subsection, we discuss the imputation of nonstationary time series.
-For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and we can apply imputation methods to the first-differenced series $\{y_{t}\}$.
-In this case, the number of missing data multiplies in the first-differenced series and could compromise the imputation performance.
-However, for TWI we can mitigate this issue by encoding the information from the raw data via the admissible set $\mathcal{C}$.
-For example, if $x_{t}$ is observed, then $y_{1}+y_{2}+\ldots+y_{t}=x_{t}-x_{0}$, so we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed).
-Finally, although we have motivated TWI from the stationarity of the underlying process, we found that it fares quite well with certain nonstationary processes such as cyclic time series.
-In Section 5, we further explore the application of TWI to nonstationary time series via simulation studies.
+First, good theoretical values for the moments may not be readily available in practice, especially when the missing values are many.
+Second, their approach promotes smoothness of the time series and hence positive autocorrelation, which does not necessarily align with the underlying time series.
+Finally, the optimization problem is highly non-convex, leading to an intense computational hurdle.
+In contrast, the proposed TWI is data-driven, completely nonparametric, and enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
+Next, we discuss the imputation of nonstationary time series.
+For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and one can apply imputation methods to the first-differenced series $\{y_{t}\}$.
+More generally, one can apply TWI to the series $y_{t}=\phi(B)x_{t}$, where $\phi(z)=(1-z)^{d}$ if $x_{t}$ is an $I(d)$ process, and $B$ is the back-shift operator.
+Then, an imputation for $x_{t}$ can be recovered by inverting $\phi(B)$.
+Note that, in this case, the number of missing data in $\{y_{t}\}$ multiplies, and we need at least $d+1$ contiguous observed values in $\{x_{t}\}$ to obtain an observed value for $\{y_{t}\}$.
+To mitigate the information loss, we can encode the information from the raw data via the admissible set $\mathcal{C}$.
+For example, consider the $I(1)$ case: If $x_{t}$ is observed, then since $y_{1}+\ldots+y_{t}=x_{t}-x_{0}$, we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed).
+Finally, although we have motivated TWI from the stationarity of the underlying process, TWI fares quite well with certain nonstationary processes such as cyclic time series.
+In Section 5, we further explore the application of TWI to nonstationary time series.
 4 Theoretical properties
 In this section, we investigate the optimization and statistical properties of TWI.
-Under some general assumptions, we first show Algorithm 1 converges to critical points.
-Then, the asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$.
-For general missing patterns, we show that identification of the underlying marginal distributions is possible under the assumptions of asymmetric missing pattern and interpolator stability (see (C1) and (C2) in Section 4.2) using a two-state Markov chain example.
-In the following, we maintain the assumption that the index set $\mathcal{M}=\mathcal{M}_{n}$, which corresponds to the missing data, is fixed and independent of the underlying time series.
-It is worth noting that we allow the number of missing entries to scale with the sample size, i.e., $\sharp(\mathcal{M}_{n})\rightarrow\infty$ as $n\rightarrow\infty$, where $\sharp(E)$ is the cardinality of a set $E$.
-This asymptotic regime is more suitable for applications where non-negligible portions of the data are missing.
-In addition, our focus is on the distributional bias induced by the imputations.
-If $\sharp(\mathcal{M}_{n})=o(n)$, the problem becomes trivial because the empirical marginal distributions of any (bounded) imputations satisfy
+We first show Algorithm 1 converges to critical points.
+Then, asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$.
+For general non-systematic missing patterns, we show that identification of the underlying marginal distributions can be achieved by TWI under the assumptions of interpolator stability using a two-state Markov chain example (see (C1) and (C2) in Section 4.2).
+Before presenting the results, it is worthwhile to clarify the asymptotic regime under consideration.
+The analysis is meaningful only when the number of missing entries is non-negligible compared to the sample size.
+Otherwise, if $\sharp(\mathcal{M})=o(n)$, where $\sharp(E)$ is the cardinality of the set $E$, then the empirical marginal distributions of any (bounded) imputations satisfy
 $\frac{1}{n}\sum_{t\in\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}+\frac{1}{n}\sum_{t\notin\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}\Rightarrow\mu_{p},$
-where $\mathcal{A}_{n}=\{t:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}_{n}\neq\emptyset,t\in\{0,1,\ldots,n-1\}\}$.
+where $\mathcal{A}_{n}=\{t\in\{p-1,\ldots,n-1\}:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}\neq\emptyset\}$.
+Therefore, assuming $\sharp(\mathcal{M})=o(n)$ would essentially preclude distributional inconsistency in the downstream.
+To avoid this trivial case, in this paper we maintain instead the assumption that $\mathcal{M}=\mathcal{M}_{n}$ is a sequence of index sets (independent of the underlying time series) such that $\sharp(\mathcal{M}_{n})/n\asymp 1$, where $a_{n}\asymp b_{n}$ means there exists $0<\underline{c}<\bar{c}<\infty$ such that $\underline{c}\leq a_{n}/b_{n}\leq\bar{c}$ for all large $n$.
 4.1 Optimization and statistical guarantees
-We begin by studying the optimization properties of Algorithm 1.
+We begin by studying the optimization properties of Algorithm 1. All proofs are relegated to Section S1 of the supplementary material.
 Proposition 3.
-Suppose $\mathcal{C}$ is a convex set. Then,
+Suppose $\mathcal{C}$ is a convex set and $F(.,.)$ is defined in (5). Then,
   • (a)
-    (monotonicity) $F(\mathbf{w}^{(t)},\mathbf{\Pi}^{(t+1)})\leq F(\mathbf{w}^{(t)},\mathbf{\Pi}^{(t)})\leq F(\mathbf{w}^{(t-1)},\mathbf{\Pi}^{(t)})$, $t=1,2,\ldots$;
+    (monotonicity) $F(\hat{\mathbf{w}}^{(t)},\hat{\mathbf{\Pi}}^{(t+1)})\leq F(\hat{\mathbf{w}}^{(t)},\hat{\mathbf{\Pi}}^{(t)})\leq F(\hat{\mathbf{w}}^{(t-1)},\hat{\mathbf{\Pi}}^{(t)})$, $t=1,2,\ldots$;
   • (b)
-    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\mathbf{w}^{(t+1)},\mathbf{\Pi}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. That is,
+    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\hat{\mathbf{w}}^{(t+1)},\hat{\mathbf{\Pi}}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. i.e.,
     $\frac{\partial F(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})}{\partial\mathbf{w}}\cdot(\mathbf{w}-\hat{\mathbf{w}})\geq 0,\quad\frac{\partial F(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})}{\partial\mathrm{vec}(\mathbf{\Pi})}\cdot\mathrm{vec}(\mathbf{\Pi}-\hat{\mathbf{\Pi}})\geq 0$
@@
     $F(\mathbf{w},\mathbf{\Pi})$ is biconvex in $\mathbf{w}$ and $\mathbf{\Pi}$.
-Proof.
+Due to the biconvexity of $F$, one can employ off-the-shelf solvers for the subproblems in Algorithm 1.
+Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$.
+Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
+A few remarks are in order.
+First, it is straightforward to generalize Proposition 3 for multivariate time series.
+Second, like most methods involving non-convex optimization, a good initialization is crucial for converging to a “good” critical point.
+In our simulation studies, we use simple linear interpolation or ARIMA-based Kalman smoothing as initializations and TWI is able to find an imputation that offers substantial improvements over these initializations.
+Third, unlike many non-convex optimization problems, we are not satisfied by an arbitrary global minimizer of $F$.
+Essentially, it is because among all processes that have stable $p$-dimensional marginal distributions, only a subset of them matches the underlying marginal distributions.
+In Section 4.2, we probe this identification issue in more detail.
+However, when the missing data happen to lie on the same side of $t=n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 4 below.
+Proposition 4.
+Assume (A1) and (A2) in Section S1 of the supplementary material.
+Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$.
+Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$.
+If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$ with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, subject to $\mathbf{w}\in\mathcal{C}$ and the coupling conditions, satisfies
+$\mathcal{W}_{k}\left(\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)+\mathcal{W}_{k}\left(\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)\rightarrow 0\quad\mbox{almost surely}.$
+[…]
+4.2 General missing pattern: A two-state Markov chain example
+In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example.
+Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations.
+We also show that, under two natural conditions, the correct marginal distributions can still be pinned down by TWI.
+To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
+Suppose the underlying time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain. Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by
+$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.6,\quad$ $\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.2.$
+The transition probabilities are specified to facilitate discussions, but the main results below are not specific to these choices.
+It is easy to check that $\{\tilde{x}_{t}\}$ has the stationary distribution given by $\mathbb{P}(\tilde{x}_{t}=1)=0.75$ and $\mathbb{P}(\tilde{x}_{t}=0)=0.25$.
+Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
+$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=0.6,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=0.15,$ (9)
+$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=0.15,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=0.1.$
+Without loss of generality, we set $n_{1}=0$.
+Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations.
+Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$.
+Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$ and $q_{2}$ are some integers larger than 2 and $q_{1}\neq q_{2}$ (see condition (C1) below).
+Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ (10)
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$
+and if $t\geq n_{1}$,
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ (11)
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$
+for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$.
+Note that $\mathbb{P}$ in (10) and (11) should be interpreted as proportions since the imputation may depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined.
+Essentially, here we parametrize all imputation methods by $a_{1},a_{2},b_{1},b_{2}$, which correspond to different 2-dimensional marginal distributions.
+Indeed, from (9)–(11), we can compute the 2-dimensional marginal distributions of the imputed series.
+For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-\frac{1}{q_{1}})+\frac{3a_{1}}{4q_{1}},\quad\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))=0.15(1-\frac{1}{q_{1}})+\frac{b_{1}}{4q_{1}},$
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))$ $=0.15(1-q_{1}^{-1})+\frac{3}{4q_{1}}(1-a_{1}),$
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))$ $=0.1(1-q_{1}^{-1})+\frac{1}{4q_{1}}(1-b_{1}).$
+For $\{\hat{w}_{t}\}_{0\leq t<\infty}$, we can obtain similar expressions with $a_{1},b_{1},q_{1}$ replaced by $a_{2},b_{2},q_{2}$.
+The “correct” imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting 2-dimensional marginal distributions equal the marginal distribution of the underlying series, given in (9).
+To achieve such identification, the correct imputation must choose
+$a_{1}^{*}=a_{2}^{*}=0.8,\quad b_{1}^{*}=b_{2}^{*}=0.6.$ (12)
+Note that this calibration does not depend on $q_{1},q_{2}$; it only depends on the distributional properties of the underlying process.
+Assuming the optimization is carried out perfectly [^1] This is where we ignore finite sample complications and assume a zero Wasserstein loss is achieved.
+One may use, for example, the hamming distance $L(x,y)=\mathbf{1}\{x\neq y\}$ as the ground metric for the discrete data., the temporal Wasserstein imputation (with $p\geq 2$) calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$.
+Hence, $\{a_{1},a_{2},b_{1},b_{2}\}$ satisfies
+[…]
+If $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ would satisfy (13). Naturally, identification is not possible in this case, which leads to our first condition.
+Non-systematic missing pattern: $q_{1}\neq q_{2}$.
+Condition (C1) precludes the special case of down-sampling, which is notoriously hard to be treated as a missing data problem.
+However, (C1) is inadequate for identification because the equations in (13) are not linearly independent.
+The lack of uniqueness of the solutions to (13) is at the crux of the identification problem in imputation.
+The blue lines in Figure 1 plots the set of solutions to (13).
+Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same.
+Nonetheless, (12) hints an additional requirement for the correct imputation, giving rise to the following condition.
+Interpolator stability: $a_{1}=a_{2}$, $b_{1}=b_{2}$.
+Intuitively, (C2) requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$.
+Together, (C1) and (C2) imply a unique solution to the system (13) and are sufficient for identification.
+Figure 1: Set of calibrations that correspond to temporal Wasserstein imputation solutions (blue line) and the correct calibrations $(a_{1}^{*},a_{2}^{*})$ and $(b_{1}^{*},b_{2}^{*})$ (red dots).
+The dashed orange line is the 45 degree line.
+$q_{1}=3$ and $q_{2}=4$ are used to plot this figure.
+Proposition 5.
+Under (C1) and (C2), the only solution to (13) is
+$a_{1}=a_{2}=a_{1}^{*},\quad b_{1}=b_{2}=b_{1}^{*}.$
+That is, the only 2-dimensional marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
+As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the solution set of temporal Wasserstein imputations at the correct calibrations.
+This example shows, under two natural requirements, TWI can achieve distributional identification through minimizing the Wasserstein distance between the marginal distributions.
+In practice, appropriate tuning of $n_{1}$ and $p$ is needed for TWI to find an imputation with marginal distributions approximating the underlying ones.
+As shown in Proposition 5, it is not necessary to use an $n_{1}$ such that all missing data lie on the same side.
+Instead, we can select $n_{1}$ from a pre-specified interval $[t_{1},t_{2}]$ which yields the smallest Wasserstein loss, where $t_{2}-t_{1}$, $t_{2}$, and $n-t_{1}$ are sufficiently large.
+Alternatively, a simple procedure to promote interpolator stability, as suggested by (C2), is to run TWI with different cut-off $n_{1}$, using previously obtained imputations as initializations.
+This procedure, which we call $k$-TWI, is summarized in Algorithm 2.
+Finally, as a nonparametric method, $p$ can be chosen to diverge with the sample size with $p/n\rightarrow 0$.
+Input: Cut-off points $n_{1},n_{2},\ldots,n_{k}$, number of lags $p$
+Initialization: $\hat{\mathbf{w}}^{(0)}\in\mathcal{C}$
+1 for $h=1,2,\ldots,k$ do
+2      Run Algorithm 1 with initialization $\hat{\mathbf{w}}^{(0)}$ and cut-off $n_{h}$ to obtain $\hat{\mathbf{w}}$
+3      Set $\hat{\mathbf{w}}^{(0)}=\hat{\mathbf{w}}$
+4      end for
+     Output: Imputed time series $\hat{\mathbf{w}}$
+Algorithm 2 $k$-TWI
+5 Simulation Studies
+In this section, we apply TWI to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance.
+For comparison, we employ imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations.
+In addition, the following two missing patterns are considered.
+  • Pattern I: 300 observations are omitted at random;
+  • Pattern II: For every 20 observations, 6 consecutive observations are omitted.
+Pattern I simulates sporadic missing data while Pattern II generates patches of missing data.
+The total length of the time series is set to $n=1000$.
+Thus, for both missing patterns, 30% of the data are missing.
+For each data generating process and missing pattern considered, 1000 Monte Carlo simulations are carried out.
+We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward.
+Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (6).
+Throughout we fix $n_{1}=0.4n$ and $p=6$ for TWI.
+In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($n_{1}\in\{0.25n,0.5n,0.75n\}$) to the simulated data.
+For the initializations of TWI, both linear interpolation and Kalman smoothing are used in our experiments.
+Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R.
+The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series.
+Its implementation details are documented in Section S2 of the supplementary material.
+5.1 Linear and nonlinear univariate time series
+In this subsection, we consider the following data-generating processes (DGPs) for univariate time series, which can exhibit linear or nonlinear dynamic patterns.
+In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d.
+sequences of standard Gaussian variables.
+DGP 1 (AR).
+$x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\phi$ is set to 0.8.
+DGP 2 (ARMA).
+$(1-\phi_{1}B)x_{t}=(1+\phi_{2}B)\epsilon_{t}$, where $B$ is the backshift operator and $(\phi_{1},\phi_{2})=(0.8,-0.6)$.
+DGP 3 (TAR; Tong,, 1983).
+$x_{t}=\left\{\begin{array}[]{cc}\phi_{1}x_{t-1}+\epsilon_{t},&x_{t-1}\leq\tau\\ \phi_{2}x_{t-1}+0.5\epsilon_{t},&x_{t-1}>\tau\end{array}\right.,$
+where $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$
+DGP 4 (I(1) process).
+Data $\{x_{t}\}$ are generated according to $x_{t}=x_{t-1}+z_{t}+\epsilon_{t,2}$, where $z_{t}=\phi z_{t-1}+0.5\epsilon_{t,1}$ with $\phi=-0.7$.
+DGP 5 (Cyclic series with noise (CYC)).
+$x_{t}=10\cos(t(0.23\pi))+6\cos(t(0.17\pi))+0.5\epsilon_{t}$.
+DGPs 1 and 2 are well-known linear time series models.
+Note that DGP 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution is inadequate to fully characterize the distribution of the underlying process.
+DGP 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018).
+DGP 4 generates an I(1) series which typically produces a highly persistent stochastic trend.
+Finally, DGP 5 features a deterministic cyclic trend with non-standard periods.
+Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
+Three measures of performance are considered.
+First, we report the Wasserstein distance (of order 2) between the empirical 3-dimensional marginal distributions of the imputed series and the full data.
+Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{x_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$.
+We report
+$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{({x}_{t}^{(i)},{x}_{t-1}^{(i)},{x}_{t-2}^{(i)})}\right),$ (14)
+as a measure of how well the imputed series approximates the original series.
+Second, for DGP 14, the model parameters are estimated using the imputed series, and their root mean square errors (RMSEs) are reported.
+Third, we compare the RMSEs of the estimated autocovariance functions (ACFs).
+However, due to space constraints, we relegate the comparison of ACFs to Section S3.1 in the supplementary material.
+The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1.
+Across all five DGPs employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series.
+While other methods may yield good imputations for some specific DGPs, the result is not consistent across all processes.
+For instance, linear and spline interpolations show low Wasserstein loss for DGP 1, but their performance is less competitive when applied to other models.
+In contrast, TWI methods, carefully initialized, consistently deliver solid performance across models and missing patterns.
+For the nonstationary Model 4 (I(1)), we first obtain an imputation $\{\hat{w}_{t}\}$ and consider the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$.
+Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$.
+In this case, TWI consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
+Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often significantly.
+For example, for DGP 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74).
+Similarly, for DGP 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79).
+Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed under DGP 3, alongside $(x_{t},x_{t-1})$ of the original data.
+While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth.
+Note that Kalman smoothing is done via an ARIMA state-space representation; hence the model is misspecified if the underlying process is nonlinear.
+Nevertheless, it already serves as a good initialization of TWI.
+Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations.
+TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization.
+Similar notations are used for $k$-TWI.
+For DGP 4 (I(1)), the loss is computed using the first-differenced series.
+The smallest value for each model and missing pattern is marked in boldface.
+Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
+Missing pattern I
+AR 0.41 0.41 0.42 0.50 0.40 0.44 0.41 0.44
+ARMA 0.48 0.58 0.47 0.48 0.40 0.38 0.40 0.38
+TAR 1.12 1.51 1.38 1.13 0.96 0.81 0.74 0.63
+I(1) 0.71 0.75 0.62 4.38 0.58 0.53 0.54 0.50
+CYC 1.96 0.82 0.77 1.97 0.79 0.77 0.60 0.70
+NLVAR 2.98 3.42 3.01 2.94 2.25 2.14 2.19 2.12
+AL (×$\times 10$) 0.70 1.00 4.17 0.55 0.43 0.37 0.38 0.35
+Missing pattern II
+AR 0.44 0.59 0.50 0.76 0.39 0.44 0.43 0.45
+ARMA 0.47 1.29 0.51 0.53 0.36 0.34 0.40 0.35
+TAR 1.04 3.11 1.25 0.99 0.84 0.73 0.76 0.61
+I(1) 0.67 0.65 0.60 8.35 0.51 0.45 0.49 0.44
+CYC 2.58 1.81 0.77 3.79 2.62 1.60 0.62 0.72
+NLVAR 3.00 6.07 3.33 3.67 2.13 2.02 2.31 2.04
+AL (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
+(a) Under Missing pattern I
+(b) Under Missing pattern II
+Figure 2: Scatter plot of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
+Next, we turn to the model parameters estimated from the imputed series.
+Table 2 presents the RMSEs for each model parameter.
+For DGP 1 (AR) and DGP 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements.
+For the nonlinear DGP 3 (TAR), the results are drastically different.
+TWI produces the most accurate parameter estimates and can achieve more than 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing.
+These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics.
+For DGP 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization.
+Table 2: Estimation root mean square errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
+Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
+Missing pattern I
+AR $\phi$ 0.06 0.06 0.05 0.03 0.03 0.04 0.03 0.04
+ARMA $\phi_{1}$ 0.19 0.47 0.12 0.08 0.17 0.16 0.11 0.11
+$\phi_{2}$ 0.46 0.95 0.18 0.10 0.33 0.26 0.16 0.15
+TAR $\phi_{1}$ 1.02 1.67 0.82 0.63 0.71 0.58 0.53 0.44
+$\phi_{2}$ 0.08 0.07 0.20 0.04 0.05 0.03 0.03 0.02
+$\tau$ 0.01 0.75 0.35 0.20 0.05 0.05 0.03 0.03
+I(1) $\phi$ 0.96 0.47 0.45 0.68 0.48 0.32 0.30 0.24
+NLVAR $\phi_{11}$ 0.19 0.27 0.15 0.03 0.06 0.03 0.05 0.03
+$\phi_{12}$ 3.10 3.16 2.87 1.67 1.34 1.02 1.22 0.97
+$\phi_{21}$ 0.03 0.07 0.03 0.03 0.03 0.03 0.03 0.03
+$\phi_{21}$ 0.18 0.23 0.10 0.04 0.05 0.04 0.04 0.04
+Missing pattern II
+AR $\phi$ 0.05 0.07 0.04 0.04 0.03 0.02 0.03 0.02
+ARMA $\phi_{1}$ 0.04 0.08 0.09 0.08 0.05 0.06 0.09 0.09
+$\phi_{2}$ 0.16 0.61 0.11 0.10 0.12 0.11 0.11 0.12
+TAR $\phi_{1}$ 0.72 2.44 0.38 0.25 0.25 0.20 0.29 0.21
+$\phi_{2}$ 0.09 0.12 0.05 0.03 0.06 0.05 0.02 0.01
+$\tau$ 0.01 1.72 0.01 0.01 0.01 0.01 0.01 0.01
+I(1) $\phi$ 0.08 1.08 0.11 0.70 0.09 0.10 0.10 0.10
+NLVAR $\phi_{11}$ 0.19 0.44 0.08 0.02 0.02 0.01 0.02 0.01
+$\phi_{12}$ 2.67 3.51 1.68 1.26 0.61 0.46 0.61 0.43
+$\phi_{21}$ 0.02 0.09 0.03 0.03 0.06 0.05 0.04 0.03
+$\phi_{21}$ 0.15 0.34 0.05 0.04 0.04 0.04 0.04 0.04
+5.2 Multivariate time series
+In this subsection, we consider multivariate nonlinear time series models.
+DGP 6 (Nonlinear VAR; NLVAR).
+$x_{t,1}=\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}\quad x_{t,2}=\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
+where $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ is the centered sigmoid function, and $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
+DGP 6 is a VAR model with a nonlinear term.
+In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function.
+As depicted in Figure 3, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships.
+This is validated by the results shown in Table 1, where TWI achieves the lowest Wasserstein distance.
+Next, we assess the effectiveness of the imputed series in estimating the model parameters $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$, assuming the nonlinear functional form is known.
+The root mean square estimation errors are documented in Table 2.
+Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$.
+It attains more than 50% reduction in RMSE compared to the scalar filter, the best-performing benchmark.
+We also consider a challenging multivariate nonlinear model, which generates the so-called compositional time series.
+The data are generated from the additive logistic (AL) model of Brunsdon and Smith, (1998).
+Due to space constraints, we present its simulation results and discussion in Section S3.2 of the supplementary material.
+(a) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,1}$
+(b) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,1}$
+(c) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,2}$
+(d) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,2}$
+Figure 3: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 under missing pattern II.
+6 Groundwater data application
+In this section, we apply the proposed TWI to a groundwater dataset from Taiwan, which was collected and analyzed by Hsu et al., (2020).
+Given the island’s steep topography and varying rainfall distribution, groundwater management is of considerable importance for Taiwan, particularly as it is a critical water source.
+However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells.
+We aim to study the seasonal groundwater variations, for which we first transform the hourly data from 352 sites between October 1992 and August 2020 into monthly records.
+For each site, monthly groundwater level is calculated as the average of the observations in a given month, with months classified as missing if more than one-third of the observations are missing.
+Our analysis focuses on sites with a less-than-median amount of missing data, reducing the sample to 176 sites.
+Figure 4 shows the monthly groundwater levels for these sites, after subtracting the mean and normalizing by the standard deviation.
+We make two observations.
+First, many series exhibit strong seasonality due to the rainfall pattern in Taiwan.
+Monsoon and typhoons increase rainfall in summer, while winters are drier.
+Second, there are many outlying values and anomalies in the observed series, including abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean.
+We also treat such extreme values as missing data.
+Figure 4: Time plot of the groundwater levels of the 176 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
+During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites.
+Figure 5 shows the fluctuations in the percentage of sites with missing data over time.
+Although missing data are mostly prevalent in the early years, there are some time during which the percentage surges, such as between 2003 and 2005.
+In addition, the percentage is gradually increasing in more recent years.
+Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
+We consider two downstream tasks using the imputed series.
+First, following Hsu et al., (2020), we decompose the de-trended time series using the independent component analysis (ICA), for which we apply the fastICA algorithm of Hyvärinen and Oja, (2000).
+Because direct application of ICA becomes infeasible when the series are severely corrupted by missing data, the analysis of Hsu et al., (2020) is confined to a relatively short time span.
+The second downstream task considers autoregressive modeling, which is presented in Section S3.3 in the supplementary material.
+In implementation, we chose $n_{1}=160$ and $p=6$ for TWI, and for $k$-TWI we use $n_{1}\in\{160,225,110\}$.
+Figure 5: Monthly percentage of sites recording a missing value.
+Figure 6 shows the leading two independent components (ICs) estimated using the imputed series.
+ICs estimated from series imputed by Kalman smoothing have peaks around August-September (red shaded area) and May-June (blue shaded area), which was also observed in Hsu et al., (2020).
+These peaks are attributed to the dual rain seasons in Taiwan, which involve typhoons in August-September and the Meiyu period from mid-May to mid-June.
+ICs estimated from TWI-imputed series with Kalman initialization largely preserve these seasonal peaks, but with improved performance before 2002.
+Recall that missing data are prevalent during this period (Figure 5).
+Especially for the first IC, those computed from TWI-imputed series have clear peaks around August-September with a more consistent amplitude.
+On the other hand, the ICs computed from the linearly interpolated series could not effectively isolate and identify the Meiyu season, with peaks occurring slightly earlier than expected.
+Nonetheless, ICs estimated from TWI-imputed series with linear interpolation substantially alleviates this shortcoming and produced more interpretable results.
+ICs computed from spline-interpolated series also identifies the two rain seasons, albeit they show more rugged trajectories instead of clean periodic patterns.
+For the scalar filter, the resulting ICs similarly appear noisy, and the Meiyu season is not well-isolated, showing that the imputation induces distortions in the dynamic pattern.
+Figure 6: Independent components estimated using the imputed series.
+The dashed red lines in each plot mark January 2002.
+The red and blue shaded areas correspond to August-September (typhoon season) and May-June (Meiyu season), respectively.
+Overall, TWI took a more holistic view which accounts for the underlying dynamics, such as complex seasonality, and produced interpretable results.
+In Section S3.3 of the supplementary material, we conduct further experiments to show that the marginal distributions of TWI-imputed series, in terms of AR modeling, are also more consistent across time compared to other methods.
+These results show TWI performs reliably in practice and is a useful method in handling missing data in time series.
+SUPPLEMENTARY MATERIAL
+Supplementary material:
+The file contains the proofs of the propositions presented in Section 4 and some technical results.
+It also documents the scalar filter algorithm of Peña and Tsay, (2021) and complementary numerical results, including the simulation results for autocovariance estimation and for data generated from a compositional time series model.
+Finally, it includes the analysis of the groundwater data via autoregressive modeling.
+Supplementary material to “Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series”
+This supplementary material contains three sections.
+Section S1 collects the proofs of the propositions presented in main text as well as some technical results.
+Section S2 documents the scalar filter algorithm of Peña and Tsay, (2021) for completeness.
+Section S3 contains complementary numerical results, including the simulation results for autocovariance estimation and imputing data generated from a compositional time series model.
+It also includes the analysis of autoregressive modeling using the imputed groundwater data.
+S1 Technical proofs
+Proof of Proposition 3.
 Clearly, (a) holds by construction of the algorithm.
@@
 Moreover,
-$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}^{2}}=$ $\mathrm{diag}\left(\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):i=p-1,\ldots,n_{1}\right),$ (10)
-$\frac{\partial^{2}F_{0}}{\partial\mathbf{v}^{2}}=$ $\mathrm{diag}\left(\sum_{i=p-1}^{n_{1}}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):j=n_{1}+1,\ldots,n-1\right),$ (11)
-$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}\partial\mathbf{v}}=$ $-\left(\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j})\right)_{p-1\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}.$ (12)
+$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}^{2}}=$ $\mathrm{diag}\left(\sum_{j=n_{1}+1}^{n-1}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):i=p-1,\ldots,n_{1}\right),$ (S1.1)
+$\frac{\partial^{2}F_{0}}{\partial\mathbf{v}^{2}}=$ $\mathrm{diag}\left(\sum_{i=p-1}^{n_{1}}\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j}):j=n_{1}+1,\ldots,n-1\right),$ (S1.2)
+$\frac{\partial^{2}F_{0}}{\partial\mathbf{u}\partial\mathbf{v}}=$ $-\left(\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j})\right)_{p-1\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}.$ (S1.3)
 Since $\ell$ is convex, ′′$\ell^{\prime\prime}\geq 0$.
-From (10)–(12), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries.
+From (S1.1)–(S1.3), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries.
 It follows from the Gershgorin circle theorem that $\mathbf{H}$ is positive semi-definite.
@@
 ∎
-Remark 5.
-It is straightforward to generalize Proposition 4 for multivariate time series. Details are omitted.
-Due to the biconvexity of $F$, one can employ many off-the-shelf solvers for the subproblems in Algorithm 1.
-Moreover, it follows directly from the proof that if $\lambda>0$, $F(\mathbf{w},\mathbf{\Pi})$ is $\lambda$-strongly convex in $\mathbf{w}$.
-Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
-In general, not all critical points of $F$ yield desirable imputations, and even its global minimizer may not correspond to the correct marginal distributions.
-For (an extreme) example, imputing a fixed constant when all data are missing results in a zero transport cost, which is clearly arbitrary.
-However, when the missing data happen to lie on the same side of $n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 6 below.
-Before stating the proposition, we make the following assumptions, which are satisfied by many linear and nonlinear time series models.
+Before proving Proposition 4, we first state our assumptions and establish Lemma S6.
 The underlying time series $\{\tilde{x}_{t}\}$ is stationary and ergodic.
 $\mathbb{E}|\tilde{x}_{t}|^{k}<\infty$ for some $k\geq 1$.
-Proposition 6.
-Assume (A1) and (A2) hold.
-Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$.
-Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$.
-If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$, subject to relevant constraints such as the coupling conditions and with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, satisfies
-$\mathcal{W}_{k}\left(\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)+\mathcal{W}_{k}\left(\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)\rightarrow 0\quad\mbox{almost surely},$
-where $\mu_{p}$ is the $p$-dimensional marginal distribution of $\{\tilde{x}_{t}\}$.
-Proposition 6 implies the empirical marginal distributions of the imputed time series converge to the underlying marginal distribution in the Wasserstein distance. This implies
-$\hat{\mu}_{\mathrm{pre}}(\hat{\mathbf{w}})=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}\Rightarrow\mu_{p},\mbox{ and}\quad\hat{\mu}_{\mathrm{post}}(\hat{\mathbf{w}})=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}\Rightarrow\mu_{p}.$
-In fact, convergence in Wasserstein distance further ensures for any continuous function $\varphi$ with $|\varphi(x)|\leq C(1+|x|^{k})$ for all $x\in\mathbb{R}$ and some $C<\infty$, we have
-[…]
-[…]
-In practice, we may not have all missing data lying on the same side of $n_{1}$, even though we have some freedom to choose $n_{1}$.
-Unfortunately, asymptotic consistency results for general missing pattern that are in similar spirits as Proposition 6 can hardly be derived.
-The reason, as we shall see in the next subsection, is that matching the underlying marginal distribution by equating the empirical marginal distributions essentially suffers from an identification problem.
-4.2 General missing pattern: A two-state Markov chain example
-In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example.
-Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations.
-We also show that, under two favorable conditions, the correct marginal distributions can still be pinned down by TWI.
-To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
-Suppose the underlying time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain. Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by
-$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.2,$
-$\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.6.$
-The transition probabilities are specified to facilitate discussions; the main results below are not specific to these choices.
-It is easy to check that $\{\tilde{x}_{t}\}$ has the following stationary distribution.
-$\mathbb{P}(\tilde{x}_{t}=1)=0.75,\quad\mathbb{P}(\tilde{x}_{t}=0)=0.25.$
-Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
-$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=0.6,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=0.15,$ (13)
-$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=0.15,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=0.1.$
-Without loss of generality, we set $n_{1}=0$.
-Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations.
-Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$.
-Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$, $q_{2}$ are some integers larger than 2.
-Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ (14)
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$
-and if $t\geq n_{1}$,
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ (15)
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$
-for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$.
-Note that $\mathbb{P}$ in (14) and (15) should be interpreted as proportions since the imputation may still depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined.
-Imputation methods corresponding to varying $a_{1},a_{2},b_{1},b_{2}$ will lead to different 2-dimensional marginal distributions.
-Indeed, from (13)–(15), we can compute the 2-dimensional marginal distributions of the imputed series.
-For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-q_{1}^{-1})+\frac{3}{4q_{1}}a_{1},$
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))$ $=0.15(1-q_{1}^{-1})+\frac{1}{4q_{1}}b_{1},$
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))$ $=0.15(1-q_{1}^{-1})+\frac{3}{4q_{1}}(1-a_{1}),$
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))$ $=0.1(1-q_{1}^{-1})+\frac{1}{4q_{1}}(1-b_{1}).$
-For $\{\hat{w}_{t}\}_{0\leq t<\infty}$, we can obtain similar expressions with $a_{1},b_{1},q_{1}$ replaced by $a_{2},b_{2},q_{2}$.
-The “correct” imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting 2-dimensional marginal distributions equal the marginal distribution of the underlying series, given in (13).
-To achieve such identification, the correct imputation must choose
-$a_{1}^{*}=a_{2}^{*}=0.8,\quad b_{1}^{*}=b_{2}^{*}=0.6.$ (16)
-Note that this calibration does not depend on $q_{1},q_{2}$; it only depends on the distributional properties of the underlying process.
-Assuming the optimization is carried out perfectly [^1] This is where we ignore finite sample complications and assume a zero Wasserstein loss is achieved.
-One may use, for example, the hamming distance $L(x,y)=\mathbf{1}\{x\neq y\}$ as the ground metric for the discrete data., the temporal Wasserstein imputation (with $p\geq 2$) calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$.
-Hence, $\{a_{1},a_{2},b_{1},b_{2}\}$ satisfies
-[…]
-Clearly, if $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ satisfies the above system of equations.
-Therefore, identification is not possible in this case, which leads to our first condition.
-Asymmetric missing pattern: $q_{1}\neq q_{2}$.
-However, condition (C1) is still inadequate for identification because the equations in (17) are not linearly independent.
-The lack of uniqueness of the solutions to (17) is at the crux of the identification issue of the imputation problem.
-The blue lines in Figure 1 plots the set of solutions to (17).
-Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same.
-Figure 1: Set of calibrations that correspond to temporal Wasserstein imputation solutions (blue line) and the correct calibrations $(a_{1}^{*},a_{2}^{*})$ and $(b_{1}^{*},b_{2}^{*})$ (red dots).
-The dashed orange line is the 45 degree line.
-$q_{1}=3$ and $q_{2}=4$ are used to plot this figure.
-Nonetheless, (16) hints an additional requirement for the correct imputation, giving rise to the following condition.
-Interpolator stability: $a_{1}=a_{2}$, $b_{1}=b_{2}$.
-Intuitively, (C2) requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$.
-Together, (C1) and (C2) imply a unique solution to the system (17) and are sufficient for identification.
-Proposition 7.
-Under (C1) and (C2), the only solution to (17) is
-$a_{1}=a_{2}=a_{1}^{*},\quad b_{1}=b_{2}=b_{1}^{*}.$
-That is, the only marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
-As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the set of temporal Wasserstein imputations at the correct calibrations.
-The above example illustrates when TWI can potentially identify the underlying marginal distribution.
-The crucial interpolator stability condition (C2) thus provides some guidance on the selection of TWI solutions.
-Indeed, Algorithm 1 may converge to varying solutions depending on the initialization.
-(C2) implies that we shall choose the solution so that the relationship between the imputed values and the neighboring observations stays stable before and after $n_{1}$.
-In higher dimensions, summary statistics can be used.
-For example, fit a linear regression of imputed values on neighboring values using data before and after $n_{1}$ and compare the similarity of the coefficients.
-To promote interpolator stability, we suggest running TWI with different cut-off $n_{1}$, using the previously obtained imputations as initializations.
-This leads to a simple procedure termed $k$-TWI henceforth, which is summarized in Algorithm 2 below.
-Input: Cut-off points $n_{1},n_{2},\ldots,n_{k}$, number of lags $p$
-Initialization: $\hat{\mathbf{w}}^{(0)}\in\mathcal{C}$
-1 for $h=1,2,\ldots,k$ do
-2      Run Algorithm 1 with initialization $\hat{\mathbf{w}}^{(0)}$ and cut-off $n_{h}$ to obtain $\hat{\mathbf{w}}$
-3      Set $\hat{\mathbf{w}}^{(0)}=\hat{\mathbf{w}}$
-4      end for
-     Output: Imputed time series $\hat{\mathbf{w}}$
-Algorithm 2 $k$-TWI
-5 Simulation Studies
-In this section, we apply the proposed temporal Wasserstein imputation to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance.
-For comparison, we employ alternative imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations.
-In addition, the following two missing patterns are considered.
-  • Pattern I: 300 observations are omitted at random, which corresponds to sporadic missing data.
-  • Pattern II: For every 20 observations, 6 consecutive observations are omitted, which mimics patches of missing data.
-The total length of the time series is set to $n=1000$.
-Thus, for both missing patterns, 30% of the data are missing.
-For each model and missing pattern considered in this section, 1000 Monte Carlo simulations are carried out to compute the measure of performance.
-We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward.
-Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (7).
-In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($0.25n$, $0.5n$, and $0.75n$) to the simulated data.
-Both linear interpolation and Kalman smoothing are used as initialization of TWI in our experiments.
-Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R.
-The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series.
-The details are documented in Appendix B.
-5.1 Linear and nonlinear univariate time series
-In this subsection, we focus on univariate time series exhibiting linear or nonlinear dynamic patterns.
-Specifically, we consider the following time series models.
-In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d.
-sequences of standard Gaussian variables.
-Model 1 (AR).
-$x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\phi$ is set to 0.8.
-Model 2 (ARMA).
-$(1-\phi_{1}B)x_{t}=(1+\phi_{2}B)\epsilon_{t}$, where $B$ is the backshift operator and $(\phi_{1},\phi_{2})=(0.8,-0.6)$.
-Model 3 (TAR; Tong,, 1983).
-$x_{t}=\left\{\begin{array}[]{cc}\phi_{1}x_{t-1}+\epsilon_{t},&x_{t-1}\leq\tau\\ \phi_{2}x_{t-1}+0.5\epsilon_{t},&x_{t-1}>\tau\end{array}\right.,$
-where $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$
-Model 4 (I(1) process).
-Data $\{x_{t}\}$ are generated according to
-$z_{t}=$ $\phi z_{t-1}+0.5\epsilon_{t,1},$
-$x_{t}=$ $x_{t-1}+z_{t}+\epsilon_{t,2},$
-with $\phi=-0.7$.
-Model 5 (Deterministic cyclic series with noise (CYC)).
-$x_{t}=10\cos(t(0.23\pi))+6\cos(t(0.17\pi))+0.5\epsilon_{t}.$
-Models 1 and 2 are linear time series models commonly seen in practice.
-Kalman smoothing is particularly adept at imputing such series.
-Note that Model 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution, the key objects TWI and $k$-TWI work with, is inadequate to fully characterize the distribution of the underlying process.
-Model 3 is a celebrated threshold AR model which has found many applications in modeling nonlinear dynamics in the econometrics and statistics literature (Montgomery et al.,, 1998; Li and Ling,, 2012; Tsay and Chen,, 2018).
-Model 4 generates an I(1) series which typically produces a highly persistent stochastic trend.
-To remove the unit root, difference transform is usually applied to the data $\{x_{t}\}$; that is, one often deals with $\nabla x_{t}=x_{t}-x_{t-1}$ instead of the original series.
-However, this may result in excessive missing values in $\{\nabla x_{t}\}$, which can cause difficulty in imputation.
-Finally, Model 5 features a deterministic cyclic trend with non-standard periods.
-Specifically, it includes two sinusoidal components with periods of 8.70 and 11.77, respectively.
-Three measures of performance are considered.
-First, we report the Wasserstein distance (of order 2) between the empirical 3-dimensional marginal distributions of the imputed series and the full data.
-Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{x_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$.
-We report
-$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{({x}_{t}^{(i)},{x}_{t-1}^{(i)},{x}_{t-2}^{(i)})}\right),$ (18)
-as a measure of how well the imputed series approximates the original series.
-Second, we compare the root mean square estimation errors (RMSE) of the estimated autocovariance functions.
+These assumptions are satisfied by many linear and nonlinear time series models.
+Lemma S6.
+Suppose $\{\tilde{x}_{t}\}$ satisfies (A1) and (A2). Then
+$\mathcal{W}_{k}\left(\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})},\mu_{p}\right)\xrightarrow{n\rightarrow\infty}0\quad\mathrm{a.s.},$
+where $\tilde{\mathbf{x}}=(\tilde{x}_{0},\ldots,\tilde{x}_{n-1})$.
+Proof.
+By Theorem 7.12 of Villani, (2021), it suffices to show that with probability one,
+$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}\Rightarrow\mu_{p},$ (S1.4)
+and
+$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|^{k}.$ (S1.5)
+By Theorem 36.4 of Billingsley, (1995), the processes $\{\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\}_{t}$ and $\{\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\}_{t}$ are stationary and ergodic for any $\mathbf{h}\in\mathbb{R}^{p}$, where $\mathrm{i}=\sqrt{-1}$.
+Thus (S1.5) follows immediately from the Ergodic theorem.
+To show (S1.4), it suffices to show, with probability one,
+$\hat{\varphi}_{n}(\mathbf{h}):=\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\rightarrow\int\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v})d\mu_{p}(\mathbf{v})=:\varphi(\mathbf{h})\quad\mbox{for all }\mathbf{h}\in\mathbb{R}^{p}.$
+By the Ergodic theorem, we have
+$\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q}),\quad\mbox{for all }\mathbf{q}\in\mathbb{Q}^{p},$ (S1.6)
+with probability one. Let
+=$\mathcal{E}=\{\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q})\mbox{ for all }\mathbf{q}\in\mathbb{Q}^{p}\}\cap\left\{\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|\right\},$
+noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2).
+By the Ergodic theorem and (S1.6), $\mathbb{P}(\mathcal{E})=1$.
+Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary.
+Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
+$|\varphi(\mathbf{h})-\varphi(\mathbf{q})|<\epsilon,$ (S1.7)
+for all $\mathbf{q}\in\mathbb{Q}^{p}$ with $\|\mathbf{h}-\mathbf{q}\|<\delta$. In addition, using $|\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{z})-1|\leq\|\mathbf{h}\|\|\mathbf{z}\|$, we have
+$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (S1.8)
+for all $\mathbf{h}$, $\mathbf{q}$ such that $\|\mathbf{h}-\mathbf{q}\|<\delta$. Combining (S1.7)–(S1.8), we have
+$|\hat{\varphi}_{n}(\mathbf{h})-\varphi(\mathbf{h})|\leq$ $|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}(\mathbf{q})|+|\hat{\varphi}_{n}(\mathbf{q})-\varphi(\mathbf{q})|+|\varphi(\mathbf{q})-\varphi(\mathbf{h})|$
+$\leq$ $\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|+o(1)+\epsilon$
+on $\omega\in\mathcal{E}$.
+Note, in particular, that $\epsilon$ and $\delta$ do not depend on $\mathbf{h}$.
+Letting $\epsilon\rightarrow 0$ and $\delta\rightarrow 0$ proves (S1.4) holds on $\mathcal{E}$, and hence with probability one.
+∎
+Proof of Proposition 4.
+We only consider the case $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ here as the other case follows from a similar argument. Define
+$\hat{\mu}_{\mathrm{pre}}$ $:=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}=:\mu_{\mathrm{pre}},$
+$\hat{\mu}_{\mathrm{post}}$ $:=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}$
+$\mu_{\mathrm{post}}$ $:=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}$
+Since the Wasserstein distance of order $k$ is a metric on the space of probability measures with finite $k$-th moments, we have
+$\mathcal{W}_{k}(\hat{\mu}_{\mathrm{post}},\mu_{p})\leq$ $\mathcal{W}_{k}(\hat{\mu}_{\mathrm{post}},\hat{\mu}_{\mathrm{pre}})+\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p})$
+$\leq$ $\mathcal{W}_{k}({\mu}_{\mathrm{post}},{\mu}_{\mathrm{pre}})+\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p})$
+$\leq$ $\mathcal{W}_{k}(\mu_{\mathrm{post}},\mu_{p})+2\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p}).$
+Now the desired result follows from Lemma S6. ∎
+S2 Scalar filter algorithm
+The following summarizes the scalar filter algorithm of Peña and Tsay, (2021).
+  • Initialize the imputation $\hat{\mathbf{w}}^{(0)}=(\hat{w}_{0}^{(0)},\hat{w}_{1}^{(0)},\ldots,\hat{w}_{n-1}^{(0)})^{\top}$ with $\hat{w}_{t}^{(0)}=0$ if $t\in\mathcal{M}$.
+  • For each $s\in\mathcal{M}$, let $I^{(s)}_{t}$ be the indicator function. That is,
+    $I_{t}^{(s)}=\left\{\begin{array}[]{cc}1,&t=s\\ 0,&t\neq s\end{array}\right.$
+    and estimate the intervention model using $\hat{\mathbf{w}}^{(0)}$. For example, one can fit an ARX model with $\{I_{t}^{(s)}:s\in\mathcal{M}\}$ as exogenous predictors.
+  • Let $\hat{\beta}^{(s)}$ be the intervention effect associated with $I_{t}^{(s)}$. Estimate the missing data by
+    $\hat{w}_{s}=\hat{w}_{s}^{(0)}-\hat{\beta}^{(s)},\quad s\in\mathcal{M}.$
+As one can see from the algorithm above, the scalar filter algorithm proposed by Peña and Tsay, (2021) is based on intervention analysis.
+It can be iterated in order to improve the estimates.
+However, no convergence guarantee is provided by the authors.
+When fitting an ARX model with the indicators as exogenous predictors, if there are many missing values, standard approaches may be infeasible or unreliable due to near collinearity.
+In this case, we employ a small $\ell_{2}$ penalty to avoid numerical instability.
+S3 Complementary numerical results
+S3.1 Autocovariance estimation
+In this subsection, we compare the ACFs estimated from the imputed series.
 Define $\gamma(\mathbf{z},h)=(n-h)^{-1}\sum_{t=h}^{n-1}(z_{t}-\bar{z})(z_{t-h}-\bar{z})$, where $\mathbf{z}=(z_{0},z_{1},\ldots,z_{n-1})$ and $\bar{z}=n^{-1}\sum_{t=0}^{n-1}z_{t}$.
 Then the RMSE is defined as
-$\mathrm{RMSE}_{\gamma}(h)=\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}(\gamma(\hat{\mathbf{w}}^{(i)},h)-\gamma_{\mathbf{x}}(h))^{2}},$ (19)
+$\mathrm{RMSE}_{\gamma}(h)=\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}(\gamma(\hat{\mathbf{w}}^{(i)},h)-\gamma_{\mathbf{x}}(h))^{2}},$ (S3.9)
 where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(x_{0}^{(i)},\ldots,x_{n-1}^{(i)})$.
-Third, for Models 14, the model parameters are estimated using the imputed series, and their root mean square errors are reported.
-The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1.
-Across all five models employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series.
-While other methods may yield good imputations for some specific models, the result is not consistent across all models.
-For instance, linear and spline interpolations show low Wasserstein loss for Model 1, but their performance is less competitive when applied to other models.
-Similarly, the Kalman smoothing performs well under Model 1 with missing pattern I, but its advantage diminishes under missing pattern II.
-In contrast, TWI methods consistently deliver solid performance across models and missing patterns, provided they are carefully initialized.
-Additionally, Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often siginificantly.
-For example, for Model 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74).
-Similarly, for Model 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79).
-Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed (${x}_{t}=\bigstar$ or ${x}_{t-1}=\bigstar$) under Model 3, alongside $(x_{t},x_{t-1})$ of the original data.
-While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth.
-Figure 3 similarly plots $\hat{w}_{t}$ against $\hat{w}_{t-h}$ (and $x_{t}$ against $x_{t-h}$) for $h=1,2,6,10$ under Model 5.
-As shown in the figure, TWI markedly improves from its linear initialization by “denoising” the marginal distributions, leading to a distribution that closely match the ground truth.
-For the nonstationary Model 4 (I(1)), we apply the methods to obtain an imputation $\{\hat{w}_{t}\}$ and focus on the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$.
-Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$.
-As noted before, there are many more missing values in $\{\nabla x_{t}\}$ than originally in $\{x_{t}\}$.
-In this case, TWI still consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
-Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations.
-TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization.
-Similar notations are used for $k$-TWI.
-For Model 4 (I(1)), the loss is computed using the first-differenced series.
-The smallest value for each model and missing pattern is marked in boldface.
-Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
-Missing pattern I
-AR 0.41 0.41 0.42 0.50 0.40 0.44 0.41 0.44
-ARMA 0.48 0.58 0.47 0.48 0.40 0.38 0.40 0.38
-TAR 1.12 1.51 1.38 1.13 0.96 0.81 0.74 0.63
-I(1) 0.71 0.75 0.62 4.38 0.58 0.53 0.54 0.50
-CYC 1.96 0.82 0.77 1.97 0.79 0.77 0.60 0.70
-NLVAR 2.98 3.42 3.01 2.94 2.25 2.14 2.19 2.12
-ALR (×$\times 10$) 0.70 1.00 4.17 0.55 0.43 0.37 0.38 0.35
-Missing pattern II
-AR 0.44 0.59 0.50 0.76 0.39 0.44 0.43 0.45
-ARMA 0.47 1.29 0.51 0.53 0.36 0.34 0.40 0.35
-TAR 1.04 3.11 1.25 0.99 0.84 0.73 0.76 0.61
-I(1) 0.67 0.65 0.60 8.35 0.51 0.45 0.49 0.44
-CYC 2.58 1.81 0.77 3.79 2.62 1.60 0.62 0.72
-NLVAR 3.00 6.07 3.33 3.67 2.13 2.02 2.31 2.04
-ALR (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
-(a) Marginal distribution under missing pattern I
-(b) Marginal distribution under missing pattern II
-Figure 2: Scatter plot of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
-(a) $h=1$
-(b) $h=2$
-(c) $h=6$
-(d) $h=10$
-Figure 3: Scatter plots of $(x_{t-h},x_{t})$ of the original data and $(\hat{w}_{t-h},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 5 (CYC) under missing pattern I.
-$h\in\{1,2,6,10\}$.
-Next, we examine the sample autocovariance functions (ACFs) estimated from the imputed data.
-Table 2 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$.
+Table S1 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$.
 For Model 1 (AR), most methods perform reasonably well.
@@
 Notably, the scalar filter also yields accurate ACFs for $h=1,2$.
-Figure 4 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
+Figure S1 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
 The advantage of TWI becomes even more pronounced in nonstationary models 4 (I(1)) and 5 (CYC), for which it achieves low ACF estimation errors.
@@
 (b) Missing pattern II
-Figure 4: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR).
+Figure S1: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR).
 The blue line is the smoothed density estimated from the histogram of the ACFs estimated from the ground truth.
-Table 2: $\mathrm{RMSE}_{\gamma}(h)$, defined in (19), for $h=0,1,2$, with $h$ denoting lag.
+Table S1: $\mathrm{RMSE}_{\gamma}(h)$, defined in (S3.9), for $h=0,1,2$, with $h$ denoting lag.
 The results are averaged over 1000 simulations, and the smallest values in each row are marked in boldface.
-See Table 1 for notations.
+See Table 1 in the main text for notations.
 Model $h$ Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@
 2 11.18 5.09 0.79 7.90 1.28 1.04 0.28 0.31
-Finally, we turn to the model parameters estimated from the imputed series.
-Table 3 presents the RMSEs for each model parameter.
-For Model 1 (AR) and Model 2 (ARMA), Kalman smoothing and the scalar filter yield satisfactory estimates among the benchmarks, with TWI offering only limited improvements.
-For the nonlinear Model 3 (TAR), the results are drastically different.
-TWI produces the most accurate parameter estimates and can achieve approximately 50% reduction in estimation errors compared to linear interpolation and Kalman smoothing.
-In some cases, such as with linear initialization under missing pattern II, the improvement far exceeds 50%.
-These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics.
-For Model 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization.
-Figure 5 shows the histogram of the estimated parameter $\phi_{1}$ in Model 4.
-It is visible from the figure that TWI helps neutralize the biases from the initialization and produces reliable estimates.
-Table 3: Estimation root mean square errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
-Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
-Missing pattern I
-AR $\phi$ 0.06 0.06 0.05 0.03 0.03 0.04 0.03 0.04
-ARMA $\phi_{1}$ 0.19 0.47 0.12 0.08 0.17 0.16 0.11 0.11
-$\phi_{2}$ 0.46 0.95 0.18 0.10 0.33 0.26 0.16 0.15
-TAR $\phi_{1}$ 1.02 1.67 0.82 0.63 0.71 0.58 0.53 0.44
-$\phi_{2}$ 0.08 0.07 0.20 0.04 0.05 0.03 0.03 0.02
-$\tau$ 0.01 0.75 0.35 0.20 0.05 0.05 0.03 0.03
-I(1) $\phi$ 0.96 0.47 0.45 0.68 0.48 0.32 0.30 0.24
-NLVAR $\phi_{11}$ 0.19 0.27 0.15 0.03 0.06 0.03 0.05 0.03
-$\phi_{12}$ 3.10 3.16 2.87 1.67 1.34 1.02 1.22 0.97
-$\phi_{21}$ 0.03 0.07 0.03 0.03 0.03 0.03 0.03 0.03
-$\phi_{21}$ 0.18 0.23 0.10 0.04 0.05 0.04 0.04 0.04
-Missing pattern II
-AR $\phi$ 0.05 0.07 0.04 0.04 0.03 0.02 0.03 0.02
-ARMA $\phi_{1}$ 0.04 0.08 0.09 0.08 0.05 0.06 0.09 0.09
-$\phi_{2}$ 0.16 0.61 0.11 0.10 0.12 0.11 0.11 0.12
-TAR $\phi_{1}$ 0.72 2.44 0.38 0.25 0.25 0.20 0.29 0.21
-$\phi_{2}$ 0.09 0.12 0.05 0.03 0.06 0.05 0.02 0.01
-$\tau$ 0.01 1.72 0.01 0.01 0.01 0.01 0.01 0.01
-I(1) $\phi$ 0.08 1.08 0.11 0.70 0.09 0.10 0.10 0.10
-NLVAR $\phi_{11}$ 0.19 0.44 0.08 0.02 0.02 0.01 0.02 0.01
-$\phi_{12}$ 2.67 3.51 1.68 1.26 0.61 0.46 0.61 0.43
-$\phi_{21}$ 0.02 0.09 0.03 0.03 0.06 0.05 0.04 0.03
-$\phi_{21}$ 0.15 0.34 0.05 0.04 0.04 0.04 0.04 0.04
-(a) missing pattern I
-(b) missing pattern II
-Figure 5: Histogram of estimated model parameter $\phi_{1}$ when the data are generated from Model 4 (I(1)).
-The blue line is the smoothed density estimated from the histogram of the parameter estimated from the ground truth.
-In conclusion, under the univariate time series models considered in this subsection, which include linear, nonlinear, and nonstationary time series, TWI consistently fares well in all three comparing metrics.
-It can capture nonlinear dynamics as illustrated by the low Wasserstein error in the imputed marginal distributions, which provides substantial improvements in downstream ACF and model estimates.
-5.2 Multivariate time series
-In this subsection, we consider the following two multivariate nonlinear time series models.
-Model 6 (Nonlinear VAR; NLVAR).
-$x_{t,1}=$ $\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}$
-$x_{t,2}=$ $\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
-where $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ is the centered sigmoid function, and $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
-Model 7 (Additive logistic model; AL, Brunsdon and Smith,, 1998).
-$\begin{pmatrix}y_{t,1}\\ y_{t,2}\end{pmatrix}=\begin{pmatrix}0.1\\ 0.1\end{pmatrix}+\begin{pmatrix}0.7&-0.5\\ 0&-0.7\end{pmatrix}\begin{pmatrix}y_{t-1,1}\\ y_{t-1,2}\end{pmatrix}+0.2\begin{pmatrix}\epsilon_{t,1}\\ \epsilon_{t,2}\end{pmatrix},$
+S3.2 Compositional time series
+In this subsection, we discuss the simulation results associated with the following model.
+DGP 7 (Additive logistic model; AL, Brunsdon and Smith,, 1998).
+$y_{t,1}=$ $0.1+0.7y_{t-1,1}-0.5y_{t-1,2}+0.2\epsilon_{t,1}$
+$y_{t,2}=$ $0.1-0.7y_{t-1,2}+0.2\epsilon_{t,2}$
 and the observed data are
 $x_{t,j}=\left\{\begin{aligned} &\frac{\exp(y_{t,j})}{1+\exp(y_{t,1})+\exp(y_{t,2})},\quad j=1,2,\\ &\frac{1}{1+\exp(y_{t,1})+\exp(y_{t,2})},\quad j=3.\end{aligned}\right.$
-Model 6 is a VAR model with a nonlinear term.
-In particular, the impact of $x_{t-1,2}$ on $x_{t,1}$ is nonlinear and is modeled by a sigmoid function.
-Model 7 generates the so-called compositional time series, which refers to a multivariate time series $\{\mathbf{x}_{t}\}$, where $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$, $d>1$, such that for each $t$, $\sum_{j=1}^{d}x_{t,j}=1$, and $x_{t,j}\geq 0$ for all $j$.
+DGP 7 generates the so-called compositional time series, which refers to a multivariate time series $\mathbf{x}_{t}=(x_{t,1},\ldots,x_{t,d})^{\top}$, $d>1$, such that for each $t$, $\sum_{j=1}^{d}x_{t,j}=1$, and $x_{t,j}\geq 0$ for all $j$.
 Such data are encountered in a wide array of fields such as the environmental sciences, economics, geology and political science.
-Here, we adopt the additive logistic (AL) model of Brunsdon and Smith, (1998) in Model 7.
+Here, we adopt the additive logistic (AL) model of Brunsdon and Smith, (1998) in DGP 7.
 The AL model postulates that after some nonlinear transformation, the data follow a standard VARMA model.
-In fact, one can view compositional data as a time series on some low-dimensional manifold.
 Alternative models for composition time series include the recent works of Zheng and Chen, (2017), Harvey et al., (2024), and Zhu and Müller, (2024).
-Focus on Model 6 (NLVAR) first.
-The Wasserstein distances between the 3-dimensional marginal distributions of the imputed and the original series, defined in (18), are again reported Table 1.
-TWI methods clearly produce the lowest Wasserstein distances.
-As depicted in Figure 6, TWI successfully captures the underlying nonlinear dynamics, whereas other imputation methods fail to preserve meaningful relationships.
-Next, we assess the effectiveness of the imputed series in estimating the model parameters, assuming the nonlinearity is known.
-That is, we estimate $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$ by plugging in the imputed series.
-The root mean square estimation errors are documented in Table 3.
-Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$.
-It attains more than 50% reduction in RMSE from initialization under both missing patterns I and II, and can achieve 50% lower RMSE than the scalar filter, the best-performing benchmark.
-(a) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,1}$
-(b) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,1}$
-(c) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,2}$
-(d) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,2}$
-Figure 6: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 (NLVAR) under missing pattern II.
-Next, we turn to Model 7 (AL).
-Since the components must sum to unity, they exhibit negative correlation when one component approaches the endpoints 0 and 1.
-However, such correlation tends to weaken when the series are near the midpoint.
-Table 1 records the Wasserstein distance between the 3-dimensional marginal distribution of the imputed and original series.
+Table 1 in the main paper records the Wasserstein distance between the 3-dimensional marginal distributions of the imputed and original series.
 With the smallest Wasserstein loss, TWI methods again demonstrate their ability to adapt to nonlinear dependence structures.
@@
 By restricting the admissible set, TWI ensures that the imputed series remain compositional.
-As shown in Figure 7, linear interpolation falsely promotes negative correlation between $\hat{w}_{t,3}$ and $\hat{w}_{t-1,2}$ and Kalman smoothing often deviates from the plausible range of a compositional series.
+As shown in Figure S2, linear interpolation falsely promotes negative correlation between $\hat{w}_{t,3}$ and $\hat{w}_{t-1,2}$ and Kalman smoothing often deviates from the plausible range of a compositional series.
 TWI amends these peculiarities, leading to imputations that align more closely with the underlying data.
-Figure 7: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data (plotted in blue) and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series (plotted in red) for $i,j=1,2,3$, $i\neq j$, when the data are generated from Model 7 (AL) under missing pattern I.
-6 Groundwater data application
-In this section, we apply the proposed temporal Wasserstein imputation to a groundwater dataset from Taiwan.
-Groundwater serves as a critical water source for Taiwan, and its management is of considerable importance to the policymakers as well as the industrial sector, particularly due to the island’s steep topography and varying rainfall distribution (Hsu et al.,, 2020).
-However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells.
-To address this, imputation of missing values is essential to facilitate reliable downstream analyses.
-6.1 Preprocessing and exploratory analysis
-Our goal is to study the seasonal groundwater variations, for which we first transform the data into monthly records.
-The data contain hourly observations from 352 monitoring sites between October, 1992 and August, 2020.
-For each site, we define monthly groundwater level by the average of the observations in a given month.
-If more than one-third of the observations for a month are missing, we classify that month as missing.
-For our analysis, we focus on sites with a less-than-median amount of missing data, reducing the sample to 176 sites.
-Figure 8 shows the monthly groundwater levels for these sites, after subtracting the mean and normalized by the standard deviation.
-We make two observations.
-First, many series exhibit strong seasonality, which is due to the rainfall pattern in Taiwan.
-There is more rain in the summer due to monsoon and typhoons, while winters are generally drier.
-Second, there are many outlying values and anomalies in the observed series.
-For instance, there are abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean.
-For the purpose of our analysis, we treat such extreme values as missing data.
-Figure 8: Time plot of the groundwater levels of the 176 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
-During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites.
-Figure 9 shows the fluctuations in the percentage of sites with missing data over time.
-Although missing data are mostly prevalent in the early years with more than 50% of the sites generating missing data, there are some time during which the percentage surges, such as between 2003 and 2005.
-In addition, the percentage of missing data is gradually increasing in more recent years.
-Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
-Figure 9: Monthly percentage of sites recording a missing value.
-6.2 Results
-Similar to Section 5, we employ linear interpolation, spline interpolation, Kalman smoothing, and the scalar filter as benchmark imputation methods.
-For TWI and $k$-TWI, linear interpolation and Kalman smoothing are again used as initializations.
-We consider two downstream tasks using the imputed series.
-First, we perform dimension reduction via the time series factor model of Lam et al., (2011).
-The factor model postulates that the observations $\mathbf{x}_{t}\in\mathbb{R}^{176}$ are driven by a small number of latent factor processes.
-Specifically,
-$\mathbf{x}_{t}=\mathbf{A}\mathbf{f}_{t}+\bm{\epsilon}_{t},$
-where $\mathbf{f}_{t}\in\mathbb{R}^{r}$, $r\ll 176$, is the factor process and $\{\bm{\epsilon}_{t}\}$ is white noise.
-Factor modeling is helpful for understanding the high-dimensional time series and its complex temporal patterns.
-However, estimating the factor model requires eigenanalysis of the autocovariance matrices $\mathbf{\Sigma}(h)=\mathbb{E}(\mathbf{x}_{t}\mathbf{x}_{t-h})$, which becomes infeasible when the series are severely corrupted by missing data.
-Figure 10 shows the leading two factors estimated using series imputed by the benchmark methods.
-The linear and spline interpolation methods, both favoring trend preservation, result in the first estimated factor displaying strong trends, while the second factor reflects more the seasonal patterns.
-Notably, this is reversed when factors are estimated based on series imputed by Kalman smoothing and scalar filter.
-In these cases, the leading factor process exhibits strong seasonality, and the second factor displays trend patterns with a slight U-shape, which is quite different from those estimated with linear and spline methods.
-Figure 10: Factor processes estimated using series imputed by the benchmark methods. The dashed red line marks January 2002.
-Figure 11 displays the leading two factors estimated using series imputed by TWI and $k$-TWI.
-Note that while linear interpolation has led to a strong trend factor, as shown in Figure 10, the leading estimated factor of TWI initialized by linear interpolations exhibits significant seasonal pattern, which aligns more closely with those estimated by Kalman smoothing and scalar filter.
-Moreover, the second estimated factor also shows trends resembling those of the Kalman smoothing and scalar filter, but with a more apparent U-shape.
-To compare TWI with Kalman smoothing, we focus on the data before January 2002 (marked by the red dashed lines in Figure 10 and 11).
-Recall that missing data are prevalent during this period (Figure 9).
-Factors estimated by Kalman smoothing are affected by the abundance of missing data and tend to show smaller variations in the early periods, which may be the result of small variances in the optimal interpolator.
-In contrast, factors estimated using data imputed by TWI (initialized by Kalman smoothing) show consistent variational patterns across time and appear more trustworthy in understanding the dynamics.
-Overall, TWI took a more holistic view of the series which accounts for the prominent seasonality and trend in the data.
-Figure 11: Factor processes estimated using series imputed by TWI. The dashed red line marks January 2002.
-Our second downstream task concerns autoregreesive modeling.
-Specifically, we fit an AR(6) model using data imputed by the methods considered.
-The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including missing values.
+Figure S2: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data (plotted in blue) and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series (plotted in red) for $i,j=1,2,3$, $i\neq j$, when the data are generated from Model 7 (AL) under missing pattern I.
+S3.3 Autoregressive modeling with groundwater data
+As our second downstream task for the groundwater data analysis, we fit an AR(6) model using data imputed by the methods considered.
+The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including imputed values.
 Then, we assess the impact of imputation quality by the out-of-sample prediction performance in predicting the remaining data.
-Note that the missing data in the test data are removed for fair comparison.
-Figure 12 presents, in violin plots, the distribution of the root mean squared prediction errors (RMSE) for each method across the 176 series under study.
-The TWI-type methods effectively alleviate the large RMSEs caused by the benchmark imputation methods and produce lower RMSEs in general.
-Figure 12: The root mean squared prediction errors (RMSEs) distribution for each method across the 176 series, ranked by median from left to right.
-In conclusion, the temporal Wasserstein imputation preserves important temporal features such as the seasonal patterns and enables complex downstream statistical analyses such as factor modeling.
-Our analysis show that the results obtained from TWI appear reliable.
-Thus, we expect the series imputed by TWI are also suitable for further analyses, especially when nonlinear models are employed.
-7 Conclusion
-In this work, we proposed an imputation method, the temporal Wasserstein imputation, for imputing missing entries in time series data.
-The proposed method complements the existing approaches by effectively capturing nonlinear dynamics and reducing the distributional bias typically observed for the conventional methods in handling missing time series data.
-Hence, the imputation of the proposed method is more reliable in downstream analysis.
-Our simulation studies and the analysis of groundwater data support the practical utility of the proposed method.
-Appendix A Proof of Proposition 6
-Lemma 8.
-Suppose $\{\tilde{x}_{t}\}$ satisfies (A1) and (A2). Then
-$\mathcal{W}_{k}\left(\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})},\mu_{p}\right)\xrightarrow{n\rightarrow\infty}0\quad\mathrm{a.s.},$
-where $\tilde{\mathbf{x}}=(\tilde{x}_{0},\ldots,\tilde{x}_{n-1})$.
-Proof.
-By Theorem 7.12 of Villani, (2021), it suffices to show that with probability one,
-$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}\Rightarrow\mu_{p},$ (20)
-and
-$\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|^{k}.$ (21)
-By Theorem 36.4 of Billingsley, (1995), the processes $\{\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|^{k}\}_{t}$ and $\{\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\}_{t}$ are stationary and ergodic for any $\mathbf{h}\in\mathbb{R}^{p}$, where $\mathrm{i}=\sqrt{-1}$.
-Thus (21) follows immediately from the Ergodic theorem.
-To show (20), it suffices to show, with probability one,
-$\hat{\varphi}_{n}(\mathbf{h}):=\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v}_{t}(\tilde{\mathbf{x}}))\rightarrow\int\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{v})d\mu_{p}(\mathbf{v})=:\varphi(\mathbf{h})\quad\mbox{for all }\mathbf{h}\in\mathbb{R}^{p}.$
-By the Ergodic theorem, we have
-$\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q}),\quad\mbox{for all }\mathbf{q}\in\mathbb{Q}^{p},$ (22)
-with probability one. Let
-=$\mathcal{E}=\{\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q})\mbox{ for all }\mathbf{q}\in\mathbb{Q}^{p}\}\cap\left\{\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|\right\},$
-noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2).
-By the Ergodic theorem and (22), $\mathbb{P}(\mathcal{E})=1$.
-Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary.
-Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
-$|\varphi(\mathbf{h})-\varphi(\mathbf{q})|<\epsilon,$ (23)
-for all $\mathbf{q}\in\mathbb{Q}^{p}$ with $\|\mathbf{h}-\mathbf{q}\|<\delta$. In addition, using $|\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{z})-1|\leq\|\mathbf{h}\|\|\mathbf{z}\|$, we have
-$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (24)
-for all $\mathbf{h}$, $\mathbf{q}$ such that $\|\mathbf{h}-\mathbf{q}\|<\delta$. Combining (23)–(24), we have
-$|\hat{\varphi}_{n}(\mathbf{h})-\varphi(\mathbf{h})|\leq$ $|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}(\mathbf{q})|+|\hat{\varphi}_{n}(\mathbf{q})-\varphi(\mathbf{q})|+|\varphi(\mathbf{q})-\varphi(\mathbf{h})|$
-$\leq$ $\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|+o(1)+\epsilon$
-on $\omega\in\mathcal{E}$.
-Note, in particular, that $\epsilon$ and $\delta$ do not depend on $\mathbf{h}$.
-Letting $\epsilon\rightarrow 0$ and $\delta\rightarrow 0$ proves (20) holds on $\mathcal{E}$, and hence with probability one.
-∎
-Proof of Proposition 6.
-We only consider the case $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ here as the other case follows from a similar argument. Define
-$\hat{\mu}_{\mathrm{pre}}$ $:=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}=\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}=:\mu_{\mathrm{pre}},$
-$\hat{\mu}_{\mathrm{post}}$ $:=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})}$
-$\mu_{\mathrm{post}}$ $:=\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\tilde{\mathbf{x}})}$
-Since the Wasserstein distance of order $k$ is a metric on the space of probability measures with finite $k$-th moments, we have
-$\mathcal{W}_{k}(\hat{\mu}_{\mathrm{post}},\mu_{p})\leq$ $\mathcal{W}_{k}(\hat{\mu}_{\mathrm{post}},\hat{\mu}_{\mathrm{pre}})+\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p})$
-$\leq$ $\mathcal{W}_{k}({\mu}_{\mathrm{post}},{\mu}_{\mathrm{pre}})+\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p})$
-$\leq$ $\mathcal{W}_{k}(\mu_{\mathrm{post}},\mu_{p})+2\mathcal{W}_{k}(\mu_{\mathrm{pre}},\mu_{p}).$
-Now the desired result follows from Lemma 8. ∎
-Appendix B Scalar filter algorithm
-The following summarizes the scalar filter algorithm of Peña and Tsay, (2021).
-  • Initialize the imputation $\hat{\mathbf{w}}^{(0)}=(\hat{w}_{0}^{(0)},\hat{w}_{1}^{(0)},\ldots,\hat{w}_{n-1}^{(0)})^{\top}$ with $\hat{w}_{t}^{(0)}=0$ if $t\in\mathcal{M}$.
-  • For each $s\in\mathcal{M}$, let $I^{(s)}_{t}$ be the indicator function. That is,
-    $I_{t}^{(s)}=\left\{\begin{array}[]{cc}1,&t=s\\ 0,&t\neq s\end{array}\right.$
-    and estimate the intervention model using $\hat{\mathbf{w}}^{(0)}$. For example, one can fit an ARX model with $\{I_{t}^{(s)}:s\in\mathcal{M}\}$ as exogenous predictors.
-  • Let $\hat{\beta}^{(s)}$ be the intervention effect associated with $I_{t}^{(s)}$. Estimate the missing data by
-    $\hat{w}_{s}=\hat{w}_{s}^{(0)}-\hat{\beta}^{(s)},\quad s\in\mathcal{M}.$
-As one can see from the algorithm above, the scalar filter algorithm proposed by Peña and Tsay, (2021) is based on intervention analysis.
-It can be iterated in order to improve the estimates.
-However, no convergence guarantee is provided by the authors.
-When fitting an ARX model with the indicators as exogenous predictors, if there are many missing values, standard approaches may be infeasible or unreliable due to near collinearity.
-In this case, we employ a small $\ell_{2}$ penalty to avoid numerical instability.
+The missing data in the test data are removed for fair comparison.
+Figure S3 presents, in violin plots, the distribution of the logarithm of the root mean squared prediction errors ($\log$ RMSE) for each method across the 176 series under study.
+The TWI-type methods effectively alleviate some large RMSEs caused by the benchmark imputation methods and generally produce lower RMSEs.
+This shows the series imputed by TWI have marginal distributions that are stable (across time), meaning the method does not artificially induce structural shifts.
+This further corroborates the practical utility of TWI.
+Figure S3: The distributions of the logarithm of the root mean squared prediction errors (RMSEs) for each method across the 176 series, ranked by median from left to right.

### v3

@@
-Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series
+Temporal Wasserstein Imputation: A Versatile Method for Time Series Imputation
 Abstract
-Missing data can significantly hamper standard time series analysis, yet in practice they are frequently encountered.
+Missing data can significantly hamper standard time series analysis, yet they occur frequently in applications.
 In this paper, we introduce temporal Wasserstein imputation, a novel method for imputing missing data in time series.
-Unlike existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for potential nonlinear dynamics.
+Unlike most existing techniques, our approach is fully nonparametric, circumventing the need for model specification prior to imputation, making it suitable for empirical applications even with nonlinear dynamics.
 Its principled algorithmic implementation can seamlessly handle univariate or multivariate time series with any non-systematic missing pattern.
 In addition, the plausible range and side information of the missing entries (such as box constraints) can easily be incorporated.
-As a key advantage, our method mitigates the distributional bias typical of many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series.
-Leveraging the benign landscape of the optimization formulation, we establish the convergence of an alternating minimization algorithm to critical points.
+Furthermore, our method mitigates the distributional bias common among many existing approaches, ensuring more reliable downstream statistical analysis using the imputed series.
+We establish the convergence of an alternating minimization algorithm to critical points.
 We also provide conditions under which the marginal distributions of the underlying time series can be identified.
-Numerical experiments, including extensive simulations covering linear and nonlinear time series models and a real-world groundwater dataset laden with missing values, corroborate the practical usefulness of the proposed method.
-Keywords: Optimal transport, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
+Numerical experiments, including extensive simulations covering both linear and nonlinear time series and an analysis on a real-world groundwater dataset, corroborate the practical usefulness of the proposed method.
+Keywords: Optimal transport, Missing values, Nonlinear time series, Threshold AR model, Compositional time series, Nonstationary time series
 1 Introduction
-In many scientific fields, data collection is often subject to various disruptions such as idiosyncratic or system-level instrument failures or changes in data collection policies, leading to missing entries in the data (Little and Rubin,, 2019).
-In time series analysis, missing data pose serious challenges to the application of many statistical tools since it is typically presumed that observations are available at equally spaced time points.
-Moreover, since the dynamic relationship is of crucial importance in time series analysis, one cannot proceed by throwing away the missing data, even if the missing data appear completely at random (i.e., the missing pattern is independent of the series).
-To tackle this issue, one approach is to impute the missing entries, after which the vast array of time series methods becomes readily applicable for statistical inference in the downstream.
-Naturally, the estimation of the optimal interpolator, which is the expectation of the missing data conditional on the observed data, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021; McElroy and Politis,, 2020, 2022; McElroy,, 2022).
-Since the optimal interpolator would depend on the probabilistic structure of the underlying process, most methods involve EM-type iterations that alternate between model identification and missing data imputation (see, for example, Chpt.
-4 of Peña and Tsay,, 2021).
-Besides the optimal interpolator, ad hoc deterministic imputation methods, such as smoothing with linear interpolation, B-splines and wavelets (see, for instance, Hastie et al.,, 2009), are also used in practice for their simplicity.
-However, the optimal interpolator can cause distributional biases that are detrimental to downstream statistical analysis.
-To illustrate, consider an AR(1) process, $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $10<|\phi|<$ 1 and $\{\epsilon_{t}\}$ is a sequence of i.i.d.
-standard normal random variables.
-The optimal interpolator can be expressed in terms of the inverse autocorrelation functions (Peña and Maravall, 1991; McElroy and Politis, 2022; see also Chpt.
-6 of Peña et al.,, 2011).
-For instance, suppose that data at time indices $T$ and $T+1$ are missing.
-Then the optimal interpolators are given by
-$\hat{x}_{T}=\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T+2},\quad\hat{x}_{T+1}=\frac{\phi^{2}}{1+\phi^{2}+\phi^{4}}x_{T-1}+\frac{\phi(1+\phi^{2})}{1+\phi^{2}+\phi^{4}}x_{T+2}.$
-Using the fact that $\mathbb{E}(x_{t}x_{t-h})=\phi^{h}/(1-\phi^{2})$, for $h\geq 0$, it is not difficult to show that
-$\mathbb{E}(\hat{x}_{T}^{2})=\mathbb{E}(\hat{x}_{T+1}^{2})=\frac{\phi^{2}+3\phi^{4}+3\phi^{6}+2\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}\mathbb{E}(x_{t}^{2})<\mathbb{E}(x_{t}^{2}),\mbox{ and}$
-$|\mathbb{E}(\hat{x}_{T}\hat{x}_{T+1})|=\frac{2\phi^{2}+3\phi^{4}+3\phi^{6}+\phi^{8}}{(1+\phi^{2}+\phi^{4})^{2}}|\mathbb{E}(x_{t}x_{t+1})|<|\mathbb{E}(x_{t}x_{t+1})|.$
-Hence, the variance and the autocovariances of the imputed data are different from those of the original AR(1) process.
-In general, the joint distribution of the optimally interpolated time series is often quite different from that of the original process, leading to questionable downstream statistical analysis, especially when the number of missing data is non-negligible.
-Similar biases are also observed for some deterministic imputation methods (Carrizosa et al.,, 2013).
-In addition to the distributional bias, the above example also reveals further predicaments of the optimal interpolator.
-First, in order to compute the optimal interpolator, one often needs to choose a model class before imputation.
-However, such a decision relies heavily on domain expertise and may be ill-informed when there are many missing values.
-While it is possible, for certain time series models, to estimate the model parameters without imputation, it still engages in model specification before imputation.
-Second, it may be difficult to compute the optimal interpolator if some nonlinear model is employed.
-With different missing patterns and different chosen models, the computational procedure may change drastically.
-Finally, incorporating side information about the missing data into this methodology appears less straightforward.
-Side information about the missing data restricts the set of admissible imputations.
-For instance, a compositional time series is known to lie in $[0,1]$ and sums to unity at each time point.
-Besides the optimal interpolator, many deterministic interpolators such as B-spline smoothing are also inclined to over-(or under-)shoot, leading to imputations that are outside plausible ranges.
-In this paper, we present the temporal Wasserstein imputation (TWI), a novel approach for time series imputation which circumvents the aforementioned limitations.
-As a nonparametric method, TWI is highly versatile and can be applied to a wide range of time series data.
-In particular, as a key advantage of TWI, it requires no model class specification prior to imputation, allowing it to perform well even when the underlying time series is nonlinear, such as the threshold autoregressive processes (TAR, Tong,, 1983).
+Data collection is often subject to various disruptions at either the idiosyncratic or system level, leading to missing values.
+In time series analysis, missing data pose serious challenges to the application of many statistical tools because they typically assume that observations were taken at equally spaced time points.
+Indeed, the problem of missing data in time series has received wide attention from numerous fields, including environmental and climate studies (Afrifa-Yamoah et al.,, 2020; Junger and Ponce de Leon,, 2015), geophysics (Shen et al.,, 2015; Schoellhamer,, 2001), and engineering (Li et al.,, 2019; Racault et al.,, 2014).
+Existing methods to cope with missing data abound, which can be classified into two general categories.
+The methods of the first category estimate the missing data before any further analysis.
+For example, the optimal interpolator, which is the expectation of the missing data conditional on the observed data and a given model, has been extensively studied (Harvey and Pierse,, 1984; Pourahmadi,, 1989; Peña and Maravall,, 1991; Gómez et al.,, 1999; Alonso and Sipols,, 2008; Peña and Tsay,, 2021; McElroy and Politis,, 2020, 2022; McElroy,, 2022).
+Without any model assumption, some nonparametric imputation methods are also available.
+For instance, the singular spectrum analysis (SSA, Kondrashov and Ghil,, 2006; Shen et al.,, 2015; Golyandina et al.,, 2001), which relies on the spectral analysis of the trajectory matrix, is a nonparametric method to impute missing time series data.
+Other nonparametric methods include smoothing with polynomial interpolation or B-splines and moment matching (see Carrizosa et al.,, 2013 and Hastie et al.,, 2009).
+The methods of the second category circumvent the imputation stage and directly learns the model parameters.
+For example, by specifying a Gaussian likelihood function, one can use the Kalman filter to carry out estimation.
+See Harvey, (1990) and Little and Rubin, (2019), among others.
+However, the aforementioned methods have some drawbacks.
+On the one hand, the required model specification of the optimal interpolator and the likelihood-based approaches may be ill-informed when there are many missing values.
+Furthermore, since computing the optimal interpolator under a nonlinear model can be challenging either computationally or analytically, simple linear models are often used in practice.
+Finally, the imputed series generally has a different distribution compared to the original process, which biases downstream statistical analysis especially when the number of missing data is non-negligible.
+On the other hand, the nonparametric SSA methods work well with structural time series featuring deterministic trends and seasonality, but are not effective in handling processes described by stochastic models, such as the celebrated ARIMA process.
+This is particularly so when the percentage of missing values is high.
+In this paper, we present the temporal Wasserstein imputation (TWI), a novel approach for time series imputation which addresses the aforementioned limitations.
+As a nonparametric method, TWI requires no model specification prior to imputation, allowing the underlying time series to be nonlinear, such as the threshold autoregressive process (TAR, Tong,, 1983).
+Hence, TWI is highly versatile and can be applied to a wide range of stochastic time series, which nicely complements the SSA-based methods.
 Formulated as an optimization problem, TWI can be implemented efficiently through an alternating minimization algorithm.
-This principled algorithmic approach allows for the seamless integration of side information by restricting the admissible imputations, and its extension to multivariate time series is also straightforward.
-The optimization and statistical underpinnings of TWI are closely related to the optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013), which is briefly reviewed in Section 2.
-Section 3 maps out the optimization formulation and introduces the computational scheme in detail.
-The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series, thereby lowering the distributional bias discussed above.
+This principled algorithmic approach not only is computationally efficient, but can also be easily generalized to handle multivariate or non-stationary time series (see Section 3).
+Additionally, it can seamlessly integrate the side information by restricting the admissible imputations, which can be used to enforce natural constraints such as nonnegativity constraints.
+The optimization and statistical underpinnings of TWI are rooted in the well-known optimal transport problem (Villani,, 2008, 2021; Cuturi,, 2013).
+The core idea of TWI is that by minimizing the Wasserstein distance (over all admissible imputations) between the marginal distributions before and after a specified time point, the marginal distributions implied by the imputation shall resemble those of the original series, which facilitates downstream and exploratory analyses.
+Recently, optimal transport techniques have gained traction in imputation problems under different contexts.
+In their pioneering work, Muzellec et al., (2020) first advocated that “ideal” imputations should preserve the distribution of the underlying data.
+Zhao et al., (2023) called this idea “distribution matching,” and devised a deep invertible function approach to imputation, leveraging the statistical and computational properties of optimal transport (see also Wu et al.,, 2023).
+Hur and Liang, (2024) proposed a new method for individual treatment effect imputation based on computational optimal transport.
+In the time series context, after the completion of this paper we became aware of a concurrent and independent work by Wang et al., (2025), who incorporated a new Fourier-based discrepancy measure into the framework of Muzellec et al., (2020).
+However, because of batch sampling and the non-convexity of the optimization problem, neither Wang et al., (2025) nor Muzellec et al., (2020) has provided satisfactory optimization or statistical guarantees.
+In this paper, we propose an alternating minimization algorithm and study its optimization and statistical properties of TWI.
+Our result shows that the algorithm monotonically reduces the loss and is guaranteed to converge to a critical point.
+Moreover, we show, in a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under certain assumptions.
+This simple example encapsulates the identification problem and provides an original insight into the distribution matching framework.
+The rest of the paper is organized as follows.
+Section 2 provides a brief review of the elements of optimal transport.
+Section 3 maps out the optimization formulation and introduces the computational scheme of TWI in detail.
+It also discusses some extensions and comparisons with certain existing methods.
 In Section 4, we investigate the optimization and statistical properties of TWI.
-Our analysis shows that the optimization landscape of TWI is biconvex and the alternating algorithm is guaranteed to converge to a critical point.
-However, from the statistical viewpoint, missing data imputation suffers from an identification issue: Many distributions (of the stochastic process) may be equally plausible in describing the corrupted series.
-This raises the question of whether and under what conditions TWI can identify the underlying marginal distributions.
-In Section 4.2, we show, via a two-state Markov chain example, that the correct underlying distributions can indeed be identified by TWI under assumptions on the missing pattern and the stability of the imputation.
-The latter condition motivates a simple modification of TWI, termed $k$-TWI, which shows further improvement in imputation quality from TWI in our numerical experiments.
+We also propose a simple modification of TWI based on our theoretical results, termed $k$-TWI, which shows further improvement over TWI in our numerical experiments.
 In Section 5, we evaluate the performance of TWI on synthetic data generated from a wide range of time series models, including nonlinear, multivariate, and nonstationary settings.
-Across all processes considered, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics such as autocovariance and model parameter estimates.
+Across all processes considered, TWI successfully recovers the underlying dynamics and yields favorable downstream statistics.
 Its advantage is particularly prominent in the presence of nonlinear dynamics.
-In Section 6, we apply TWI to a real-world groundwater dataset with many missing data.
-After imputation, we conduct independent component analysis and autoregressive modeling in the downstream.
-Compared to existing methods, TWI effectively captures the underlying dynamics, producing interpretable components as well as accurate autoregressive predictions.
-These numerical results support the practical contributions of TWI.
+In Section 6, we apply TWI to a real-world groundwater dataset.
+We use the imputed series to compute summary statistics and perform autoregressive modeling.
+Compared to existing methods, TWI can effectively capture the underlying dynamics, especially when the missing ratio is high.
 2 Preliminary: Optimal transport
@@
 Given two probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{p}$ and a cost function $L:\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}_{+}$, the optimal transport cost between $\mu$ and $\nu$ is defined as
-[…]
-[…]
-If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$, for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we formally denote as $\mathcal{W}_{k}(\mu,\nu)$.
-Namely,
-[…]
+$\mathsf{OTC}(\mu,\nu)=\inf_{\pi\in\mathcal{M}(\mu,\nu)}\int L(\mathbf{x},\mathbf{y})d\pi(\mathbf{x},\mathbf{y}),$
+where $\mathcal{M}$ is the set of couplings between $\mu$ and $\nu$, i.e., probability measures over $\mathbb{R}^{p}\times\mathbb{R}^{p}$ with marginals $\mu$ and $\nu$.
+The classic interpretation for the minimization problem is finding the most cost-effective way to move a pile of earth distributed as $\mu$ (with volume normalized to one) to fill a pit whose shape is distributed as $\nu$.
+It also has a matching or assignment interpretation that has led to fruitful results in the economics literature (Galichon,, 2016).
+If $L(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|^{k}$, for some $k\geq 1$, the resulting optimal transport cost is known as the Wasserstein distance of order $k$ (to the $k$-th power), which we denote as
+$\mathcal{W}_{k}(\mu,\nu):=\left\{\inf_{\pi\in\mathcal{M}(\mu,\nu)}\int\|\mathbf{x}-\mathbf{y}\|^{k}d\pi(\mathbf{x},\mathbf{y})\right\}^{1/k}=\{\mathsf{OTC}(\mu,\nu)\}^{1/k}.$
 The Wasserstein distance is a metric on the space of probabilities with finite $k$-th moment on $\mathbb{R}^{p}$, and metrizes the weak topology on this space (Villani,, 2021, Chpt.
@@
 The constraints in (1) enforce the joint distribution $\{\pi_{ij}\}$ to be a coupling between $\mu$ and $\nu$, and will henceforth be called the coupling constraints.
-In the discrete case, the optimal coupling $\{\pi_{ij}\}$ typically involves split of masses for discrete measures and is no longer the push-forward under an one-to-one map.
-Recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale parallelization and can be accelerated with GPUs; see Cuturi, (2013) and Peyré and Cuturi, (2019).
+Recent research has made significant progress in the fast computation of the solution $\{\pi_{ij}\}$ through entropic regularization and some algorithms are amenable to large-scale or GPU parallelization; see Cuturi, (2013) and Peyré and Cuturi, (2019).
 3 Temporal Wasserstein imputation
-In this section, we introduce the proposed temporal Wasserstein imputation (TWI) method for time series imputation.
-First, we provide an optimization formulation of the proposed method that directly connects to the optimal transport problem.
-Then, we describe an alternating minimization algorithm to impute the missing values.
-The algorithm also reveals some connections between the proposed method and some existing approaches.
+This section introduces the temporal Wasserstein imputation (TWI) method for time series imputation.
+We first present its optimization formulation, which shows close connection to optimal transport.
+Then, we describe an alternating minimization algorithm, and highlight its relation to some existing approaches.
 3.1 Optimization formulation
-Let $\{\tilde{x}_{t}\}$ be the time series of interest and we observe the time series $\{x_{t}:t=0,1,\ldots,n-1\}$.
-Denote by $\bigstar$ a generic missing value, and $\mathcal{M}$ the set of time indices corresponding to the missing observations.
+Let $\{\tilde{x}_{t}\}$ be the time series of interest and we observe the data $\{x_{t}:t=0,1,\ldots,n-1\}$.
+Denote by $\bigstar$ a generic missing value and $\mathcal{M}$ the set of time indices corresponding to the missing observations.
 Thus, $x_{t}=\bigstar$ if $t\in\mathcal{M}$ and $x_{t}=\tilde{x}_{t}$ otherwise.
 We discuss below the case that $x_{t}$ is stationary.
-The case of unit-root nonstationary series is discussed at the end of this section.
+The case of unit-root nonstationary series is discussed at the end of the next subsection.
 If $\{\tilde{x}_{t}\}$ is stationary, then its $p$-dimensional marginal distribution $\mu_{p}$, i.e.
@@
 The discrepancy between the two empirical distributions can then be measured by the optimal transport cost, which is the minimum value of the following linear program.
-$\mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij},\pi_{ij}\geq 0}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t.
+$\left.\begin{aligned} \mathsf{OTC}(\hat{\mu}_{\mathrm{pre}}(\mathbf{w}),\hat{\mu}_{\mathrm{post}}(\mathbf{w}))=\min_{\mathbf{\Pi}=(\pi_{ij})_{ij},\pi_{ij}\geq 0}&\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\pi_{ij}L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))\\ \mbox{s.t.
 }&\sum_{i=p-1}^{n_{1}}\pi_{ij}=\frac{1}{n-n_{1}-1}\quad\mbox{for all }j\\ &\sum_{j=n_{1}+1}^{n-1}\pi_{ij}=\frac{1}{n_{1}-p+2}\quad\mbox{for all }i,\end{aligned}\right\}$ (2)
@@
 where the constraint set $\mathcal{C}$ delineates the admissible imputations.
-Typically, it enforces the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$.
+Typically, it ensures the imputation to agree with the observed data, in which case $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$.
 It may also reflect the side information available.
@@
 where the regularization parameter $\lambda>0$ is small.
-As we will see in the next subsection, the $\ell_{2}$-regularization term makes a subproblem in the alternating minimization algorithm strongly convex, enhancing numerical stability.
-In addition, in the important special case corresponding to the Wasserstein distance of order 2, i.e., $L(\mathbf{v}_{i}(\mathbf{w}),\mathbf{v}_{j}(\mathbf{w}))=\|\mathbf{v}_{i}(\mathbf{w})-\mathbf{v}_{j}(\mathbf{w})\|^{2}$, the regularization allows one to solve the subproblem by finding a fixed point of a linear system.
-Since $L$ is often chosen to be some power of the Euclidean norm, which induces the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (3) or (4) is called the temporal Wasserstein imputation (TWI) throughout this paper.
+As we will see in the next subsection, the $\ell_{2}$-regularization term enhances the numerical stability of the alternating minimization algorithm.
+Since $L$ is often chosen to induce the Wasserstein distances between the marginal distributions of a time series, the solution $\hat{\mathbf{w}}$ to (3) or (4) is called the temporal Wasserstein imputation (TWI) throughout this paper.
 In this formulation, it is clear that the temporal Wasserstein imputation, which seeks to equate the marginal distributions on both sides of $n_{1}$, is nonparametric.
@@
 To begin, observe that with an initialized $\hat{\mathbf{w}}^{(0)}$, problem (2) is exactly the discrete optimal transport problem.
-Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (2) efficiently by some existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019).
+Thus, at iteration $k=1,2,\ldots$, treating $\mathbf{w}=\hat{\mathbf{w}}^{(k-1)}$ as fixed, we may compute the optimizer $\hat{\mathbf{\Pi}}^{(k)}$ of (2) efficiently by existing algorithms (Cuturi,, 2013; Peyré and Cuturi,, 2019).
 This corresponds to the inner optimization problem of (4).
 Next, we optimize (4) over $\mathbf{w}$, taking $\mathbf{\Pi}=\hat{\mathbf{\Pi}}^{(k)}$ as fixed.
-This procedure is repeated until convergence.
+This procedure is repeated until convergence (discussed in the next section).
 Let
@@
 and the procedure described above can be summarized in Algorithm 1.
-One advantage of the alternating minimization scheme is that the subproblems (steps (a) and (b) in Algorithm 1) are easy to optimize when $L$ satisfies some regularity conditions.
-Optimization guarantees are discussed in the next subsection.
 Input: Cut-off $n_{1}$, number of lags $p$
@@
 Algorithm 1 Temporal Wasserstein imputation with alternating minimization
-Next, we investigate the properties of Algorithm 1 in an important special case where closed forms can be obtained.
+One advantage of the alternating minimization scheme is that the subproblems (steps (a) and (b) in Algorithm 1) are easy to optimize when $L$ satisfies some regularity conditions.
 Suppose $L(\mathbf{v}_{i},\mathbf{v}_{j})=\|\mathbf{v}_{i}-\mathbf{v}_{j}\|^{2}$.
@@
 $\mathbf{A}_{h}(\mathbf{\Pi})$ $=\begin{pmatrix}\mathbf{0}_{p-1-h}&&&\\ &\frac{1}{n_{1}-p+2}\mathbf{I}_{n_{1}-p+2}&-\mathbf{\Pi}&\\ &-\mathbf{\Pi}^{\top}&\frac{1}{n-n_{1}-1}\mathbf{I}_{n-n_{1}-1}&\\ &&&\mathbf{0}_{h}\end{pmatrix}\in\mathbb{R}^{n\times n},\quad h=0,1,\ldots,p-1,$
-in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero, and $\mathbf{0}_{h}$ is an $h\times h$ matrix of zeros.
+in which all unspecified entries in $\mathbf{A}_{h}(\mathbf{\Pi})$ are zero.
 Hence $F$ is quadratic in $\mathbf{w}$.
-In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints.
-Moreover, $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 3 in Section 4).
-Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004; Parikh et al.,, 2014).
-If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e., $\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$ for some known $\mathbf{K}$ and $\mathbf{b}$, then step (b) in Algorithm 1 has the following closed form
+In fact, it can be shown that $\mathbf{H}(\mathbf{\Pi})$ is nonnegative definite for any $\mathbf{\Pi}$ satisfying the coupling constraints, and $F$ is strongly convex in $\mathbf{w}$ if $\lambda>0$ (see also Proposition 3 in Section 4).
+Thus many efficient solvers are applicable in this step (Boyd and Vandenberghe,, 2004).
+If, in addition, the constraint set $\mathcal{C}$ is defined through a system of linear equations, i.e., $\mathcal{C}=\{\mathbf{w}\in\mathbb{R}^{n}:\mathbf{Kw}=\mathbf{b}\}$ where $\mathbf{K}$ and $\mathbf{b}$ are known, then step (b) in Algorithm 1 has the closed form
 $\hat{\mathbf{w}}^{(k)}=\mathbf{H}^{-1}\mathbf{K}^{\top}(\mathbf{K}\mathbf{H}^{-1}\mathbf{K}^{\top})^{-1}\mathbf{b},$ (6)
 where $\mathbf{H}=\mathbf{H}(\hat{\mathbf{\Pi}}^{(k)})$.
-Suppose further that $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$.
-That is, there is no side information about the missing values.
-In this case, the solution $(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ to the subproblem in step (b) must satisfy the following system of linear equations which has the same number of unknowns and equations.
-$\hat{w}_{s}=&\frac{\sum_{i=s\vee(p-1)}^{s+p-1}\sum_{n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}}{\frac{p\wedge(s+1)}{n_{1}-p+2}+\frac{\lambda}{2}},\quad\mbox{if }0\leq s\leq n_{1}-p+1,s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=s}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}+\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{s+p-1}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{n_{1}-s+1}{n_{1}-p+2}+\frac{s+p-n_{1}+1}{n-n_{1}-1}+\frac{\lambda}{2}},\\ &\hskip 142.26378pt\mbox{if }n_{1}-p+2\leq s\leq n_{1},s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=p-1}^{n_{1}}\sum_{j=s}^{(s+p-1)\wedge(n-1)}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{p\wedge(n-s)}{n-n_{1}-1}+\frac{\lambda}{2}},\quad\mbox{if }n_{1}+1\leq s\leq n-1,s\in\mathcal{M}\\ \hat{w}_{s}=&x_{s},\quad\mbox{if }s\notin\mathcal{M}\end{aligned}\right.$ (7)
-where $\hat{\mathbf{\Pi}}=(\hat{\pi}_{ij})$ is treated as fixed. We make a few remarks regarding (7).
+Suppose further that $\mathcal{C}=\{\mathbf{w}:w_{t}=x_{t},t\notin\mathcal{M}\}$, corresponding to the case of no side information.
+The solution $(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ to the subproblem in step (b) must satisfy the following system of linear equations.
+$\left\{\begin{aligned} \hat{w}_{s}=&\frac{\sum_{i=s\vee(p-1)}^{s+p-1}\sum_{n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}}{\frac{p\wedge(s+1)}{n_{1}-p+2}+\frac{\lambda}{2}},\quad\mbox{if }0\leq s\leq n_{1}-p+1,s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=s}^{n_{1}}\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}\hat{w}_{j-i+s}+\sum_{i=p-1}^{n_{1}}\sum_{j=n_{1}+1}^{s+p-1}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{n_{1}-s+1}{n_{1}-p+2}+\frac{s+p-n_{1}+1}{n-n_{1}-1}+\frac{\lambda}{2}},\\ &\hskip 142.26378pt\mbox{if }n_{1}-p+2\leq s\leq n_{1},s\in\mathcal{M}\\ \hat{w}_{s}=&\frac{\sum_{i=p-1}^{n_{1}}\sum_{j=s}^{(s+p-1)\wedge(n-1)}\hat{\pi}_{ij}\hat{w}_{i-j+s}}{\frac{p\wedge(n-s)}{n-n_{1}-1}+\frac{\lambda}{2}},\quad\mbox{if }n_{1}+1\leq s\leq n-1,s\in\mathcal{M}\\ \hat{w}_{s}=&x_{s},\quad\mbox{if }s\notin\mathcal{M}\end{aligned}\right$.
+(7)
+where $\hat{\mathbf{\Pi}}=(\hat{\pi}_{ij})$ is treated as fixed, and for $a,b\in\mathbb{R}$, $a\wedge b=\min\{a,b\}$, $a\vee b=\max\{a,b\}$. We make a few remarks regarding (7).
 Remark 1.
 Let us contrast TWI with the optimal interpolator.
-(7) shows, in general, the temporal Wasserstein imputation is a linear function of many observed data $\{x_{t}:t\notin\mathcal{M}\}$, even when $p$ is small.
-On the other hand, the optimal interpolator is usually a linear combination of nearby observed data when the underlying process is a finite-order AR model.
-For instance, suppose the underlying time series is an AR(1) process: $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\{\epsilon_{t}\}$ is a mean-zero, unit-variance white noise.
-If $x_{n_{1}}=\bigstar$ is the only missing value in the time series, then the optimal interpolator (Peña and Maravall,, 1991) of $x_{n_{1}}$ is
-$w_{n_{1},\mathrm{opt}}=\frac{\phi}{1+\phi^{2}}(x_{n_{1}-1}+x_{n_{1}+1}).$
+Equation (7) shows, in general, the temporal Wasserstein imputation is a linear function of many observed data $\{x_{t}:t\notin\mathcal{M}\}$, even when $p$ is small.
+In contrast, the optimal interpolator is often a function of nearby observed data.
+For instance, suppose the underlying time series is a stationary AR(1) process: $x_{t}=\phi x_{t-1}+\epsilon_{t}$, where $\{\epsilon_{t}\}$ is a mean-zero, unit-variance white noise.
+If $x_{n_{1}}=\bigstar$ is the only missing value in the time series, then the optimal interpolator (Peña and Maravall,, 1991) of $x_{n_{1}}$ is $w_{n_{1},\mathrm{opt}}=\phi(x_{n_{1}-1}+x_{n_{1}+1})/(1+\phi^{2})$.
 In contrast, (7) shows that the temporal Wasserstein imputation (with $p=2$ and $\lambda=0$) is given by
 $\hat{w}_{n_{1}}=\frac{1}{n_{1}^{-1}+(n-n_{1}-1)^{-1}}\left(\sum_{t=0}^{n_{1}-1}\hat{\pi}_{t+1,n_{1}+1}x_{t}+\sum_{t=n_{1}+1}^{n-1}\hat{\pi}_{n_{1},t}x_{t}\right),$
-which is a convex combination of all observed values. This is somewhat expected because, as a nonparametric method, TWI uses all data to learn the underlying dynamics.
+which is a convex combination of all observed values.
+This is expected because, as a nonparametric method, TWI uses all data to learn the underlying dynamics.
+Note that convex combination appears naturally in imputation problems.
+For instance, in synthetic control, a convex combination of untreated units is used to estimate the counterfactual (see Hur and Liang,, 2024; Abadie,, 2021).
 Remark 2.
-(EM interpretation) Let $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ be a fixed point of the alternating minimization algorithm (Algorithm 1). That is,
-$\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}\in\mathcal{C}}F(\mathbf{w},\hat{\mathbf{\Pi}}),\quad\hat{\mathbf{\Pi}}\in\arg\min_{\mathbf{\Pi}}F(\hat{\mathbf{w}},\mathbf{\Pi}).$
-In addition, suppose $\lambda=0$ for a moment.
-Let $\mathbf{u}=(u_{0},u_{1},\ldots,u_{p-1})$ be a random vector whose distribution is $\hat{\mu}_{\mathrm{pre}}$ and $\mathbf{v}=(v_{0},\ldots,v_{p-1})$ a random vector whose distribution is $\hat{\mu}_{\mathrm{post}}$.
-Furthermore, let the joint distribution of $(\mathbf{u},\mathbf{v})$ be given by the optimal transport coupling between $\hat{\mu}_{\mathrm{pre}}$ and $\hat{\mu}_{\mathrm{post}}$, identified as $\hat{\mathbf{\Pi}}$ (by viewing it as a joint distribution between the two discrete measures).
-Consider an index $s\in\mathcal{M}$ with $p-1\leq s\leq n_{1}-p+1$.
-By (7), it is not difficult to see that
+[…]
 $\hat{w}_{s}=$ $\frac{1}{p}\sum_{i=s}^{s+p-1}\sum_{j=n_{1}+1}^{n-1}\frac{\hat{\pi}_{ij}}{\sum_{j=n_{1}+1}^{n-1}\hat{\pi}_{ij}}\hat{w}_{j-i+s}$
 $=$ $\frac{1}{p}\sum_{h=0}^{p-1}\mathbb{E}_{\hat{\mathbf{\Pi}}}(v_{h}|\mathbf{u}=(\hat{w}_{s+h},\ldots,\hat{w}_{s},\ldots,\hat{w}_{s+h-p+1})).$ (8)
-(8) shows that the temporal Wasserstein imputation $\hat{w}_{s}$ is essentially an average of conditional expectations.
+Equation (8) shows that the temporal Wasserstein imputation $\hat{w}_{s}$ is essentially an average of conditional expectations.
 Thus, Algorithm 1 is closely related to the EM algorithm: To minimize the Wasserstein distance $\mathbb{E}_{\mathbf{\Pi}}\|\mathbf{u}-\mathbf{v}\|^{2}$, step (b) computes the conditional expectations under the current optimal transport coupling $\hat{\mathbf{\Pi}}$ while step (a) minimizes over the coupling space.
-Formula like (8) can be derived for other time indices $s$ with appropriate modifications.
+Similar formulas to (8) can be derived for other time indices $s$ with appropriate modifications.
 Because of the optimization formulation, it is straightforward to extend TWI to multivariate time series.
@@
 Then Algorithm 1 can still be applied.
-Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 is identical as the scalar case.
+Since the multivariate information is summarized by the distance matrix $\mathbf{L}=(l_{ij})_{0\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}$, where $l_{ij}=L(\mathbf{v}_{i}(\mathbf{W}),\mathbf{v}_{j}(\mathbf{W}))$, regardless of the dimension $d$, step (a) in Algorithm 1 remains unchanged.
 In the notable special case where $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{2}$, (5) becomes
@@
 Straightforward calculations show that fixing $\mathbf{\Pi}=\hat{\mathbf{\Pi}}$, if $\hat{\mathbf{W}}$ is a solution to step (b) of Algorithm 1, then for each $l\in\{1,2,\ldots,d\}$, $(\hat{w}_{0,l},\hat{w}_{1,l},\ldots,\hat{w}_{n-1,l})$ also satisfies the system of linear equations (7) in the case of no side information.
-Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel to lower the computational overhead and only the values of individual series are involved in computing the imputation in each component.
-The cross-sectional information is exploited when learning the optimal transport coupling in step (a).
+Therefore, in step (b) of Algorithm 1, each component in the multivariate time series can be imputed in parallel, lowering the computational overhead.
 Thus, the algorithm is fast to implement in the multivariate setting.
-We shall compare TWI with the approach proposed by Carrizosa et al., (2013), which also aims to alleviate the corrupted statistical patterns caused by imputation.
-Their method is based on penalizing the deviation of some moments from the theoretical values.
-As a result, the moments, such as the autocorrelation functions, of the imputed series are close to the pre-specified values.
-However, this approach has a number of drawbacks.
-First, good theoretical values for the moments may not be readily available in practice, especially when the missing values are many.
-Second, their approach promotes smoothness of the time series and hence positive autocorrelation, which does not necessarily align with the underlying time series.
-Finally, the optimization problem is highly non-convex, leading to an intense computational hurdle.
-In contrast, the proposed TWI is data-driven, completely nonparametric, and enjoys the computational advantages brought by the alternating minimization algorithm and recent advances in computational optimal transport.
-Next, we discuss the imputation of nonstationary time series.
+Before closing this subsection, we discuss the imputation of nonstationary time series.
 For an $I(1)$ process $\{x_{t}\}$, $y_{t}=x_{t}-x_{t-1}$ is stationary and one can apply imputation methods to the first-differenced series $\{y_{t}\}$.
@@
 Then, an imputation for $x_{t}$ can be recovered by inverting $\phi(B)$.
-Note that, in this case, the number of missing data in $\{y_{t}\}$ multiplies, and we need at least $d+1$ contiguous observed values in $\{x_{t}\}$ to obtain an observed value for $\{y_{t}\}$.
-To mitigate the information loss, we can encode the information from the raw data via the admissible set $\mathcal{C}$.
-For example, consider the $I(1)$ case: If $x_{t}$ is observed, then since $y_{1}+\ldots+y_{t}=x_{t}-x_{0}$, we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed).
-Finally, although we have motivated TWI from the stationarity of the underlying process, TWI fares quite well with certain nonstationary processes such as cyclic time series.
-In Section 5, we further explore the application of TWI to nonstationary time series.
+However, in this case, the number of missing data in $\{y_{t}\}$ multiplies, and at least $d+1$ contiguous observed values in $\{x_{t}\}$ is needed to obtain an observed value for $\{y_{t}\}$.
+To mitigate the information loss, we can encode the information from the raw data via an admissible set $\mathcal{C}$.
+For example, consider the $I(1)$ case: If $x_{t}$ is observed, then since $y_{1}+\ldots+y_{t}=x_{t}-x_{0}$, we may restrict the imputation $\{w_{t}\}$ for $\{y_{t}\}$ to satisfy $w_{1}+\ldots+w_{t}=x_{t}-x_{0}$ (assuming $x_{0}$ is observed), which constitutes a set of linear constraints $\mathcal{C}$.
+Finally, although we have motivated TWI using the stationarity of the underlying process, it fares quite well with certain nonstationary processes, in addition to unit-root series, such as cyclic time series (see Section 5 for a simulation study).
+3.3 Related works
+We shall compare TWI with some existing nonparametric imputation approaches.
+Carrizosa et al., (2013) proposed a moment-matching method, which tries to minimize the discrepancies between the moments (e.g.
+autocorrelations) of the imputed series and some pre-specified target values.
+However, their non-convex algorithm is computationally intensive, and good target values for the moments may not be readily available in practice.
+In contrast, TWI is completely data-driven, and enjoys the computational efficiency of the alternating minimization algorithm and computational optimal transport.
+Singular spectrum analysis (SSA) methods (Golyandina et al.,, 2018, 2001; Kondrashov and Ghil,, 2006) are nonparametric techniques that perform spectral analysis on the trajectory matrix of the time series $\mathbf{x}=(x_{0},x_{1},\ldots,x_{n-1})$.
+Specifically, the trajectory matrix is defined as $\mathbf{T}=(\mathbf{y}_{L-1},\mathbf{y}_{L},\ldots,\mathbf{y}_{n-1})$, where $\mathbf{y}_{j}=(x_{j-L+1},x_{j-L+2},\ldots,x_{j})^{\top}$ for $j=L-1,\ldots,n-1$, and $L$ is called the window length.
+SSA methods exploit the possible low-rank structure of $\mathbf{T}$.
+When $\mathbf{T}$ contains missing values, columns with missing entries are typically removed (shaped SSA, Golyandina et al.,, 2018), followed by direct or iterative imputation (Kondrashov and Ghil,, 2006).
+For some missing patterns this could result in too few complete columns, forcing $L$ to be small (violating a common guideline of $L\approx n/2$) and compromising the effectiveness of SSA.
+In contrast, TWI accommodates any missing pattern without discarding substantial data.
+While SSA is typically effective for imputing structural time series, TWI prevails in learning stochastic time series (see Section 5 for our simulation study).
+However, due to the convex nature of TWI (see Remark 1), it cannot handle series with explosive deterministic trends.
+Hence, it is advisable to remove the trends before applying TWI, or use SSA-based methods when the deterministic patterns are of primary interest.
+A concurrent and independent work of Wang et al., (2025) proposed the Proximal Spectrum Wasserstein for Imputation (PSW-I), which, following the framework of Muzellec et al., (2020), imputes time series using optimal transport techniques.
+While both PSW-I and TWI focus on distribution matching, their implementations differ significantly.
+PSW-I uses a gradient descent algorithm that relies on automatic differentiation and some ad hoc gradient-stabilizing techniques, due to the difficulty in computing the gradient of the (regularized) Wasserstein distance with respect to the imputation.
+TWI, by comparison, uses a straightforward algorithm with minimal tuning.
+Although Wang et al., (2025) includes a convergence result, its guarantees hinge on restrictive assumptions.
+In contrast, TWI’s convergence and statistical properties are established under more standard conditions, as detailed in the next section.
 4 Theoretical properties
-In this section, we investigate the optimization and statistical properties of TWI.
+This section investigates the optimization and statistical properties of TWI.
 We first show Algorithm 1 converges to critical points.
-Then, asymptotic consistency of TWI is established when the missing data concentrate on one side of $n_{1}$.
-For general non-systematic missing patterns, we show that identification of the underlying marginal distributions can be achieved by TWI under the assumptions of interpolator stability using a two-state Markov chain example (see (C1) and (C2) in Section 4.2).
-Before presenting the results, it is worthwhile to clarify the asymptotic regime under consideration.
-The analysis is meaningful only when the number of missing entries is non-negligible compared to the sample size.
-Otherwise, if $\sharp(\mathcal{M})=o(n)$, where $\sharp(E)$ is the cardinality of the set $E$, then the empirical marginal distributions of any (bounded) imputations satisfy
-$\frac{1}{n}\sum_{t\in\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}+\frac{1}{n}\sum_{t\notin\mathcal{A}_{n}}\delta_{(w_{t},w_{t-1},\ldots,w_{t-p+1})}\Rightarrow\mu_{p},$
-where $\mathcal{A}_{n}=\{t\in\{p-1,\ldots,n-1\}:\{t,t-1,\ldots,t-p+1\}\cap\mathcal{M}\neq\emptyset\}$.
-Therefore, assuming $\sharp(\mathcal{M})=o(n)$ would essentially preclude distributional inconsistency in the downstream.
-To avoid this trivial case, in this paper we maintain instead the assumption that $\mathcal{M}=\mathcal{M}_{n}$ is a sequence of index sets (independent of the underlying time series) such that $\sharp(\mathcal{M}_{n})/n\asymp 1$, where $a_{n}\asymp b_{n}$ means there exists $0<\underline{c}<\bar{c}<\infty$ such that $\underline{c}\leq a_{n}/b_{n}\leq\bar{c}$ for all large $n$.
+We then establish the asymptotic consistency of TWI when the missing data concentrate on one side of $n_{1}$.
+For general non-systematic missing patterns, a case study of a two-state Markov chain example shows that TWI can identify the underlying marginal distributions under some natural assumptions.
 4.1 Optimization and statistical guarantees
-We begin by studying the optimization properties of Algorithm 1. All proofs are relegated to Section S1 of the supplementary material.
+We begin by studying the optimization properties of Algorithm 1.
+All proofs are relegated to Section S1 of the supplementary material.
+The following result shows Algorithm 1 is guaranteed to find a local minimum.
 Proposition 3.
@@
   • (b)
-    every limit point $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $\{(\hat{\mathbf{w}}^{(t+1)},\hat{\mathbf{\Pi}}^{(t)}):t=1,2,\ldots\}$ is a critical point of $F$. i.e.,
+[…]
     $\frac{\partial F(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})}{\partial\mathbf{w}}\cdot(\mathbf{w}-\hat{\mathbf{w}})\geq 0,\quad\frac{\partial F(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})}{\partial\mathrm{vec}(\mathbf{\Pi})}\cdot\mathrm{vec}(\mathbf{\Pi}-\hat{\mathbf{\Pi}})\geq 0$
     for any $\mathbf{w}\in\mathcal{C}$ and $\mathbf{\Pi}$ satisfying the coupling constraints.
-If, in addition, the cost function $L$ takes the form $L(\mathbf{u},\mathbf{v})=\sum_{k=1}^{p}\ell(u_{k}-v_{k})$, where $\mathbf{u}=(u_{1},\ldots,u_{p})^{\top}$, $\mathbf{v}=(v_{1},\ldots,v_{p})^{\top}$, and $\ell$ is twice differentiable and convex, then
   • (c)
-    $F(\mathbf{w},\mathbf{\Pi})$ is biconvex in $\mathbf{w}$ and $\mathbf{\Pi}$.
+    If, in addition, the cost function $L$ takes the form $L(\mathbf{u},\mathbf{v})=\sum_{k=1}^{p}\ell(u_{k}-v_{k})$, where $\mathbf{u}=(u_{1},\ldots,u_{p})^{\top}$, $\mathbf{v}=(v_{1},\ldots,v_{p})^{\top}$, and $\ell$ is twice differentiable and convex, then $F(\mathbf{w},\mathbf{\Pi})$ is biconvex in $\mathbf{w}$ and $\mathbf{\Pi}$.
 Due to the biconvexity of $F$, one can employ off-the-shelf solvers for the subproblems in Algorithm 1.
@@
 Hence, step (b) in Algorithm 1 admits a unique solution in each iteration and can be solved efficiently.
-A few remarks are in order.
-First, it is straightforward to generalize Proposition 3 for multivariate time series.
-Second, like most methods involving non-convex optimization, a good initialization is crucial for converging to a “good” critical point.
-In our simulation studies, we use simple linear interpolation or ARIMA-based Kalman smoothing as initializations and TWI is able to find an imputation that offers substantial improvements over these initializations.
-Third, unlike many non-convex optimization problems, we are not satisfied by an arbitrary global minimizer of $F$.
-Essentially, it is because among all processes that have stable $p$-dimensional marginal distributions, only a subset of them matches the underlying marginal distributions.
-In Section 4.2, we probe this identification issue in more detail.
-However, when the missing data happen to lie on the same side of $t=n_{1}$, the global minimizer of $F$ achieves asymptotic consistency in the sense of Proposition 4 below.
+It is straightforward to generalize Proposition 3 for multivariate time series, as discussed in Section 3.
+Unlike typical non-convex optimization problems, we are only interested in the subset of the global minimizers of $F$ which have the correct marginal distributions (see Section 4.2).
+The problem simplifies when the missing data happen to lie on the same side of $t=n_{1}$, in which case the global minimizer of $F$ achieves asymptotic consistency.
 Proposition 4.
-Assume (A1) and (A2) in Section S1 of the supplementary material.
-Suppose for any $\mathbf{w}\in\mathcal{C}$, $w_{t}=\tilde{x}_{t}$ if $t\notin\mathcal{M}_{n}$.
-Suppose further $n$ and $n_{1}$ satisfy $n_{1}\rightarrow\infty$ and $n-n_{1}\rightarrow\infty$.
-If either $\mathcal{M}_{n}\subseteq\{0,1,\ldots,n_{1}\}$ or $\mathcal{M}_{n}\subseteq\{n_{1}+1,\ldots,n-1\}$ (but not both), then, the global minimizer $(\hat{\mathbf{w}},\hat{\mathbf{\Pi}})$ of $F$ with $\lambda=0$ and $L(\mathbf{u},\mathbf{v})=\|\mathbf{u}-\mathbf{v}\|^{k}$, subject to $\mathbf{w}\in\mathcal{C}$ and the coupling conditions, satisfies
+[…]
 $\mathcal{W}_{k}\left(\frac{1}{n_{1}-p+2}\sum_{t=p-1}^{n_{1}}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)+\mathcal{W}_{k}\left(\frac{1}{n-n_{1}-1}\sum_{t=n_{1}+1}^{n-1}\delta_{\mathbf{v}_{t}(\hat{\mathbf{w}})},\mu_{p}\right)\rightarrow 0\quad\mbox{almost surely}.$
 […]
+Corollary 5.
+Assume that the conditions of Proposition 4 hold.
+Then for continuous $\varphi:\mathbb{R}^{p}\rightarrow\mathbb{R}$ with $|\varphi(x)|\leq C(1+\|x\|^{k})$ for all $x\in\mathbb{R}^{p}$ and some $C<\infty$, we have $\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\varphi(\hat{w}_{t},\ldots,\hat{w}_{t-p+1})\rightarrow\mathbb{E}\varphi(\tilde{x}_{t},\ldots,\tilde{x}_{t-p+1})$ a.s., where $\hat{\mathbf{w}}=(\hat{w}_{0},\ldots,\hat{w}_{n-1})$ is defined in Proposition 4.
 4.2 General missing pattern: A two-state Markov chain example
-In this subsection, the identification issue of the temporal Wasserstein imputation method is studied via a two-state Markov chain example.
-Particularly, we connect the non-identification of the underlying marginal distributions to the lack of uniqueness of the solutions to a system of linear equations.
+In this subsection, we study the identification of the underlying marginal distributions via a two-state Markov chain example under the general missing pattern where missing data are not confined to one side of time index $n_{1}$.
+We connect the non-identification problem to the lack of uniqueness of the solutions to a system of linear equations.
 We also show that, under two natural conditions, the correct marginal distributions can still be pinned down by TWI.
 To free our discussion from finite-sample complications, we will focus on the distribution level in this subsection, or intuitively speaking, assume “infinite sample.”
-Suppose the underlying time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain. Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by
-$\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=0.6,\quad$ $\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=0.2.$
-The transition probabilities are specified to facilitate discussions, but the main results below are not specific to these choices.
-It is easy to check that $\{\tilde{x}_{t}\}$ has the stationary distribution given by $\mathbb{P}(\tilde{x}_{t}=1)=0.75$ and $\mathbb{P}(\tilde{x}_{t}=0)=0.25$.
-Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is specified by
-$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=0.6,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=0.15,$ (9)
-$\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=0.15,\quad\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=0.1.$
-Without loss of generality, we set $n_{1}=0$.
-Suppose that before time $n_{1}$, there is a missing value every $q_{1}$ observations.
-Likewise, there is a missing value every $q_{2}$ observations after time $n_{1}$.
-Specifically, $x_{t}=\bigstar$ if $-t$ is a positive integer multiple of $q_{1}$, and similarly, $x_{t}=\bigstar$ if $t$ is a positive integer multiple of $q_{2}$, where $q_{1}$ and $q_{2}$ are some integers larger than 2 and $q_{1}\neq q_{2}$ (see condition (C1) below).
-Next, suppose an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ has been made such that if $t<n_{1}$,
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ (10)
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$
-and if $t\geq n_{1}$,
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1)|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ (11)
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0)|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$
+Suppose the time series $\{\tilde{x}_{t}\}_{-\infty<t<\infty}$ under study is a two-state Markov chain.
+Specifically, $\tilde{x}_{t}\in\{0,1\}$ and the transition probabilities are given by $\mathbb{P}(\tilde{x}_{t}=1|\tilde{x}_{t-1}=0)=p$, and $\mathbb{P}(\tilde{x}_{t}=0|\tilde{x}_{t-1}=1)=q$, where $p,q\in(0,1)$.
+The associated stationary distribution is denoted by $\lambda_{1}:=\mathbb{P}(\tilde{x}_{t}=0)$ and $\lambda_{2}:=\mathbb{P}(\tilde{x}_{t}=1)$.
+Consequently, the 2-dimensional marginal distribution of $\{\tilde{x}_{t}\}$ is
+$\begin{array}[]{llll}\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,1))=\lambda_{2}(1-q),&\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(1,0))=\lambda_{1}p,\\ \mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,1))=\lambda_{2}q,&\mathbb{P}((\tilde{x}_{t},\tilde{x}_{t-1})=(0,0))=\lambda_{1}(1-p).\end{array}$ (9)
+We consider the following missing pattern.
+Without loss of generality, set $n_{1}=0$.
+Suppose that before time $n_{1}$, there is a missing value every $k_{1}$ observations.
+Likewise, there is a missing value every $k_{2}$ observations after time $n_{1}$.
+Specifically, $x_{t}=\bigstar$ if t-t is a positive integer multiple of $k_{1}$, or if $t$ is a positive integer multiple of $k_{2}$, where $k_{1}$ and $k_{2}$ are some integers larger than 2 and $k_{1}\neq k_{2}$ (see condition (C1) below).
+Next, we parametrize an imputation $\{\hat{w}_{t}\}_{-\infty<t<\infty}$ as follows.
+Specifically, let
+$\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,1))=a_{1},$ $\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,0))=b_{1},$ $\mbox{ if }t<n_{1},$ (10)
+$\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,1))=a_{2},$ $\mathbb{P}(\hat{w}_{t}=1|(x_{t},x_{t-1})=(\bigstar,0))=b_{2},$ $\mbox{ if }t\geq n_{1},$ (11)
 for some $a_{1},a_{2},b_{1},b_{2}\in[0,1]$.
-Note that $\mathbb{P}$ in (10) and (11) should be interpreted as proportions since the imputation may depend on the whole sequence of $\{x_{t}\}$, and it is assumed that they are well-defined.
-Essentially, here we parametrize all imputation methods by $a_{1},a_{2},b_{1},b_{2}$, which correspond to different 2-dimensional marginal distributions.
-Indeed, from (9)–(11), we can compute the 2-dimensional marginal distributions of the imputed series.
-For $\{\hat{w}_{t}\}_{-\infty<t<0}$,
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))$ $=0.6(1-\frac{1}{q_{1}})+\frac{3a_{1}}{4q_{1}},\quad\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))=0.15(1-\frac{1}{q_{1}})+\frac{b_{1}}{4q_{1}},$
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))$ $=0.15(1-q_{1}^{-1})+\frac{3}{4q_{1}}(1-a_{1}),$
-$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))$ $=0.1(1-q_{1}^{-1})+\frac{1}{4q_{1}}(1-b_{1}).$
-For $\{\hat{w}_{t}\}_{0\leq t<\infty}$, we can obtain similar expressions with $a_{1},b_{1},q_{1}$ replaced by $a_{2},b_{2},q_{2}$.
-The “correct” imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting 2-dimensional marginal distributions equal the marginal distribution of the underlying series, given in (9).
-To achieve such identification, the correct imputation must choose
-$a_{1}^{*}=a_{2}^{*}=0.8,\quad b_{1}^{*}=b_{2}^{*}=0.6.$ (12)
-Note that this calibration does not depend on $q_{1},q_{2}$; it only depends on the distributional properties of the underlying process.
+Essentially, we parametrize all possible imputations by $a_{1},a_{2},b_{1},b_{2}$, each corresponding to different 2-dimensional marginal distributions.
+The correct imputation should calibrate $a_{i}$’s and $b_{i}$’s such that the resulting marginal distributions equal those of the underlying series, given in (9).
+It is not difficult to see that the correct imputation must choose
+$a_{1}^{*}=a_{2}^{*}=1-q,\quad b_{1}^{*}=b_{2}^{*}=p.$ (12)
+Note that this calibration does not depend on the missing pattern $k_{1},k_{2}$.
 Assuming the optimization is carried out perfectly [^1] This is where we ignore finite sample complications and assume a zero Wasserstein loss is achieved.
-One may use, for example, the hamming distance $L(x,y)=\mathbf{1}\{x\neq y\}$ as the ground metric for the discrete data., the temporal Wasserstein imputation (with $p\geq 2$) calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$.
+One may use, for example, the hamming distance as the ground metric for the optimal transport cost., TWI calibrates $a_{i},b_{i},i=1,2$, by equating the 2-dimensional marginal distributions of $\{\hat{w}_{t}\}_{-\infty<t<0}$ and $\{\hat{w}_{t}\}_{0\geq t<\infty}$.
 Hence, $\{a_{1},a_{2},b_{1},b_{2}\}$ satisfies
-[…]
-If $q_{1}=q_{2}$, then any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ would satisfy (13). Naturally, identification is not possible in this case, which leads to our first condition.
-Non-systematic missing pattern: $q_{1}\neq q_{2}$.
-Condition (C1) precludes the special case of down-sampling, which is notoriously hard to be treated as a missing data problem.
-However, (C1) is inadequate for identification because the equations in (13) are not linearly independent.
-The lack of uniqueness of the solutions to (13) is at the crux of the identification problem in imputation.
-The blue lines in Figure 1 plots the set of solutions to (13).
-Each point on the blue line represents a legitimate TWI imputation whose marginal distributions before and after time $n_{1}$ are the same.
-Nonetheless, (12) hints an additional requirement for the correct imputation, giving rise to the following condition.
+$\left.\begin{aligned} \frac{1}{k_{1}}a_{1}-\frac{1}{k_{2}}a_{2}=&(1-q)(\frac{1}{k_{1}}-\frac{1}{k_{2}}),\\ \frac{1}{k_{1}}b_{1}-\frac{1}{k_{2}}b_{2}=&p(\frac{1}{k_{1}}-\frac{1}{k_{2}}).\end{aligned}\right\}$ (13)
+Derivation of (13) is relegated to the Section S1 of the supplementary material. Now we introduce two conditions.
+Non-systematic missing pattern: $k_{1}\neq k_{2}$.
 Interpolator stability: $a_{1}=a_{2}$, $b_{1}=b_{2}$.
-Intuitively, (C2) requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$.
-Together, (C1) and (C2) imply a unique solution to the system (13) and are sufficient for identification.
-Figure 1: Set of calibrations that correspond to temporal Wasserstein imputation solutions (blue line) and the correct calibrations $(a_{1}^{*},a_{2}^{*})$ and $(b_{1}^{*},b_{2}^{*})$ (red dots).
-The dashed orange line is the 45 degree line.
-$q_{1}=3$ and $q_{2}=4$ are used to plot this figure.
+Clearly, identification is not possible when $k_{1}=k_{2}$, in which case any imputation with $a_{1}=a_{2}$ and $b_{1}=b_{2}$ would satisfy (13), so (C1) is necessary for identification.
+Intuitively, (C1) precludes the special case of down-sampling, which is notoriously hard to be treated as a missing data problem.
+However, even with (C1) the solutions to (13) are not unique.
+This lack of uniqueness is at the crux of the identification problem for time series in the distribution matching framework (Muzellec et al.,, 2020; Wang et al.,, 2025).
+There are many imputations whose marginal distributions before and after $n_{1}$ are the same, but they may not equal to the one given in (9).
+Inspired by (12), (C2) imposes an additional requirement for the ideal imputation, which can be verified without knowledge about the missing data.
+It requires that, given the same neighboring observations, one should impute the missing entries in the same way (at least probabilistically) before and after $n_{1}$.
+Together, (C1) and (C2) imply a unique solution to the system (13), which guarantees identification.
 Proposition 6.
-Under (C1) and (C2), the only solution to (13) is
-$a_{1}=a_{2}=a_{1}^{*},\quad b_{1}=b_{2}=b_{1}^{*}.$
+Under (C1) and (C2), the only solution to (13) is $a_{1}=a_{2}=a_{1}^{*}$, $b_{1}=b_{2}=b_{1}^{*}$.
 That is, the only 2-dimensional marginal distribution that achieves zero Wasserstein loss and satisfies (C2) is that of the underlying distribution.
-As shown in Figure 1, the 45 degree line (dashed orange), which imposes (C2), intersects with the solution set of temporal Wasserstein imputations at the correct calibrations.
-This example shows, under two natural requirements, TWI can achieve distributional identification through minimizing the Wasserstein distance between the marginal distributions.
-In practice, appropriate tuning of $n_{1}$ and $p$ is needed for TWI to find an imputation with marginal distributions approximating the underlying ones.
 As shown in Proposition 6, it is not necessary to use an $n_{1}$ such that all missing data lie on the same side.
-Instead, we can select $n_{1}$ from a pre-specified interval $[t_{1},t_{2}]$ which yields the smallest Wasserstein loss, where $t_{2}-t_{1}$, $t_{2}$, and $n-t_{1}$ are sufficiently large.
-Alternatively, a simple procedure to promote interpolator stability, as suggested by (C2), is to run TWI with different cut-off $n_{1}$, using previously obtained imputations as initializations.
+In practice, appropriate tuning of $n_{1}$ and $p$ is needed.
+We can select $n_{1}$ from a pre-specified interval $t_{1},t_{2}]$ which yields the smallest Wasserstein loss, where $t_{1}$ and $n-t_{2}$ are sufficiently large.
+Alternatively, a simple procedure to promote interpolator stability is to run TWI with different splitting cut-offs, using previously obtained imputations as initializations.
 This procedure, which we call $k$-TWI, is summarized in Algorithm 2.
-Finally, as a nonparametric method, $p$ can be chosen to diverge with the sample size with $p/n\rightarrow 0$.
+Finally, similar to other nonparametric methods, $p$ can be chosen to diverge with the sample size so long as $p/n\rightarrow 0$.
 Input: Cut-off points $n_{1},n_{2},\ldots,n_{k}$, number of lags $p$
@@
 5 Simulation Studies
-In this section, we apply TWI to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its performance.
-For comparison, we employ imputation methods such as Kalman smoothing, the scalar filtering algorithm of Peña and Tsay, (2021), and deterministic methods such as linear and spline interpolations.
-In addition, the following two missing patterns are considered.
+In this section, we apply TWI to data generated from various time series models, including multivariate, nonlinear, and nonstationary ones, to examine its finite sample performance.
+For comparison, we employ Kalman smoothing, the iterative gap filling by SSA (Kondrashov and Ghil,, 2006), the scalar filtering algorithm (Peña and Tsay,, 2021), and linear interpolations.
+The following two missing patterns are considered:
   • Pattern I: 300 observations are omitted at random;
   • Pattern II: For every 20 observations, 6 consecutive observations are omitted.
-Pattern I simulates sporadic missing data while Pattern II generates patches of missing data.
+Pattern I simulates data missing at random while Pattern II generates patches of missing data within the same positions of every 20 observations.
 The total length of the time series is set to $n=1000$.
 Thus, for both missing patterns, 30% of the data are missing.
-For each data generating process and missing pattern considered, 1000 Monte Carlo simulations are carried out.
-We employ the proximal gradient descent (Parikh et al.,, 2014) to efficiently solve step (b) in Algorithm 1 when no side information is available since the projection step is straightforward.
-Conversely, when $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (6).
+For each model and missing pattern considered, 1000 Monte Carlo simulations are carried out.
+The proximal gradient descent (Parikh et al.,, 2014) is used to efficiently solve step (b) in Algorithm 1 when no side information is available because the projection step is straightforward.
+When $\mathcal{C}$ is defined by a system of linear equations, we solve step (b) using the closed form (6).
+The POT package (Flamary et al.,, 2021) in Python is used to solve the discrete optimal transport problem.
 Throughout we fix $n_{1}=0.4n$ and $p=6$ for TWI.
-In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points ($n_{1}\in\{0.25n,0.5n,0.75n\}$) to the simulated data.
+In addition, we apply the $k$-TWI (Algorithm 2) with three cut-off points $n_{1}\in\{0.25n,0.5n,0.75n\}$.
 For the initializations of TWI, both linear interpolation and Kalman smoothing are used in our experiments.
-Kalman smoothing and the deterministic smoothing techniques are implemented via the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R.
-The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, and can be applied to multivariate and nonstationary series.
-Its implementation details are documented in Section S2 of the supplementary material.
+Kalman smoothing and gap filling with iterative SSA (iSSA henceforth) are implemented, respectively, via the imputeTS (Moritz and Bartz-Beielstein,, 2017) and Rssa (Golyandina et al.,, 2018) packages in R.
+As discussed in Section 3.3, the window length $L$ for which iSSA can be implemented is restricted.
+Here we use $L=6$ for iSSA.
+In the reconstruction stage for iSSA, the first three eigentriples are used for grouping with the linear interpolation as initialization.
+The scalar filter algorithm (denoted as ScalarF henceforth) utilizes intervention analysis, which is detailed in Section S2 of the supplementary material.
+Codes for the numerical experiments in this and the next section can be found in the author’s Github page: https://github.com/shuochieh/Wasserstein_imputation.
 5.1 Linear and nonlinear univariate time series
-In this subsection, we consider the following data-generating processes (DGPs) for univariate time series, which can exhibit linear or nonlinear dynamic patterns.
-In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$ for $j=1,2,\ldots$ denote i.i.d.
+We first consider the following data-generating processes (DGPs).
+In the sequel, $\{\epsilon_{t}\}$ and $\{\epsilon_{t,j}\}$, for $j=1,2,\ldots$, denote i.i.d.
 sequences of standard Gaussian variables.
@@
 DGP 3 (TAR; Tong,, 1983).
-$x_{t}=\left\{\begin{array}[]{cc}\phi_{1}x_{t-1}+\epsilon_{t},&x_{t-1}\leq\tau\\ \phi_{2}x_{t-1}+0.5\epsilon_{t},&x_{t-1}>\tau\end{array}\right.,$
-where $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$
+If $x_{t-1}\leq\tau$, then $x_{t}=\phi_{1}x_{t-1}+\epsilon_{t}$. Otherwise, $x_{t}=\phi_{2}x_{t-1}+0.5\epsilon_{t}$. The parameters $(\phi_{1},\phi_{2},\tau)=(-2,0.7,1)$.
 DGP 4 (I(1) process).
@@
 $x_{t}=10\cos(t(0.23\pi))+6\cos(t(0.17\pi))+0.5\epsilon_{t}$.
-DGPs 1 and 2 are well-known linear time series models.
+DGPs 1 and 2 are well-known stochastic time series models.
 Note that DGP 2 admits an AR($\infty$) representation, so any finite-dimensional marginal distribution is inadequate to fully characterize the distribution of the underlying process.
@@
 Three measures of performance are considered.
-First, we report the Wasserstein distance (of order 2) between the empirical 3-dimensional marginal distributions of the imputed series and the full data.
-Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{x_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$.
+First, we report the Wasserstein distance between the empirical 3-dimensional marginal distributions of the imputed series and the full data.
+Specifically, let $\{\hat{w}_{t}^{(i)}\}$ and $\{\tilde{x}_{t}^{(i)}\}$ be the imputed series and the original series in simulation $i$, respectively, for $i=1,2,\ldots,1000$.
 We report
-$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{({x}_{t}^{(i)},{x}_{t-1}^{(i)},{x}_{t-2}^{(i)})}\right),$ (14)
+$\frac{1}{1000}\sum_{i=1}^{1000}\mathcal{W}_{2}\left(\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\hat{w}_{t}^{(i)},\hat{w}_{t-1}^{(i)},\hat{w}_{t-2}^{(i)})},\frac{1}{n-3}\sum_{t=2}^{n-1}\delta_{(\tilde{x}_{t}^{(i)},\tilde{x}_{t-1}^{(i)},\tilde{x}_{t-2}^{(i)})}\right),$ (14)
 as a measure of how well the imputed series approximates the original series.
-Second, for DGP 14, the model parameters are estimated using the imputed series, and their root mean square errors (RMSEs) are reported.
+Second, for DGP 14, the model parameters are estimated using the imputed series, and the estimation root mean square errors (RMSEs) are reported.
 Third, we compare the RMSEs of the estimated autocovariance functions (ACFs).
-However, due to space constraints, we relegate the comparison of ACFs to Section S3.1 in the supplementary material.
+However, due to space constraints, the results for the ACFs are moved to Section S3.1 in the supplementary material.
 The Wasserstein distance between the imputed and the original marginal distributions is reported in Table 1.
 Across all five DGPs employed, the smallest loss is consistently achieved by TWI-type methods, indicating that TWI effectively approximates the distributional properties of the underlying series.
-While other methods may yield good imputations for some specific DGPs, the result is not consistent across all processes.
-For instance, linear and spline interpolations show low Wasserstein loss for DGP 1, but their performance is less competitive when applied to other models.
+While other methods may yield good imputations for some specific DGPs, they are not consistent across all processes.
+For instance, linear interpolations show low Wasserstein loss for DGP 1 (AR), but it is less competitive when applied to other models.
+Similarly, Kalman smoothing and iSSA work well for DGP 5 (CYC), but both yield large Wasserstein loss for DGP 3 (TAR).
 In contrast, TWI methods, carefully initialized, consistently deliver solid performance across models and missing patterns.
-For the nonstationary Model 4 (I(1)), we first obtain an imputation $\{\hat{w}_{t}\}$ and consider the differenced series $\nabla\hat{w}_{t}=\hat{w}_{t}-\hat{w}_{t-1}$.
-Table 1 shows the Wasserstein distance between the marginal distributions implied by $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$.
-In this case, TWI consistently yields the lowest Wasserstein loss and greatly improves from its initialization.
-Table 1 also shows TWI typically improves the Wasserstein loss from its initialization, often significantly.
+For the nonstationary Model 4 (I(1)), Table 1 shows the Wasserstein distance between the marginal distributions of the differenced series $\{\nabla\hat{w}_{t}\}$ and $\{\nabla x_{t}\}$.
+TWI again yields the lowest Wasserstein loss.
+Table 1 also shows TWI often significantly improves the Wasserstein loss from its initialization.
 For example, for DGP 3 (TAR), TWI reduces the Wasserstein loss from Kalman smoothing initialization by approximately 45% under missing pattern I (from 1.38 to 0.74).
 Similarly, for DGP 5 (CYC), TWI reduces the Wasserstein loss by more than 50% from linear initialization (from 1.96 to 0.79).
-Figure 2 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed under DGP 3, alongside $(x_{t},x_{t-1})$ of the original data.
-While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the ground truth.
-Note that Kalman smoothing is done via an ARIMA state-space representation; hence the model is misspecified if the underlying process is nonlinear.
-Nevertheless, it already serves as a good initialization of TWI.
+Figure 1 plots $\hat{w}_{t}$ against $\hat{w}_{t-1}$ when either or both are imputed under DGP 3, alongside $(\tilde{x}_{t},\tilde{x}_{t-1})$ of the original data.
+While Kalman smoothing introduces substantial distortion in the marginal distribution, TWI, initialized by Kalman smoothing, significantly corrects this bias, producing imputations that capture the nonlinearity in the data generating process.
+Note that Kalman smoothing is based on an ARIMA state-space representation; hence the model is misspecified.
+Nevertheless, it already serves as a good initialization for TWI.
 Table 1: Wasserstein distance between the empirical marginal distributions of the imputed series and those of the full data, averaged over 1000 simulations.
+iSSA denotes gap filling by the iterative singular spectrum analysis.
 TWIlin denotes temporal Wasserstein imputation using linear interpolation as initialization and TWIKal denotes temporal Wasserstein imputation with Kalman smoothing as initialization.
@@
 The smallest value for each model and missing pattern is marked in boldface.
-Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
+Model Linear iSSA Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
 Missing pattern I
-AR 0.41 0.41 0.42 0.50 0.40 0.44 0.41 0.44
-ARMA 0.48 0.58 0.47 0.48 0.40 0.38 0.40 0.38
-TAR 1.12 1.51 1.38 1.13 0.96 0.81 0.74 0.63
-I(1) 0.71 0.75 0.62 4.38 0.58 0.53 0.54 0.50
-CYC 1.96 0.82 0.77 1.97 0.79 0.77 0.60 0.70
-NLVAR 2.98 3.42 3.01 2.94 2.25 2.14 2.19 2.12
-AL (×$\times 10$) 0.70 1.00 4.17 0.55 0.43 0.37 0.38 0.35
+AR 0.41 0.40 0.42 0.50 0.40 0.44 0.41 0.44
+ARMA 0.48 0.51 0.47 0.48 0.40 0.38 0.40 0.38
+TAR 1.12 1.29 1.38 1.13 0.96 0.81 0.74 0.63
+I(1) 0.71 0.68 0.62 4.38 0.58 0.53 0.54 0.50
+CYC 1.96 0.61 0.77 1.97 0.79 0.77 0.60 0.70
+NLVAR 2.98 3.04 3.01 2.94 2.25 2.14 2.19 2.12
+AL ($\times 10$) 0.70 0.85 4.17 0.55 0.43 0.37 0.38 0.35
 Missing pattern II
-AR 0.44 0.59 0.50 0.76 0.39 0.44 0.43 0.45
-ARMA 0.47 1.29 0.51 0.53 0.36 0.34 0.40 0.35
-TAR 1.04 3.11 1.25 0.99 0.84 0.73 0.76 0.61
-I(1) 0.67 0.65 0.60 8.35 0.51 0.45 0.49 0.44
-CYC 2.58 1.81 0.77 3.79 2.62 1.60 0.62 0.72
-NLVAR 3.00 6.07 3.33 3.67 2.13 2.02 2.31 2.04
-AL (×$\times 10$) 0.56 1.83 3.81 0.46 0.36 0.33 0.40 0.34
+AR 0.44 0.43 0.50 0.76 0.39 0.44 0.43 0.45
+ARMA 0.47 0.69 0.51 0.53 0.36 0.34 0.40 0.35
+TAR 1.04 1.58 1.25 0.99 0.84 0.73 0.76 0.61
+I(1) 0.67 0.60 0.60 8.35 0.51 0.45 0.49 0.44
+CYC 2.58 1.47 0.77 3.79 2.62 1.60 0.62 0.72
+NLVAR 3.00 3.17 3.33 3.67 2.13 2.02 2.31 2.04
+AL ($\times 10$) 0.56 1.04 3.81 0.46 0.36 0.33 0.40 0.34
 (a) Under Missing pattern I
 (b) Under Missing pattern II
-Figure 2: Scatter plot of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
+Figure 1: Scatter plots of $(x_{t-1},x_{t})$ of the original data and $(\hat{w}_{t-1},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 3 (TAR).
 Next, we turn to the model parameters estimated from the imputed series.
@@
 These estimation results demonstrate TWI’s strength in learning the underlying nonlinear dynamics.
-For DGP 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman filter being a strong initialization.
-Table 2: Estimation root mean square errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
-Model Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
+For DGP 4 (I(1)), TWI either outperforms or stays on par with the best benchmark, with Kalman smoothing being a strong initialization.
+Table 2: Estimation root mean squared errors for parameter estimation errors. The results are averaged over 1000 simulations. See Table 1 for notations.
+Model Linear iSSA Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
 Missing pattern I
 AR $\phi$ 0.06 0.06 0.05 0.03 0.03 0.04 0.03 0.04
-ARMA $\phi_{1}$ 0.19 0.47 0.12 0.08 0.17 0.16 0.11 0.11
-$\phi_{2}$ 0.46 0.95 0.18 0.10 0.33 0.26 0.16 0.15
-TAR $\phi_{1}$ 1.02 1.67 0.82 0.63 0.71 0.58 0.53 0.44
-$\phi_{2}$ 0.08 0.07 0.20 0.04 0.05 0.03 0.03 0.02
-$\tau$ 0.01 0.75 0.35 0.20 0.05 0.05 0.03 0.03
-I(1) $\phi$ 0.96 0.47 0.45 0.68 0.48 0.32 0.30 0.24
-NLVAR $\phi_{11}$ 0.19 0.27 0.15 0.03 0.06 0.03 0.05 0.03
-$\phi_{12}$ 3.10 3.16 2.87 1.67 1.34 1.02 1.22 0.97
+ARMA $\phi_{1}$ 0.19 0.28 0.12 0.08 0.17 0.16 0.11 0.11
+$\phi_{2}$ 0.46 0.63 0.18 0.10 0.33 0.26 0.16 0.15
+TAR $\phi_{1}$ 1.02 1.41 0.82 0.63 0.71 0.58 0.53 0.44
+$\phi_{2}$ 0.07 0.07 0.20 0.04 0.05 0.03 0.03 0.02
+$\tau$ 0.01 0.26 0.35 0.20 0.05 0.05 0.03 0.03
+I(1) $\phi$ 0.96 1.03 0.45 0.68 0.48 0.32 0.30 0.24
+NLVAR $\phi_{11}$ 0.19 0.11 0.15 0.03 0.06 0.03 0.05 0.03
+$\phi_{12}$ 3.10 2.43 2.87 1.67 1.34 1.02 1.22 0.97
 $\phi_{21}$ 0.03 0.07 0.03 0.03 0.03 0.03 0.03 0.03
-$\phi_{21}$ 0.18 0.23 0.10 0.04 0.05 0.04 0.04 0.04
+$\phi_{22}$ 0.18 0.09 0.10 0.04 0.05 0.04 0.04 0.04
 Missing pattern II
-AR $\phi$ 0.05 0.07 0.04 0.04 0.03 0.02 0.03 0.02
-ARMA $\phi_{1}$ 0.04 0.08 0.09 0.08 0.05 0.06 0.09 0.09
-$\phi_{2}$ 0.16 0.61 0.11 0.10 0.12 0.11 0.11 0.12
-TAR $\phi_{1}$ 0.72 2.44 0.38 0.25 0.25 0.20 0.29 0.21
-$\phi_{2}$ 0.09 0.12 0.05 0.03 0.06 0.05 0.02 0.01
-$\tau$ 0.01 1.72 0.01 0.01 0.01 0.01 0.01 0.01
-I(1) $\phi$ 0.08 1.08 0.11 0.70 0.09 0.10 0.10 0.10
-NLVAR $\phi_{11}$ 0.19 0.44 0.08 0.02 0.02 0.01 0.02 0.01
-$\phi_{12}$ 2.67 3.51 1.68 1.26 0.61 0.46 0.61 0.43
+AR $\phi$ 0.05 0.06 0.04 0.04 0.03 0.02 0.03 0.02
+ARMA $\phi_{1}$ 0.03 0.11 0.09 0.08 0.05 0.06 0.09 0.09
+$\phi_{2}$ 0.16 0.46 0.11 0.10 0.12 0.11 0.11 0.12
+TAR $\phi_{1}$ 0.72 1.66 0.38 0.25 0.25 0.20 0.29 0.21
+$\phi_{2}$ 0.09 0.10 0.05 0.03 0.06 0.05 0.02 0.01
+$\tau$ 0.01 0.34 0.01 0.01 0.01 0.01 0.01 0.01
+I(1) $\phi$ 0.08 0.53 0.11 0.70 0.09 0.10 0.10 0.10
+NLVAR $\phi_{11}$ 0.19 0.04 0.08 0.02 0.02 0.01 0.02 0.01
+$\phi_{12}$ 2.67 1.49 1.68 1.26 0.61 0.46 0.61 0.43
 $\phi_{21}$ 0.02 0.09 0.03 0.03 0.06 0.05 0.04 0.03
-$\phi_{21}$ 0.15 0.34 0.05 0.04 0.04 0.04 0.04 0.04
+$\phi_{22}$ 0.15 0.11 0.05 0.04 0.04 0.04 0.04 0.04
 5.2 Multivariate time series
-In this subsection, we consider multivariate nonlinear time series models.
+We consider next models for multivariate nonlinear time series.
 DGP 6 (Nonlinear VAR; NLVAR).
-$x_{t,1}=\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1}\quad x_{t,2}=\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2}$
-where $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ is the centered sigmoid function, and $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
+Let $\mathfrak{s}(z)=1/(1+\exp(-z))-0.5$ be the centered sigmoid function. The data are generated as
+$x_{t,1}=\phi_{11}x_{t-1,1}+\phi_{12}\mathfrak{s}(3x_{t-1,2})+0.25\epsilon_{t,1},\quad x_{t,2}=\phi_{21}x_{t-1,1}+\phi_{22}x_{t-1,2}+3\epsilon_{t,2},$
+where $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})=(0.3,8,0,0.4)$.
 DGP 6 is a VAR model with a nonlinear term.
@@
 This is validated by the results shown in Table 1, where TWI achieves the lowest Wasserstein distance.
-Next, we assess the effectiveness of the imputed series in estimating the model parameters $(\phi_{11},\phi_{12},\phi_{21},\phi_{22})$, assuming the nonlinear functional form is known.
-The root mean square estimation errors are documented in Table 2.
-Compared to the benchmarks, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$.
+Moreover, as shown in Table 2, TWI provides considerable error reduction in the parameter associated with the nonlinear effect, $\phi_{12}$, compared to the benchmarks.
 It attains more than 50% reduction in RMSE compared to the scalar filter, the best-performing benchmark.
@@
 The data are generated from the additive logistic (AL) model of Brunsdon and Smith, (1998).
-Due to space constraints, we present its simulation results and discussion in Section S3.2 of the supplementary material.
+Due to space constraints, we present the simulation results and discussion in Section S3.2 of the supplementary material.
 (a) $\hat{w}_{t,1}$ on $\hat{w}_{t-1,1}$
@@
 (d) $\hat{w}_{t,2}$ on $\hat{w}_{t-1,2}$
-Figure 3: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 under missing pattern II.
+Figure 2: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series for $i,j\in\{1,2\}$, when the data are generated from Model 6 under missing pattern $I$.
 6 Groundwater data application
-In this section, we apply the proposed TWI to a groundwater dataset from Taiwan, which was collected and analyzed by Hsu et al., (2020).
-Given the island’s steep topography and varying rainfall distribution, groundwater management is of considerable importance for Taiwan, particularly as it is a critical water source.
-However, the statistical analysis of groundwater levels is challenged by frequent missing data, especially from deeper wells.
-We aim to study the seasonal groundwater variations, for which we first transform the hourly data from 352 sites between October 1992 and August 2020 into monthly records.
-For each site, monthly groundwater level is calculated as the average of the observations in a given month, with months classified as missing if more than one-third of the observations are missing.
-Our analysis focuses on sites with a less-than-median amount of missing data, reducing the sample to 176 sites.
-Figure 4 shows the monthly groundwater levels for these sites, after subtracting the mean and normalizing by the standard deviation.
-We make two observations.
-First, many series exhibit strong seasonality due to the rainfall pattern in Taiwan.
-Monsoon and typhoons increase rainfall in summer, while winters are drier.
-Second, there are many outlying values and anomalies in the observed series, including abrupt changes in groundwater levels that are more than 5 standard deviations from the sample mean.
+In this section, we apply the proposed TWI to a Taiwan groundwater dataset, which was collected and analyzed by Hsu et al., (2020).
+Because of the island’s steep topography and uneven rainfall distribution, groundwater management is of considerable importance, particularly as it is a critical water source.
+However, the statistical analysis of groundwater levels is plagued by frequent missing data, especially from deeper wells.
+To study the seasonal groundwater variations, we first transform the hourly data between October 1992 and August 2020 into monthly records.
+For each site, monthly groundwater level is calculated as the average in a given month, with months classified as missing if more than one-third of the observations are missing.
+There are many outlying values in the series, including abrupt changes in groundwater levels that are more than 5 standard deviations away from the sample mean.
 We also treat such extreme values as missing data.
-Figure 4: Time plot of the groundwater levels of the 176 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
-During the period from October, 1992 to August, 2020, few months recorded valid observations at all sites.
-Figure 5 shows the fluctuations in the percentage of sites with missing data over time.
-Although missing data are mostly prevalent in the early years, there are some time during which the percentage surges, such as between 2003 and 2005.
-In addition, the percentage is gradually increasing in more recent years.
-Thus, one may lose much information if the analysis is restricted to sites with relatively complete records.
-We consider two downstream tasks using the imputed series.
-First, following Hsu et al., (2020), we decompose the de-trended time series using the independent component analysis (ICA), for which we apply the fastICA algorithm of Hyvärinen and Oja, (2000).
-Because direct application of ICA becomes infeasible when the series are severely corrupted by missing data, the analysis of Hsu et al., (2020) is confined to a relatively short time span.
-The second downstream task considers autoregressive modeling, which is presented in Section S3.3 in the supplementary material.
-In implementation, we chose $n_{1}=160$ and $p=6$ for TWI, and for $k$-TWI we use $n_{1}\in\{160,225,110\}$.
-Figure 5: Monthly percentage of sites recording a missing value.
-Figure 6 shows the leading two independent components (ICs) estimated using the imputed series.
-ICs estimated from series imputed by Kalman smoothing have peaks around August-September (red shaded area) and May-June (blue shaded area), which was also observed in Hsu et al., (2020).
-These peaks are attributed to the dual rain seasons in Taiwan, which involve typhoons in August-September and the Meiyu period from mid-May to mid-June.
-ICs estimated from TWI-imputed series with Kalman initialization largely preserve these seasonal peaks, but with improved performance before 2002.
-Recall that missing data are prevalent during this period (Figure 5).
-Especially for the first IC, those computed from TWI-imputed series have clear peaks around August-September with a more consistent amplitude.
-On the other hand, the ICs computed from the linearly interpolated series could not effectively isolate and identify the Meiyu season, with peaks occurring slightly earlier than expected.
-Nonetheless, ICs estimated from TWI-imputed series with linear interpolation substantially alleviates this shortcoming and produced more interpretable results.
-ICs computed from spline-interpolated series also identifies the two rain seasons, albeit they show more rugged trajectories instead of clean periodic patterns.
-For the scalar filter, the resulting ICs similarly appear noisy, and the Meiyu season is not well-isolated, showing that the imputation induces distortions in the dynamic pattern.
-Figure 6: Independent components estimated using the imputed series.
-The dashed red lines in each plot mark January 2002.
-The red and blue shaded areas correspond to August-September (typhoon season) and May-June (Meiyu season), respectively.
-Overall, TWI took a more holistic view which accounts for the underlying dynamics, such as complex seasonality, and produced interpretable results.
-In Section S3.3 of the supplementary material, we conduct further experiments to show that the marginal distributions of TWI-imputed series, in terms of AR modeling, are also more consistent across time compared to other methods.
-These results show TWI performs reliably in practice and is a useful method in handling missing data in time series.
+For demonstration, our analysis focuses on the sites with less than 10 missing values, reducing the sample to 15 sites.
+As shown in Figure 3, these series exhibit seasonality and structural changes, which are quite representative of the dataset and are common real-world features.
+Figure 3: Time plots of the groundwater levels of the 15 sites under study. Each series is normalized by subtracting its mean and dividing by its standard deviation.
+In the following, we randomly omit $\lfloor\rho n\rfloor$ observations to create missing data, where $\rho\in\{0.1,0.2,0.3,0.4,0.5,0.6\}$ is the missing ratio, $n=308$ is the number of observations (including missing values) in the period under study, and $\lfloor x\rfloor$ is the largest integer $\leq x$ for $x>0$.
+However, the last 36 observations of each series are left intact because otherwise, the SSA could not be implemented due to lack of intact columns in the trajectory matrix.
+Since TWI performs best with stationary series, we first use kernel smoothing (estimated with missing data) to remove the trends, and then apply ($k$-)TWI to the residuals.
+In addition to the methods employed in Section 5, we also compare with imputing the missing values using the sample mean.
+Due to space constraints, the implementation details of the methods are provided in Section S3.3 of the supplementary material.
+We consider two downstream tasks.
+First, the imputed series are used to estimate some summary statistics, including the autocovariances, autocorrelations, and partial autocorrelations.
+Specifically, let $\gamma_{j}(h)$ be the statistic at the $h$-th lag obtained using the $j$-th series (without the artificial missing data), and $\hat{\gamma}_{j}(h)$ be the corresponding estimates from the imputed series.
+Table 3 reports the mean absolute error (MAE), $\frac{1}{45}\sum_{j=1}^{15}\sum_{h=1}^{3}|\hat{\gamma}_{j}(h)-\gamma_{j}(h)|$, for the autocovariances.
+When the missing ratio is mild, for instance $\rho\leq 0.3$, iSSA, Kalman smoothing, and linear interpolation perform reasonable well, while TWI and $k$-TWI are quite close to the best-performing methods.
+Mean imputation significantly distorts the estimates.
+However, when $\rho>0.3$, TWI and $k$-TWI show some advantages.
+This is expected because when the missing ratio is high, accurate estimation of the missing data is difficult, but distribution matching is still possible.
+Table 4 shows the MAE for estimating the autocorrelations.
+For $\rho\leq 0.2$, most methods yield similar estimates.
+When $\rho>0.2$, $k$-TWI outperforms the competing methods, among which the Kalman smoothing is a strong contender.
+Meanwhile, TWI is somewhat close to Kalman smoothing for $\rho\leq 0.4$.
+Linearly interpolated series, however, has autocorrelations quite different from the original series.
+Finally, Table 5 shows the MAE for the partial autocorrelations.
+For $\rho\geq 0.3$, TWI and $k$-TWI consistently show good performance.
+The MAE of Kalman smoothing is similar to the proposed methods only when $\rho\leq 0.4$.
+Unlike in previous cases, the partial autocorrelations of the iSSA imputations are inaccurate, showing that it often neglects stochastic dynamic relationships.
+Table 3: Mean absolute errors ($\times 10$) for estimating autocovariances. Mean stands for imputation by the sample mean. For each missing ratio $\rho$, the smallest two values are in boldface.
+$\rho$ Linear iSSA Kalman ScalarF Mean TWI $k$-TWI
+0.1 0.14 0.16 0.14 0.31 0.98 0.16 0.18
+0.2 0.21 0.20 0.38 0.72 1.91 0.27 0.34
+0.3 0.41 0.31 0.30 1.01 2.46 0.35 0.33
+0.4 0.50 0.49 0.43 1.58 3.10 0.46 0.46
+0.5 0.77 0.63 0.60 2.35 3.54 0.56 0.52
+0.6 0.98 0.91 1.25 2.90 3.84 0.99 0.79
+Table 4: Mean absolute errors ($\times 10$) for autocorrelations. See Table 3 for notations.
+$\rho$ Linear iSSA Kalman ScalarF Mean TWI $k$-TWI
+0.1 0.18 0.17 0.12 0.15 0.63 0.13 0.14
+0.2 0.27 0.25 0.28 0.35 1.37 0.23 0.26
+0.3 0.55 0.40 0.36 0.51 1.76 0.31 0.28
+0.4 0.73 0.47 0.42 0.72 2.37 0.45 0.38
+0.5 1.17 0.72 0.56 1.22 2.85 0.65 0.44
+0.6 1.52 0.92 1.09 1.81 3.17 1.23 0.76
+Table 5: Mean absolute errors ($\times 10$) for the partial autocorrelations. See Table 3 for notations.
+$\rho$ Linear iSSA Kalman ScalarF Mean TWI $k$-TWI
+0.1 1.10 1.13 1.11 1.34 2.27 1.14 1.27
+0.2 1.09 1.26 1.21 1.72 2.95 1.39 1.47
+0.3 1.26 1.60 1.50 1.91 3.19 1.34 1.34
+0.4 1.38 2.00 1.54 2.01 3.60 1.51 1.55
+0.5 1.48 1.76 1.88 2.24 3.81 1.47 1.46
+0.6 1.64 2.20 2.21 2.85 4.07 1.60 1.58
+Second, we use the first 200 observations from the imputed series to estimate an AR(15) model, where the estimated coefficients are then used in predicting, in a one-step-ahead rolling window fashion, the rest 108 observations from the true data.
+Figure 4 plots the root mean squared prediction errors.
+For most series except Series 1 and 2, TWI and $k$-TWI are among the methods yielding the lowest prediction errors.
+Their advantage is again more pronounced when the missing ratio is large.
+In addition, the TWI-type methods are stable as $\rho$ increases, whereas the iSSA and Kalman smoothing may suffer from large errors.
+In summary, the above analysis shows that the proposed TWI and $k$-TWI methods, through distribution matching, are useful in imputing missing values of time series for downstream statistical analysis.
+Compared to the existing nonparametric iSSA method and the parametric Kalman smoothing, such advantage is especially valuable when the missing ratio is high.
+Figure 4: Root mean squared prediction errors for each series.
+The AR(15) model is trained using the first 200 observations of the imputed series, which is then tested on the rest 108 data from the uncontaminated true series (excluding missing values).
+7 Conclusion
+This paper introduced a temporal Wasserstein imputation, a novel nonparametric method for imputing missing data in time series.
+Compared to existing approaches, TWI (1) does not require any model specification prior to imputation; (2) can be carried out in a principled manner, and (3) yields imputations whose distributional properties are more aligned with those of the underlying process, making downstream analyses more reliable.
+In simulation study and real data analysis, TWI has shown its versatility and contributions to the existing literature, especially for the cases of high missing ratio and of imputing stochastic, nonlinear time series.
 SUPPLEMENTARY MATERIAL
 Supplementary material:
-The file contains the proofs of the propositions presented in Section 4 and some technical results.
-It also documents the scalar filter algorithm of Peña and Tsay, (2021) and complementary numerical results, including the simulation results for autocovariance estimation and for data generated from a compositional time series model.
-Finally, it includes the analysis of the groundwater data via autoregressive modeling.
-Supplementary material to “Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series”
+The file contains the proofs for the results presented in Section 4 and some technical results.
+It also documents the scalar filter algorithm of Peña and Tsay, (2021) and complementary numerical results, including the simulation results for autocovariance estimation and for a compositional time series DGP.
+Finally, it includes the implementation details of the groundwater data analysis.
+Supplementary material to “Temporal Wasserstein Imputation: A Versatile Method for Time Series Imputation”
+Shuo-Chieh Huang, Tengyuan Liang, and Ruey S. Tsay
 This supplementary material contains three sections.
@@
 Section S3 contains complementary numerical results, including the simulation results for autocovariance estimation and imputing data generated from a compositional time series model.
-It also includes the analysis of autoregressive modeling using the imputed groundwater data.
+It also includes the details of implementation for the methods employed in the groundwater data analysis.
 S1 Technical proofs
@@
 $\frac{\partial^{2}F_{0}}{\partial\mathbf{u}\partial\mathbf{v}}=$ $-\left(\pi_{ij}\ell^{\prime\prime}(w_{i}-w_{j})\right)_{p-1\leq i\leq n_{1},n_{1}+1\leq j\leq n-1}.$ (S1.3)
-Since $\ell$ is convex, ′′$\ell^{\prime\prime}\geq 0$.
+Since $\ell$ is convex, $\ell^{\prime\prime}\geq 0$.
 From (S1.1)–(S1.3), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries.
@@
 ∎
-Before proving Proposition 4, we first state our assumptions and establish Lemma S6.
-The underlying time series $\{\tilde{x}_{t}\}$ is stationary and ergodic.
-$\mathbb{E}|\tilde{x}_{t}|^{k}<\infty$ for some $k\geq 1$.
-These assumptions are satisfied by many linear and nonlinear time series models.
+In order to prove Proposition 4, we need the following lemma.
 Lemma S7.
@@
 with probability one. Let
-=$\mathcal{E}=\{\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q})\mbox{ for all }\mathbf{q}\in\mathbb{Q}^{p}\}\cap\left\{\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|\right\},$
-noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2).
-By the Ergodic theorem and (S1.6), $\mathbb{P}(\mathcal{E})=1$.
-Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary.
-Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
+[…]
+[…]
 $|\varphi(\mathbf{h})-\varphi(\mathbf{q})|<\epsilon,$ (S1.7)
 for all $\mathbf{q}\in\mathbb{Q}^{p}$ with $\|\mathbf{h}-\mathbf{q}\|<\delta$. In addition, using $|\exp(\mathrm{i}\mathbf{h}^{\top}\mathbf{z})-1|\leq\|\mathbf{h}\|\|\mathbf{z}\|$, we have
-$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (S1.8)
+[…]
 for all $\mathbf{h}$, $\mathbf{q}$ such that $\|\mathbf{h}-\mathbf{q}\|<\delta$. Combining (S1.7)–(S1.8), we have
@@
 Now the desired result follows from Lemma S7. ∎
+In the rest of this section, the derivation of (13) in Section 4.2 is provided.
+Proof of (13) in Section 4.2.
+From (9)–(11) in the main text, we can compute the marginal distributions implied by the imputed series, given as follows. If $t<n_{1}$,
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))=$ $(1-\frac{1}{k_{1}})\lambda_{2}(1-q)+\frac{1}{k_{1}}\lambda_{2}a_{1},$
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))=$ $(1-\frac{1}{k_{1}})\lambda_{1}p+\frac{1}{k_{1}}\lambda_{1}b_{1},$
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))=$ $(1-\frac{1}{k_{1}})\lambda_{2}q+\frac{1}{k_{1}}\lambda_{2}(1-a_{1}),$
+$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))=$ $(1-\frac{1}{k_{1}})\lambda_{1}(1-p)+\frac{1}{k_{1}}\lambda_{1}(1-b_{1}),$
+and for $t\geq n_{1}$, the marginal distribution can be obtained by replacing $a_{1},b_{1},k_{1}$ by $a_{2},b_{2},k_{2}$ in the above.
+Since TWI equates the marginal distributions before and after $n_{1}$, we have
+$(1-\frac{1}{k_{1}})\lambda_{2}(1-q)+\frac{1}{k_{1}}\lambda_{2}a_{1}=$ $(1-\frac{1}{k_{2}})\lambda_{2}(1-q)+\frac{1}{k_{2}}\lambda_{2}a_{2},$
+$(1-\frac{1}{k_{1}})\lambda_{1}p+\frac{1}{k_{1}}\lambda_{1}b_{1}=$ $(1-\frac{1}{k_{2}})\lambda_{1}p+\frac{1}{k_{2}}\lambda_{1}b_{2},$
+$(1-\frac{1}{k_{1}})\lambda_{2}q+\frac{1}{k_{1}}\lambda_{2}(1-a_{1})=$ $(1-\frac{1}{k_{2}})\lambda_{2}q+\frac{1}{k_{2}}\lambda_{2}(1-a_{2}),$
+$(1-\frac{1}{k_{1}})\lambda_{1}(1-p)+\frac{1}{k_{1}}\lambda_{1}(1-b_{1})=$ $(1-\frac{1}{k_{2}})\lambda_{1}(1-p)+\frac{1}{k_{2}}\lambda_{1}(1-b_{2}).$
+After some algebraic manipulations, these equations simplify to (13) in the main text. ∎
 S2 Scalar filter algorithm
@@
 $\mathrm{RMSE}_{\gamma}(h)=\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}(\gamma(\hat{\mathbf{w}}^{(i)},h)-\gamma_{\mathbf{x}}(h))^{2}},$ (S3.9)
-where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(x_{0}^{(i)},\ldots,x_{n-1}^{(i)})$.
+where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(\tilde{x}_{0}^{(i)},\ldots,\tilde{x}_{n-1}^{(i)})$.
 Table S1 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$.
 For Model 1 (AR), most methods perform reasonably well.
-Linear interpolation produces accurate ACF estimates since the positive AR coefficient intrinsically favors linearity, while the scalar filter exhibits largest estimation errors.
+Linear interpolation produces accurate ACF estimates since the positive AR coefficient aligns well with linearity, while the scalar filter exhibits largest estimation errors.
 For Model 2 (ARMA), TWI improves from linear and Kalman initializations and yields more accurate ACF estimates: TWI reduces the biases in Kalman smoothing when $h=0$ and in linear interpolation when $h=1$, and $h=2$.
-For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series exhibit the lowest RMSEs, with substantial improvements over the initializations.
+For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series have RMSEs that are among the lowest, with substantial improvements over the initializations.
 Notably, the scalar filter also yields accurate ACFs for $h=1,2$.
 Figure S1 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
-The advantage of TWI becomes even more pronounced in nonstationary models 4 (I(1)) and 5 (CYC), for which it achieves low ACF estimation errors.
+For DGP 4 (I(1)) and DGP 5 (CYC), TWI again offers great improvement over initializations.
+Noe that for DGP 5 (CYC), which generates a structural time series with deterministic cycles, iSSA is very effective at least under missing pattern I.
+However, perhaps due to the small window length caused by the high missing ratio, the performance of iSSA under missing pattern II is severely compromised (see also Figure S2, which shows TWI captures the underlying dynamic well).
 The scalar filter performs poorly for these models, which is likely the consequence of using many indicators in the intervention analysis.
-Among the benchmarks, Kalman filter produces the best result for these two models, but it can be further improved by the proposed TWI procedure.
+Among the benchmarks, Kalman smoothing produces the best result for these two models, which can benefit from further improvement provided by the proposed TWI procedure.
 (a) Missing pattern I
@@
 Figure S1: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR).
-The blue line is the smoothed density estimated from the histogram of the ACFs estimated from the ground truth.
+The blue line is the smoothed density of the histogram of the ACFs estimated from the ground truth.
+(a) $h=1$
+(b) $h=2$
+(c) $h=6$
+(d) $h=10$
+Figure S2: Scatter plots of $(\tilde{x}_{t-h},\tilde{x}_{t})$ of the original data and $(\hat{w}_{t-h},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 5 (CYC) under missing pattern $I$.
+$h\in\{1,2,6,10\}$.
 Table S1: $\mathrm{RMSE}_{\gamma}(h)$, defined in (S3.9), for $h=0,1,2$, with $h$ denoting lag.
@@
 See Table 1 in the main text for notations.
-Model $h$ Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
+Model $h$ Linear iSSA Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
 Missing pattern I
-AR 0 0.32 0.29 0.37 0.50 0.30 0.31 0.32 0.32
-1 0.26 0.34 0.29 0.47 0.27 0.30 0.28 0.31
-2 0.26 0.27 0.29 0.38 0.25 0.28 0.27 0.29
-ARMA 0 0.14 0.19 0.29 0.31 0.13 0.13 0.22 0.18
-1 0.22 0.44 0.07 0.08 0.14 0.10 0.07 0.07
-2 0.08 0.08 0.07 0.07 0.06 0.06 0.07 0.07
-TAR 0 0.59 0.77 0.73 1.07 0.65 0.56 0.56 0.45
-1 0.81 1.68 0.19 0.14 0.46 0.32 0.17 0.13
-2 0.40 0.31 0.16 0.11 0.24 0.16 0.10 0.09
-I(1) 0 0.52 0.40 0.48 19.41 0.42 0.35 0.40 0.34
-1 0.35 0.51 0.25 9.32 0.28 0.25 0.22 0.21
-2 0.17 0.35 0.11 1.28 0.15 0.14 0.11 0.11
-CYC 0 10.94 2.40 3.90 4.67 1.81 1.25 0.68 0.62
-1 9.05 1.72 3.33 5.88 1.47 1.04 0.54 0.51
-2 1.90 0.46 0.81 2.13 0.41 0.33 0.17 0.18
+AR 0 0.32 0.28 0.37 0.50 0.30 0.31 0.32 0.32
+1 0.26 0.31 0.29 0.47 0.26 0.30 0.28 0.31
+2 0.26 0.28 0.29 0.38 0.25 0.28 0.27 0.29
+ARMA 0 0.14 0.09 0.29 0.31 0.13 0.13 0.22 0.18
+1 0.22 0.34 0.07 0.08 0.14 0.10 0.07 0.07
+2 0.08 0.11 0.07 0.07 0.06 0.06 0.07 0.07
+TAR 0 0.59 0.35 0.73 1.07 0.65 0.56 0.56 0.45
+1 0.81 1.28 0.19 0.14 0.46 0.32 0.17 0.13
+2 0.40 0.40 0.16 0.11 0.24 0.16 0.10 0.09
+I(1) 0 0.52 0.48 0.48 19.41 0.42 0.35 0.40 0.34
+1 0.35 0.40 0.25 9.32 0.28 0.25 0.22 0.21
+2 0.17 0.13 0.11 1.28 0.15 0.14 0.11 0.11
+CYC 0 10.94 0.64 3.90 4.67 1.81 1.25 0.68 0.62
+1 9.05 0.43 3.33 5.88 1.47 1.04 0.54 0.51
+2 1.90 0.15 0.81 2.13 0.41 0.33 0.17 0.18
 Missing pattern II
-AR 0 0.38 0.70 0.55 0.78 0.33 0.36 0.43 0.39
-1 0.28 0.79 0.42 0.70 0.29 0.34 0.37 0.36
-2 0.26 0.66 0.36 0.65 0.27 0.31 0.34 0.34
-ARMA 0 0.14 1.35 0.31 0.32 0.13 0.11 0.24 0.17
-1 0.17 1.49 0.08 0.09 0.10 0.08 0.07 0.07
-2 0.15 1.05 0.07 0.08 0.09 0.07 0.07 0.07
-TAR 0 0.49 5.39 0.71 1.08 0.72 0.68 0.73 0.65
-1 0.75 6.04 0.34 0.11 0.31 0.20 0.09 0.09
-2 0.77 4.36 0.34 0.10 0.42 0.28 0.19 0.14
-I(1) 0 0.47 0.27 0.47 51.84 0.38 0.30 0.36 0.29
-1 0.19 0.43 0.18 0.56 0.17 0.14 0.14 0.13
-2 0.08 0.08 0.07 0.52 0.08 0.08 0.07 0.07
-CYC 0 7.46 7.56 1.96 15.21 12.27 6.49 1.26 1.03
-1 2.63 3.99 1.45 16.23 9.57 5.07 0.98 0.83
-2 11.18 5.09 0.79 7.90 1.28 1.04 0.28 0.31
+AR 0 0.38 0.34 0.55 0.78 0.33 0.36 0.43 0.39
+1 0.28 0.42 0.42 0.70 0.29 0.34 0.37 0.36
+2 0.26 0.39 0.36 0.65 0.27 0.31 0.34 0.34
+ARMA 0 0.14 0.44 0.31 0.32 0.13 0.11 0.24 0.17
+1 0.17 0.65 0.08 0.09 0.10 0.08 0.07 0.07
+2 0.15 0.46 0.07 0.08 0.09 0.07 0.07 0.07
+TAR 0 0.49 1.29 0.71 1.08 0.72 0.68 0.73 0.65
+1 0.75 2.06 0.34 0.11 0.31 0.20 0.09 0.09
+2 0.77 1.37 0.34 0.10 0.42 0.28 0.19 0.14
+I(1) 0 0.47 0.38 0.47 51.84 0.38 0.30 0.36 0.29
+1 0.19 0.30 0.18 0.56 0.17 0.14 0.14 0.13
+2 0.08 0.05 0.07 0.52 0.08 0.08 0.07 0.07
+CYC 0 7.46 7.04 1.96 15.21 12.27 6.49 1.26 1.03
+1 2.63 4.82 1.45 16.23 9.57 5.07 0.98 0.83
+2 11.18 0.70 0.79 7.90 1.28 1.04 0.28 0.31
 S3.2 Compositional time series
@@
 Figure S3: Scatter plots of $(x_{t-1,i},x_{t,j})$ of the original data (plotted in blue) and $(\hat{w}_{t-1,i},\hat{w}_{t,j})$ of the imputed series (plotted in red) for $i,j=1,2,3$, $i\neq j$, when the data are generated from Model 7 (AL) under missing pattern I.
-S3.3 Autoregressive modeling with groundwater data
-As our second downstream task for the groundwater data analysis, we fit an AR(6) model using data imputed by the methods considered.
-The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including imputed values.
-Then, we assess the impact of imputation quality by the out-of-sample prediction performance in predicting the remaining data.
-The missing data in the test data are removed for fair comparison.
-Figure S3 presents, in violin plots, the distribution of the logarithm of the root mean squared prediction errors ($\log$ RMSE) for each method across the 176 series under study.
-The TWI-type methods effectively alleviate some large RMSEs caused by the benchmark imputation methods and generally produce lower RMSEs.
-This shows the series imputed by TWI have marginal distributions that are stable (across time), meaning the method does not artificially induce structural shifts.
-This further corroborates the practical utility of TWI.
-Figure S3: The distributions of the logarithm of the root mean squared prediction errors (RMSEs) for each method across the 176 series, ranked by median from left to right.
+S3.3 Implementation details in groundwater data analysis
+The methods employed in the groundwater data analysis are detailed as follows.
+First, linear interpolation, mean imputation, and Kalman smoothing are implemented using the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R.
+For the iterative singular spectrum analysis (iSSA) method, we implement it via the Rssa package (Golyandina et al.,, 2018).
+There are three parameters to tune for the iSSA: the initialization, the window length, and the rank in the reconstruction step.
+Mean imputation is adopted here as the initialization, which is a common choice in practice.
+We set the window length to 6, which is roughly the maximum window length allowed when the missing ratio is high.
+Finally, to determine the rank in the reconstruction step, we employ a cross-validation approach.
+Specifically, we randomly mask 10% of the observed data.
+Then, for each rank $r\in\{1,2,\ldots,6\}$, we apply the iSSA for imputation and select the rank $r^{*}$ that yields the most accurate imputations on the masked data, measured by the mean squared error.
+Finally, the iSSA is implemented again with the full sample and the selected rank.
+For the TWI-type methods, as discussed in the main text, we first apply kernel smoothing to estimate a trend function $f(t)$.
+We use the locfit package (Loader,, 2024) with the default tricube kernel and the bandwidth is set to 12 months.
+This step can be implemented regardless of the missing pattern.
+Then the ($k$-)TWI is applied to the residuals $r_{t}=x_{t}-f(t)$, which has the same missing pattern as $\{x_{t}\}$, to obtain $\{\hat{r}_{t}\}$.
+For the TWI, the cut-off point $n_{1}$ is set to $\lfloor 0.4n\rfloor$ as in the simulation section.
+For $k$-TWI, the cut-offs are $\{0.3n,0.7n,0.5n\}$.
+The simple last observation carried forward (LOCF) is used as initialization for both TWI and $k$-TWI.
+Finally, the imputation is $\hat{x}_{t}=\hat{r}_{t}+f(t)$.